### v3

# A Persistence-Driven Edit Distance for Graphs with Abstract Weights

Abstract
In this work we define a novel edit distance for graphs considered with some weights on the edges. The metric is driven by the idea of considering graphs as topological summaries in the context of persistence and topological data analysis. Several examples related to persistent sets are presented. In case the graphs are trees, the metric can be computed with a dynamical binary linear programming approach. This framework is applied and further studied in other works focused on merge trees, where the problems of stability and merge trees estimation are also assessed.
Keywords: Topological Data Analysis, Merge Trees, Reeb Graphs, Graph Edit Distance , Tree Edit Distance , Binary Optimization
1 Introduction
The use of graphs is almost ubiquitous in science: any time some system needs to be discretised or some relationships between points/statistical units/objects formalized or represented, graphs are always involved.
For this reason a huge number of mathematical tools have been defined to work with graphs: differential calculus, distances between graphs, algebraic structures to solve graph-related problems and all other sorts of results have been developed throughout the years.
In this paper we consider graphs assuming the point of view of topological data analysis (TDA) (Edelsbrunner and Harer, 2008), i.e. we consider them as representing topological information, and develop an edit distance to analyse populations of graph-shaped topological summaries. TDA is a quite recent type of data analysis which resorts to algebraic topology (Hatcher, 2000) to extract from data information regarding each statistical unit’s shape. A typical TDA pipeline starts from a topological space and then induces a family of nested subspaces, called filtration, from which some topological information is extracted via the use of some functors, for instance simplicial homology functors (Hatcher, 2000). Generalizations of this standard procedure include using family of subspaces indexed by partially ordered sets (posets), like $\mathbb{R}^{n}$ ($n\geq 2$) (Botnan and Lesnick, 2022) and the use of Laplacians (Mémoli et al., 2022), path-connected components (Curry et al., 2022, 2021; Pegoraro, 2024a; Cavinato et al., 2022), zig-zag homology (Carlsson and De Silva, 2010), cosheaves (De Silva et al., 2016), families of functors (Turner et al., 2014) etc..
Our investigation starts from objects of the form $F:(P,\leq)\rightarrow\operatorname*{Sets}$ (with $(P,\leq)$ being a poset), which generalize persistent sets as defined in Carlsson and Mémoli (2013), and arise from very general situations like $(P,\leq)\rightarrow\operatorname*{Top}\xrightarrow{\pi_{0}}\operatorname*{Sets}$ (with $\operatorname*{Top}$ being the category of topological spaces and $\pi_{0}$ being the functor of path-connected components). For instance, if $(P,\leq)$ is given by the ordered set $\{f^{-1}((-\infty,t])\}_{t\in\mathbb{R}}$ for some $f:X\rightarrow\mathbb{R}$ we recover merge trees of real valued functions (Morozov et al., 2013; Curry et al., 2021; Pegoraro, 2024a), while if $(P,\leq)$ is the poset of open intervals in $\mathbb{R}$ we recover Reeb cosheaves (De Silva et al., 2016).
Relationships between posets and graphs have been extensively studied: for instance it is well-known that any poset can be represented with a directed acyclic graph (DAG) (Aho et al., 1972). Thus, it is no surprise that a common way to represent instances of $F:(P,\leq)\rightarrow\operatorname*{Sets}$ is via some kind of graphs, like merge trees or Reeb graphs.
Metrics to compare graphs are not a novel research topic (Billera et al., 2001; Gao et al., 2010; Bille, 2005; Feragen and Nye, 2020), but recent advances in stratified spaces, optimal transport and geometric data analysis have fostered some novel techniques to deal with such objects (Mémoli, 2011; Calissano et al., 2023). Similarly, the use of graphs (possibly labeled Yan et al. (2019)) as summaries of persistent sets has produced a series of original definitions and results (Bauer et al., 2014; De Silva et al., 2016; Di Fabio and Landi, 2016; Bauer et al., 2020; Morozov et al., 2013; Gasparovic et al., 2019; Pont et al., 2022; Sridharamurthy et al., 2020; Wetzels et al., 2022; Sridharamurthy and Natarajan, 2021; Cardona et al., 2021; Touli, 2020; Curry et al., 2022) which are all driven by these novel scenarios.
To differentiate this work from the previous ones, we quickly mention that the metric here presented, when used (up to some minor modifications) to compare merge trees, satisfies some stability properties (contrary to Pont et al. (2022); Sridharamurthy et al. (2020); Wetzels et al. (2022); Sridharamurthy and Natarajan (2021)), which are analogous to the ones of the $1$-Wasserstein distance between persistence diagrams (Edelsbrunner and Harer, 2022). In particular our metric adds all the local modifications needed to turn one graph into the other, while the other stable metrics look for the biggest modification that needs to be made, making those metrics more stable, but also less sensitive to differences on a global scale. Exactly as for the bottleneck distance and the Wasserstein distances between persistence diagrams. Formal details are given in Pegoraro (2024a). Lastly, differently from the other stable metrics, we provide, implement and employ an exact algorithm to compute the metric for trees.
Main Contributions
In this paper we define a novel edit distance apt to compare graphs which are interpreted as topological summaries. Since, increasingly often, such objects are enriched with some kind of additional information which is, to some extent, attached to vertices and edges of the graphs (Ushizima et al., 2012; Curry et al., 2022; Pegoraro, 2024b; Curry et al., 2023a), we keep our dissertation as general as possible, allowing for some abstract weights to be associated to the edges in the graph. In particular, we can articulate the contributions of this paper in two distinct points:
  1. 1.
    the main contributions are given by the definition of a distance between weighted graphs, with weights lying in certain metric spaces, and a computational framework to exactly compute the distance with binary linear optimization when the graph is a tree;
  2. 2.
    we start exploring the possibilities offered by the different weights we can attach to the edges, discussing the potential applications for these abstract weights and how can they be used to tackle different data analysis scenarios (some of which have already been investigated in other works).
The edit distance we define stems from classical edit distances (Levenshtein et al., 1966; Tai, 1979; Bille, 2005; Gao et al., 2010) but add some deep modifications, which can be intuitively motivated by the following example, represented in Figure 1.
(a) A function (orange) with its associated merge tree.
(b) Another function (brown) with its associated merge tree in black.
(c) Comparison bewtween the merge tree in Figure 1(b), and the tree obtained from the one in Figure 1(a), removing the edge $(a,b)$.
Figure 1: Plots related to the example presented in the Introduction.
Consider the following function: $f(x)=x\cdot\sin(x)$ for $x\in[0,2\pi]$ Figure 1(a). Then take the sublevel set filtration $F(t)=f^{-1}((-\infty,t])$ and the persistent set $\pi_{0}\circ F$. The information of this persistent set can be represented with a tree with four vertices and three edges, as in Figure 1(a), along with a real-valued height function defined on its vertices - with the root being at height +$\infty$+ ∞. We can associate to any edge its length, obtained as the difference in height between its endpoints. Suppose that we were to remove from the (merge) tree the edge $(a,b)$, we would obtain the red (merge) tree in Figure 1(c). Vertex $a$ is now an order $2$ vertex which represents a merging between path connected components (i.e. a local maximum in the function), which is no more happening. On a function level, this can be visualized as removing the local minimum in $2\pi$, for instance restricting the function on $[0,3/2\pi]$ Figure 1(b). Usual edit distances would regard the red and black weighted trees in the green box in Figure 1(c) as very different trees, having a different number of edges and very different weights associated to the edges. As a consequence also the black trees in the green boxes in Figure 1(a) and in Figure 1(b) would be very far apart. In our case however, we would like our metric to be much less sensitive and, for instance, not to distinguish between the black and red trees in trees Figure 1(c). For this reason, differently from other edit distances, we add the possibility to remove the vertex $a$ and join the two adjacent edges.
Interconnected Works
As already pointed out, the framework we introduce here lies the foundations for a series of developments which are carried out in separate works. Some of these developments, in turn, contain results which retrospectively validate the approach we pursue in this manuscript. In particular, we have the following group of interconnected works:
  • the present work, where the main theoretic and computational framework is established, with the aim of providing a tool to work with graphs with general weights;
  • when the weights on the edges are just positive numbers, this general framework is used by Pegoraro (2024a) to define a metric for merge trees, which can be computed with the algorithm we present here. With the help of some results contained in Pegoraro and Secchi (2024), which is more focused on the statistical aspects of such metric, Pegoraro (2024a) shows that the approach we pursue here induces a metric with stability properties which are analogous to the ones of the $1$-Wasserstein distance between persistence diagrams. In Pegoraro and Secchi (2024) the metric is also tested (using the algorithm provided in this work) on a classification pipeline for a data set which is benchmark in the field of functional data analysis; ;
  • Pegoraro (2024b), instead, exploits the present work to explore the case with weights being functions defined on the edges of the trees. The stability result obtained in Pegoraro (2024b) stems from the one proved in Pegoraro and Secchi (2024);
  • finally, Cavinato et al. (2022) considers hierarchical clustering dendrograms and carries out an analysis of a radiomic data set, putting to work the framework (theoretic and computational) we define here on real data. Therein, a stability result in terms of Gromov-Hausdorff distance between point clouds is also obtained (which relies again on the one in Pegoraro and Secchi (2024)).
Outline
Section 2 introduces the idea of editable spaces, which are the spaces we want the weights of our graphs to live in. Their properties are exploited in Section 3 to define editable graphs. Section 4 develops the novel edit distance and proves that it is well defined for graphs considered up to order $2$ vertices. Section 5 presents a series of general examples connecting persistent sets, and their generalizations, to editable graphs. Section 6 specialises the previously introduced tools to the case where the graph is a tree and obtains an algorithm to compute the distance via linear binary optimization. Section 7 concludes the paper by highlighting pros and cons of the approach along with possible further developments.
2 Editable Spaces
As already stated, the aim of the manuscript is to present a novel metric structure to work with weighted (and possibly directed) graphs, where the word weighted is intended in a broad sense: weights are given by a map associating the edges of a graph to elements in a metric space.
In order to define the metric, we need these spaces of weights to satisfy some properties which are fundamental to ensure some desired behaviours when modifying graphs.
Definition 1
A set $X$ is called editable if the following conditions are satisfied:
  • (P1)
    $(X,d_{X})$ is a metric space
  • (P2)
    $(X,\odot,0_{X})$ is a monoid (that is $X$ has an associative operation $\odot$ with zero element $0_{X}$)
  • (P3)
    the map $d_{X}(0_{X},\cdot):X\rightarrow\mathbb{R}$ is a map of monoids between $(X,\odot)$ and $(\mathbb{R},+)$: $d_{X}(0_{X},x\odot y)=d_{X}(0_{X},x)+d_{X}(0_{X},y)$.
  • (P4)
    $d_{X}$ is $\odot$ invariant, that is: $d_{X}(x,y)=d_{X}(z\odot x,z\odot y)=d_{X}(x\odot z,y\odot z)$
Note that in property (P3), $d_{X}(x\odot y,0_{X})=d_{X}(x,0_{X})+d_{X}(y,0_{X})$, implies that $x\odot y\neq 0_{X}$. Moreover (P3)-(P4) imply that the points $0_{X}$, $x$, $y$ and $x\odot y$ form a rectangle which can be isometrically embedded in an Euclidean plane with the Manhattan geometry (that is, with the norm $\mid\mid\cdot\mid\mid_{1}$): $d_{X}(x,x\odot y)=d_{X}(0_{X},y)$, $d_{X}(y,x\odot y)=d_{X}(0_{X},x)$ and $d_{X}(x\odot y,0_{X})=d_{X}(0_{X},x)+d(0_{X},y)$.
Exploiting the monoidal structure of an editable space we can define a partial order which we will use in later sections.
Proposition 1
An editable space $X$ admits a poset structure given by:
$x\leq y\text{ if and only if }\exists r\in X\text{ such that }x\odot r=y.$
Proof
  • Reflexivity: $x\odot 0=x$ and thus $x\leq x$.
  • Antisymmetry: suppose $x\leq y$ and $x\geq y$. Then $x\odot r=y$ and $y\odot q=x$ for some $r,q\in X$. But then:
    $d_{X}(y,0)=d_{X}(x\odot r,0)=d_{X}(x,0)+d_{X}(r,0)=d_{X}(y\odot q,0)+d_{X}(r,0)=d_{X}(y,0)+d_{X}(q,0)+d_{X}(r,0),$
    which implies $r=q=0$ and so $x=y$.
  • Transitivity follows from associativity of $\odot$.
Corollary 1
Consider $X$ editable space. If $X\subset V$ with $V$ being a group and the inclusion being a map of monoids, then for every $x\leq y$ in $X$, there exist one and only one $r\in X$ such that $x+r=y$.
2.1 Examples of Editable Spaces
We give some examples of editable spaces.
2.1.1 Positive Real Numbers
The set $(\mathbb{R}_{\geq 0},+,\mid\cdot\mid)$ is an editable space, as well as its subsets which are monoids, like $\mathbb{N}$. This editable space can be used to recover the classical notion of (positively) weighted graphs.
2.1.2 $1$-Wasserstein Space of Persistence Diagrams
Persistence diagrams (PDs) are among the most used topological summaries in TDA. They are used to represent sequences of homology groups, often called persistence modules, arising by applying homology functors to filtrations of topological spaces. PDs are frequently used due to their interpretable nature, being finite set of points in the plane whose $x$ and $y$ coordinates - $y\geq x$ represent the birth and death of homological features along the filtration.
There are many possible ways to introduce persistence diagrams and Wasserstein metrics, like Edelsbrunner and Harer (2022), Skraba and Turner (2020), or Divol and Lacombe (2021). In particular, Edelsbrunner and Harer (2022) is suitable for interested readers not familiar with the topic. For notational convenience here we follow the approach in Bubenik and Elchesen (2022), which we use in the proof of an upcoming proposition.
Definition 2 (adapted from Bubenik and Elchesen (2022))
Define $D(\mathbb{R}^{2}_{\Delta})$ as the set of finite formal sums of elements in:
$\mathbb{R}^{2}_{\Delta}=\{(b,d)\in R^{2}\mid 0\leq b\leq d\}.$
Similarly $D(\Delta)$ is the set of formal sums of elements in $\Delta=\{(b,b)\in R^{2}\mid 0\leq b\}$. In both sets of formal sums we have an operation given by the sum of formal sums, which is commutative and has the empty formal sum as zero element. Given $p,q\in D(\mathbb{R}^{2}_{\Delta})$ we say that $s\sim q$ if there exist $r,t\in D(\Delta)$ such that $p+r=q+t$. We collect the equivalence classes of such relationships in $D(\mathbb{R}^{2}_{\Delta},\Delta)$ and call them persistence diagrams. The sum operation and the zero element are coherently defined on the quotient.
As a consequence of 2 we can write every diagram $D$ as a sum of points in $\mathbb{R}^{2}_{\Delta}$ each with coefficient $1$. This “decomposition” is clearly non-unique as it varies between elements in the same equivalence class. With an abuse of notation we may say that $x\in D$, meaning that $x=(b,d)$ is an off-diagonal point of $\mathbb{R}^{2}_{\Delta}$ (i.e. $d>b$) appearing in such decomposition of an element (in fact, in any element) in the equivalence class of $D$.
Definition 3 (adapted from Bubenik and Elchesen (2022))
The $1$-Wasserstein metric between persistence diagrams is:
$W_{1}(D,D^{\prime})=\inf_{\gamma:D\rightarrow D^{\prime}}\parallel x-\gamma(x)\parallel_{1},$
where $\gamma$ is any bijection between an element in the equivalence class of $D$ and an element in the equivalence class $D^{\prime}$. The infimum in the definition of $W_{1}(D,D^{\prime})$ is always attained by some $\gamma$, and thus, it is always a minimum.
Proposition 2
The space $(D(\mathbb{R}^{2}_{\Delta},\Delta),W_{1},+,0)$ is editable.
2.1.3 Functions in Editable Spaces
Consider an editable space $(X,d_{X},\odot_{X},0_{X})$ and a measure space $(M,\mu)$. Then, the space of functions $L_{1}(M,\mu,X):=\{f:M\rightarrow X\mid\int_{M}d_{X}(0_{X},f(p))d\mu(p)<\infty\}/\sim$ with $\sim$ being the equivalence relationship of functions up to zero measure subsets and $\odot_{W}$ being defined pointwise - is editable. In fact, the function $p\mapsto d_{X}(0,f(p))$ is always non negative, so if properties (P3) and (P4) hold for a fixed $p\in M$, then they hold also for integrals. Call $W:=L_{1}(M,X)$, we verify (P3) as follows:
[…]
$=$ $\int_{M}d_{X}(f(p),0)+d_{X}(g(p),0)d\mu(p)=d_{W}(f,0)+d_{W}(g,0).$
And similarly for (P4).
2.1.4 Finite Products of Editable Spaces
Consider two editable spaces $X$ and $X^{\prime}$, that is $(X,\odot,0_{X})$ and $(X^{\prime},\diamond,0_{X^{\prime}})$ satisfying properties (P1)-(P4). Then $(X\times X^{\prime},\star,(0_{X},0_{X^{\prime}}))$ is an editable space, with $\star$ being the component-wise operations $\odot$ and $\diamond$, and the metric $d$ on $X\times X^{\prime}$ being the (possibly weighted) sum of the component-wise metrics of $X$ and $X^{\prime}$. For instance, $\mathbb{R}^{n}_{\geq 0}$ with the $1$-norm is an editable space.
3 Editable Graphs
Now we introduce the objects we will work with. That is, graphs, possibly directed, with a function associating to each edge a weight in a fixed editable space.
Definition 4
A directed graph is made by a couple $G=(V_{G},E_{G})$ with $V$ being the set of its vertices and $E_{G}\subseteq V\times V$ the set of its edges. An undirected graph instead is a couple $G=(V_{G},E_{G})$ with the set of edges being a subset of $V\times V/\operatorname*{Sym}(2)$ where the quotient is over the action of $\operatorname*{Sym}(n)$, the symmetric group over $n$ elements.
Note that any directed graph $(V,E)$ can be turned into an undirected graph by considering its the image of the following map: $E\hookrightarrow V\times V\rightarrow V\times V/\operatorname*{Sym}(2)$.
We use the generic word graph when we don’t want to distinguish between directed and undirected. This means implied that the word “graphs” can be replaced with “directed graphs” or “directed graphs” and any edge can be considered either as $e\in V\times V/\operatorname*{Sym}(2)$ or as $e\in V\times V$. We will point out if there is something noteworthy distinguishing the directed and undirected scenarios. The upcoming definitions should clarify how the reader should interpret this simplification process.
Definition 5
An editable (undirected/directed) graph with weights in $X$ is a (undirected/directed) graph $G=(V,E)$ such that: there is a weight function defined on the edges $\varphi_{G}:E\rightarrow X$ with $X$ being an editable space and $0\notin Im(\varphi_{G})$.
Two editable (undirected/directed) graphs $G=(V_{G},E_{G}),G^{\prime}=(V_{G^{\prime}},E_{G^{\prime}})$ are isomorphic if there is a bijection between the vertices $\eta:V_{G}\rightarrow V_{G^{\prime}}$ giving an isomorphism of (undirected/directed) graphs such that $\varphi_{G}(e)=\varphi_{G^{\prime}}(\eta(e))$, with $\eta((v_{1},v_{2})):=(\eta(v_{1}),\eta(v_{2}))$.
Example 1
Positively weighted undirected graphs and positively weighted directed graphs $G=(V,E)$ induce editable graphs with values in $X=\mathbb{R}_{\geq 0}$.
Definition 6
An editable graphs space $(\mathcal{G},X)$ is the set of editable graphs with weight values in $X$ up to editable graphs isomorphisms.
Definition 7
An editable graph is finite if $V_{G}$ and $E_{G}$ are finite sets.
Given a graph $G=(V,E)$ we indicate with $\mathcal{CW}(G)$ the CW-complex built by considering as many 00-cells as vertices in $V$ and a $1$-cell for every edge in $E$ with the gluing function induced by the extremes of the edge. See Figure 2(a) and Figure 2(b). Note that $\mathcal{CW}(G)$ is well defined (up to homeomorphism) also for undirected graphs: exchanging the extremes of the gluing maps gives an homeomorphic $\mathcal{CW}(G)$.
We have a canonical map $V\hookrightarrow\mathcal{CW}(G)$. With an abuse of notation we will often write $v\in\mathcal{CW}(G)$ meaning the image of $v\in V$ via $V\hookrightarrow\mathcal{CW}(G)$. Similarly there is a map $E\hookrightarrow\mathbf{C}(\mathcal{CW}(G))$ (with $\mathbf{C}(U)$ being the set of closed sets of a topological space $U$), induced by the image of the $1$-cells.
We need to establish the following notation.
Definition 8 (Schelp (1972))
A partial semigroup (PSG) is a set $A$ with a partial binary operation: $+:A^{\prime}\rightarrow A$, with $A^{\prime}\subset A\times A$, which is associative, i.e.:
  1. 1.
    $(x,y+z)\in A^{\prime}$ if and only if $(x+y,z)\in A^{\prime}$
  2. 2.
    if $(x,y+z)\in A^{\prime}$ then $(x,y+z)=(x,y+z)$.
We want to work with particular paths defined on the $CW$-complexes $\mathcal{CW}(G)$. To avoid creating ambiguous notation related to the word “paths” and in analogy with the literature on real trees (Evans, 2008), we give the following definition.
Definition 9
Consider a graph $G=(V,E)$. A segment in $G$ is an embedding $s:[0,1]\rightarrow\mathcal{CW}(G)$ such that $s(0),s(1)\in V\hookrightarrow\mathcal{CW}(G)$ and there is a path on $G$, $\gamma_{s}=\{(v_{0},v_{1}),(v_{1},v_{2}),\ldots,(v_{n-1},v_{n})\}\subset E$ such that: $v_{0}=s(0)$, $v_{n}=s(1)$ and $Im(\gamma_{s}\hookrightarrow\mathbf{C}(\mathcal{CW}(G)))=Im(s)$.
For two segments $s,s^{\prime}$ we have $s\sim s^{\prime}$ if $Im(s)=Im(s^{\prime})$. The internal points of a segment are defined as $s^{o}:=Im(s)-\{s(0),s(1)\}$. We collect all equivalence classes of segments in $G$ in the set $\mathcal{S}(G)$.
Remark 1
We have a map $E\hookrightarrow\mathcal{S}(G)$ induced by the attachment maps of the $1$-cells of the $CW$-complex: $e\mapsto s_{e}\in\mathcal{S}(G)$ so that $s_{e}([0,1])=e$. Note that in case $G$ is an undirected graph, $s(t)\sim s(1-t)$.
Proposition 3
$\mathcal{S}(G)$ can be given a PSG structure via paths composition. Consider $s,s^{\prime}\in\mathcal{S}(G)$ such that:
  • $s(1)=s^{\prime}(0)$;
  • $Im(s)\cap Im(s^{\prime})=s(1)$;
then $s+s^{\prime}:=s^{\prime\prime}:[0,1]\rightarrow\mathcal{CW}(G)$ with $s^{\prime\prime}(t)=s(2t)$ for $t\in[0,1/2]$ and $s^{\prime\prime}(t)=s^{\prime}(2t-1)$ for $t\in[1/2,1]$.
(a) A directed graph.
(b) The CW-complex associated to the graph in Figure 2(a).
(c) The sum of segments by paths composition.
(d) Two different segments $s,s^{\prime}$ with their open stars $\mathcal{N}(s^{o})$ and $\mathcal{N}(s^{\prime o})$. We have $\mathcal{N}(s^{\prime o})=s^{\prime o}$.
Figure 2: Plots related to Section 3.
See also Figure 2(c). Note that, as a consequence of 3, the set $\mathcal{S}(G)$ is almost a directed structure on the space $\mathcal{CW}(G)$ (as defined in Fajstrup et al. (2016)), missing just the constant paths.
We state a couple of observations in the following remark.
Remark 2
$\mathcal{S}(G)$ is generated by $s_{e}\in\mathcal{S}(G)$: every $s\in\mathcal{S}(G)$ can be uniquely written as $\sum_{e\in Im(s)}s_{e}$. Thus, the weight function of any editable (directed) graph $\varphi_{G}$ can be extended to a PSG map $\varphi_{G}:\mathcal{S}(G)\rightarrow X$.
Lastly we define the open star of a subset of the $CW$-complex associated to a graph. This definition is standard for simplicial complexes and we extend it roughly by looking at the open star of the graph projected onto the $CW$-complex.
Definition 10
Given $U\subset\mathcal{CW}(G)$, the open star of $U$, $\mathcal{N}(U)\subset\mathcal{CW}(G)$, is the image under $V\hookrightarrow\mathcal{CW}(G)$ of all the vertices $v\in U$, united with $s_{e}^{o}$ for all the edges $e=(v_{1},v_{2})$ such that $v_{1}\in U$ or $v_{2}\in U$.
Particularly important segments are the ones whose open star in $\mathcal{CW}(G)$ is equal to the (open) segment itself. See also Figure 2(d).
Definition 11
We define $\mathcal{O}(G)\subset\mathcal{S}(G)$ as the set of all segments $s\in\mathcal{S}(G)$ such that $\mathcal{N}(s^{o})=s^{o}$.
4 Edits of Graphs
Now we present the main contribution of the paper.
The approach we follow is to define a distance inspired by Edit Distances (Tai, 1979), but with substantial differences in the edit operations. The philosophy of these distances is to allow certain modifications of the base object, called edits, each being associated to a cost, and to define the distance between two objects as the minimal cost that is needed to transform the first object into the second with a finite sequence of edits. In this way, up to properly setting up a set of edits, one can formalize the deformation of a graph modifying the local information induced by the weight function defined on the edges and the graph structure itself. On top of that, edit distances frequently enjoy some decomposition properties which simplify the calculations (Hong et al., 2017), which are notoriously very heavy (Hein et al., 1995). We also point out that, in literature, graphs are often considered with weights being real numbers and avoid modelling the weight space as we do with the definition of editable spaces.
Given an editable (directed) graphs space $(\mathcal{G},X)$, with $(X,\odot,0_{X})$ editable space, we can define our edits.
  • We call shrinking of an edge a change of the weight function. The new weight function must be equal to the previous one on all edges, apart from the “shrunk” one. In other words, for an edge $e$, this means changing the value $\varphi(e)$ with another non zero value in $X$.
  • A deletion is an edit with which an edge $e$ is removed from the graph. If $e=(v_{1},v_{2})$ is the only edge between $v_{1}$ and $v_{2}$, then we also remove either $v_{1}$ or $v_{2}$. The result is well defined up to (directed) graph isomorphism. The inverse of the deletion is the insertion of an edge along with, potentially, the addition of a novel vertex. We can insert an edge between two existing vertices or at a vertex $v$ specifying the name of the new neighbour of $v$, the neighbours of the newly added vertex (that can be either none, or any portion of the neighbours of $v$). Clearly we always need to define the value of the weight function on the new edge.
  • Lastly, consider $s=\sum s_{e}\in\mathcal{O}(G)$. We can remove all edges and vertices in $s^{o}$ and add to $E_{G}$ the edge $(s[0],s[1])$ with, eventually, the direction coherently induced. The weight of the new edge is already encoded in $w:\mathcal{S}(G)\rightarrow X$ and we have $w(s)=\bigodot_{e\in Im(s)}w(e)$. This transformation is called the ghosting of the vertices in $Im(s)$. Its inverse transformation is called the splitting of an edge $e=(v_{1},v_{2})$ and replaces $e$ with some novel vertices $v^{\prime}_{1},\ldots,v^{\prime}_{n}$ and edges $(v_{1},v^{\prime}_{1}),(v^{\prime}_{1},v^{\prime}_{2})\ldots,(v^{\prime}_{n},v_{2})$ which form a segment in the new graph $G^{\prime}$ such that $s\in\mathcal{O}(G^{\prime})$ with $s(0)=v_{1}$, $s(1)=v_{2}$. Note that any $s_{e}\in\mathcal{O}(G)$ induces a trivial ghosting.
(a) Starting and target weighted graphs.
(b) Deletions in red.
(c) Ghostings in yellow.
(d) Shrinkings in green.
(e) Splittings in yellow.
(f) Insertions in red.
Figure 3: (b)$\rightarrow$(e) form an edit path made from the left graph in Figure 3(a) to the right one. In each figure the edges involved in the editing are highlighted with different colors. In the following plot such vertices return black. This edit path can clearly be represented with a mapping - Section 4.2 - which first deletes all the red vertices in Figure 3(b), then ghosts all the yellow segments in Figure 3(c), then shrinks the green edges in Figure 3(d), split edges by means of the yellow vertices in Figure 3(c) and inserts all the red edges in Figure 3(f).
An editable graph $G$ can be edited to obtain another graph, on which one can apply a new edit to obtain a third graph and so on. One can think of this as composing two edits $e_{0}$, $e_{1}$ which are not defined on the same graph, since the second edit is defined on the already edited graph. This is what we mean by composition of edits. Any finite composition of edits is referred to as an edit path. The notations we use are functional notations, even if the edits are not operators, since an edit is not defined on the whole space of graph but on a single graph. For example $e_{1}\circ e_{0}(G)$ means that $G$ is edited with $e_{0}$, and then $e_{0}(G)$ with $e_{1}$. See Figure 3 for an example of an edit path between two graphs whose weights are edges lenghts - Section 2.1.1.
Remark 3
Even if shrinking and deletions are classical edits, with shrinking being usually referred to as relabeling, the ghosting edit is completely unusual and we are not aware of any previous work employing it. Its definition is largely driven by the TDA perspective presented in the Introduction and it is fundamental for the stability results contained in Pegoraro and Secchi (2024) and Pegoraro (2024a) - dealing with the case of $G$ being a tree.
We also give the following definition.
Definition 12
Graphs are equal up to order $2$ vertices if they become isomorphic after applying a finite number of ghostings or splittings. We write $(G,\varphi_{G})\cong_{2}(G^{\prime},\varphi_{G^{\prime}})$.
12 induces an equivalence relationship which identifies the set of graphs inside $(\mathcal{G},X)$ that we want to treat as equal. We call $(\mathcal{G}_{2},X)$ the space of equivalence classes of graphs in $(\mathcal{G},X)$, equal up to order $2$ vertices.
4.1 Costs of Edit Operations
Now we associate to every edit a cost, that is, a length measure in the space $(\mathcal{G},X)$. The costs of the edit operations are defined as follows:
  • if, via shrinking, an edge goes from weight $x$ to weight $y$, then the cost of such operation is $d_{X}(x,y)$;
  • for any deletion/insertion of an edge with weight $x$, the cost is equal to $d_{X}(x,0_{X})$;
  • the cost of ghosting/splitting operations is always zero. This is coherent with the properties of editable spaces: consider a ghosting merging two edges separated by an order two vertex, with $x$ being the weight of one edge, and $y$ the cost of the other. The total weight of the two edges is the same before and after the ghosting: $\mid d_{X}(x\odot y,0_{X})-d_{X}(x,0_{X})-d_{X}(y,0_{X})\mid=0$.
Definition 13
Given two graphs $G$ and $G^{\prime}$ in $(\mathcal{G},X)$, define:
  • $\Gamma(G,G^{\prime})$ as the set of all finite edit paths between $G$ and $G^{\prime}$;
  • $cost(\gamma)$ as the sum of the costs of the edits for any $\gamma\in\Gamma(G,G^{\prime})$;
  • the graph edit distance as:
    $d_{E}(G,G^{\prime})=\inf_{\gamma\in\Gamma(G,G^{\prime})}cost(\gamma)$
By definition the triangle inequality and symmetry must hold, but, up to now, this edit distance is intractable; one would have to search for all the possible finite edit paths which connect two graphs in order to find the minimal ones. On top of that, having an edit which is completely “for free”, it is not even obvious that $d_{E}(G,G^{\prime})>0$ for some graphs. However, it is clear that $d_{E}$ induces a pseudo-metric on classes of graphs up to order two vertices.
4.2 Mappings
A mapping between the editable graphs $G=(V_{G},E_{G})$ and $G^{\prime}=(V_{G^{\prime}},E_{G^{\prime}})$ is a set $M\subset\mathcal{S}(G)\times\mathcal{S}(G^{\prime})$ satisfying some constraints. Before listing such properties, we introduce the following notation: (1) given two sets $A,B$, $A-B:=\{a\in A\mid a\notin B\}$ (2) for any $M\subset\mathcal{S}(G)\times\mathcal{S}(G^{\prime})$:
$\mathcal{E}_{M}:=\coprod_{e\subset Im(s),s\in\pi_{G}(M)}e$
with $\pi_{G}:\mathcal{S}(G)\times\mathcal{S}(G^{\prime})\rightarrow\mathcal{S}(G)$ being the projection of the cartesian product. We then have the induced maps $\eta_{G}^{M}:\mathcal{E}_{M}\rightarrow E_{G}$. Lastly, $M$ is a mapping if:
  • (G1)
    the maps $\eta_{G}^{M}$ and $\eta_{G^{\prime}}^{M}$ are injective;
  • (G2)
    if $s\in\pi_{G}(M)$ such that $s\in\mathcal{S}(G)-E_{G}$, then $\mathcal{N}(s^{o})-s^{o}\subset G-\pi_{G}(M)$;
  • (G3)
    if we delete all edges in $E_{G}-Im(\eta_{G}^{M})$ and do all the ghostings induced by $s\in\pi_{G}(M)-E_{G}$ (and analogously for $G^{\prime}$), then $M$ induces a graph isomorphism between the resulting subgraphs of $G$ and $G^{\prime}$.
Let’s try to unveil the meaning of (G1)-(G3). The main idea is that any couple in a mapping $M$ identifies a segment of $G$ and a segment of $G^{\prime}$ which are paired and turned one in the other by the mapping.
Condition (G1) is asking that every edge in $G$ and $G^{\prime}$ is at most within one segment which is coupled with a segment of the other graph. We refer to such edges as coupled edges. The edges which are not coupled will be deleted. Condition (G2) ensures that once we remove all edges which are not coupled, and so which are deleted, then we are in the position to make all the ghostings to turn all the segments $s\in\pi_{G}(M)-E_{G}$ (and $s^{\prime}\in\pi_{G^{\prime}}(M)-E_{G^{\prime}}$), in edges of a graph. Lastly, (G3) forces the couples in $M$ to induce a graph isomorphism between the resulting graphs.
Remark 4
Properties (G1)-(G3) are not canonical properties of mappings, and descend from the fact that we have introduced the ghosting operation, which is a novelty of this work. To better compare properties (G1)-(G3) with more classical edit distances we refer to Section 6, where we report (G1)-(G3) for the special case in which the graph is indeed a tree. Mappings in fact are often used to take advantage of the properties of trees to reduce the computational cost of matching two trees.
In what follows we exploit (G1)-(G3) to parametrize a set of edit paths in the editable graphs space starting from $G$ and ending in $G^{\prime}$, which are collected under the name $\gamma_{M}$. We call:
  • $\gamma_{d}^{G}$ a path made by the deletions to be done in $G$, that is, the edges $E_{G}-Im(\eta_{G}^{M})$, executed in any order. So we obtain the graph $G^{M}_{d}=\gamma_{d}^{G}(G)$, which, instead, is well defined and not depending on the order of the deletions.
  • One then proceeds with ghosting all the segments $s\in\pi_{G}(M)-E_{G}$, in any order, getting a path $\gamma^{G}_{g}$ and the graph $G_{M}:=\gamma^{G}_{g}\circ\gamma^{G}_{d}(G)$.
  • By property (G3), $M$ induces an isomorphism between the two graphs $G^{\prime}_{M}$ (defined in the same way as $G_{M}$, but starting from $G^{\prime}$) and $G_{M}$. So one can shrink the edges of $G_{M}$ onto the edges of $G^{\prime}_{M}$ as described by $M$. The composition of all the shrinkings, executed in any order is an edit path $\gamma_{s}^{G}$.
By construction:
$\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G}(G)=G^{\prime}_{M},$
and:
$(\gamma_{d}^{G^{\prime}})^{-1}\circ(\gamma_{g}^{G^{\prime}})^{-1}\circ\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G}(G)=G^{\prime}.$
where the inverse of an edit path is thought as the composition of the inverses of the single edit operations, taken in the inverse order.
Lastly, we call $\gamma_{M}$ the set of all possible edit paths:
$(\gamma_{d}^{G^{\prime}})^{-1}\circ(\gamma_{g}^{G^{\prime}})^{-1}\circ\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G},$
obtained by changing the order in which the edit operations are executed inside $\gamma_{d}$, $\gamma_{g}$ and $\gamma_{s}$. Observe that, even if $\gamma_{M}$ is a set of paths, its cost is well defined:
$cost(M):=cost(\gamma_{M})=cost(\gamma_{d}^{G})+cost(\gamma_{s}^{G})+cost(\gamma_{d}^{G^{\prime}}).$
See Figure 3 for an example of a mapping between weighted graphs.
[…]
Theorem 1 (Main Theorem)
Given two editable graphs $G$ and $G^{\prime}$, for every finite edit path $\gamma$, there is a mapping $M\in\operatorname*{Mapp}(G,G^{\prime})$ such that $cost(M)\leq cost(\gamma)$.
A first corollary immediately follows.
Corollary 2
For any two finite editable graphs $G$ and $G^{\prime}$, the set $\operatorname*{Mapp}(G,G^{\prime})$ is finite. Thus we have the following well defined pseudo-metric between finite editable graphs:
$d_{E}(G,G^{\prime})=\inf\{cost(\gamma)\mid\gamma\in\Gamma(G,G^{\prime})\}=\min\{cost(M)\mid M\in\operatorname*{Mapp}(G,G^{\prime})\}$
which we will refer to as the edit distance between $G$ and $G^{\prime}$.
A second corollary is obtained observing that, if a mapping has cost equal to zero, then it must contain only ghostings.
Corollary 3
Given $G$ and $G^{\prime}$ editable graphs, $d_{E}(G,G^{\prime})=0$ if and only if $G$ and $G^{\prime}$ are equal up to order $2$ vertices. In other words $d_{E}$ is a metric for graphs considered up to order $2$ vertices.
5 A Discussion on the Use of Graphs and Different Editable Spaces
In this section we report a discussion featuring some key examples which regard editable graphs with values in different editable spaces induced from persistent sets (Carlsson and Mémoli, 2013; Curry, 2018; Curry et al., 2022). Persistent sets are objects of great interest in TDA and represent a (non-equivalent) alternative to persistence modules/diagrams when working with information collected from path connected components of filtrations. We stress again that this is to be intended as a roadmap, the outlining of a general idea and not as a formal and coherent description on how one can treat persistent sets via directed editable graphs. In fact, even in the simplest case, i.e. merge trees, there are non-trivial details which need to be taken care of when bridging between persistent sets and editable graphs (see Section 5.5 and Pegoraro (2024a)). Coherently, we explicitly point out which questions and problems should be assessed by future works.
5.1 Finite Posets, DAGs and Transitive Reductions
In this section, following previous works on graphs and posets, we obtain a graph representation of a poset, which we than employ to represent poset-indexed functors.
Consider a finite partially ordered set $(P,\leq)$. It is well known that one can associate to $P$ a unique (up to directed graph isomorphism) directed acyclic graph (DAG) $G_{P}$ so that $G_{P}$ is in bijection with the objects in $P$ and $(a,b)\in E_{G_{P}}$ if and only if $a\leq b$. Moreover, the graph $G_{P}$ can be simplified without losing ordering information.
Definition 14 (adapted from Aho et al. (1972))
A directed graph $G=(V,E)$ is called transitive if for every segment $s\in\mathcal{S}(G)$, $(s(0),s(1))\in E$. For every directed graph $G=(V,E)$ we define its transitive closure as the graph $G^{T}:=(V,E_{G^{T}})$, with $E_{G^{T}}=\bigcap\{E^{\prime}\subset V\times V\mid E\subset E^{\prime},(V,E^{\prime})\text{ transitive graph}\}$.
Theorem 2 (adapted from Aho et al. (1972))
For every $G=(V,E)$ DAG there is a unique graph $G^{t}=(V,E_{G^{t}})$ such that:
  • $(G^{t})^{T}=G^{T}$;
  • for every $E^{\prime}\subsetneq E_{G^{t}}$, $(V,E^{\prime})^{T}\neq G^{T}$.
Such graph is called the transitive reduction of $G$ and it is given by:
$E_{G^{t}}:=\bigcap\{E^{\prime}\subset V\times V\mid(V,E^{\prime})^{T}=G^{T}\}.$
Given a finite poset $P$, we call $\mathcal{G}(P)$ the transitive reduction of the associated DAG. See Figure 5 for some examples.
5.2 From Persistent Sets to Editable Graphs
Now, first we associate an editable graph structure to a poset $P$. And then we use this to work with functors of the form $S:P\rightarrow\operatorname*{Sets}$, which we call persistent sets, generalising terminology previously introduced by (Carlsson and Mémoli, 2013; Curry, 2018).
If $P$ is finite, we propose two general ways in which we can associate weights to $\mathcal{G}(P)$, relying just on the structure of $P$:
  1. 1.
    suppose now that the poset $P$ is also endowed with a pseudo-metric (see Burago et al. (2022) for a definition) $d_{P}:P\times P\rightarrow\mathbb{R}_{\geq 0}$. Suppose moreover that any time $a<b$, then $d_{P}(a,b)>0$. Then we can associate to $P$ an editable graph $(\mathcal{G}(P),E_{\mathcal{G}(P)},\varphi_{\mathcal{G}(P)})$, with $\varphi_{\mathcal{G}(P)}:E_{\mathcal{G}(P)}\rightarrow\mathbb{R}_{>0}$ defined as $\varphi_{\mathcal{G}(P)}((a,b))=d_{P}(a,b)$. Note that for any poset $P$ with a monotone increasing function $f:P\rightarrow\mathbb{R}$ we can define $d_{P}(a,b):=\mid f(a)-f(b)\mid$ and this is a pseudo metric which satisfies our requirements since, if $a<b$, then $f(a)<f(b)$;
  2. 2.
    instead, suppose there is a map of posets $f:P\rightarrow X$, with $X$ editable space, and the poset structure on $X$ is as defined in 1. Suppose moreover we are in the situation to apply 1. Then, for every edge $(a,b)\in\mathcal{G}(P)$ there is one and only one $r\in X$ with $f(a)+r=f(b)$. Thus we can define $\varphi_{\mathcal{G}(P)}((a,b))=r$.
As further step consider the case of functors of the form $S:P\rightarrow\operatorname*{Sets}$, with $P$ being a finite poset. We might refer to $P$ as the indexing poset. The functor $S$ can be used to push forward onto its image the partial order we have on $P$. We do so associating to $S$ a poset $D_{S}$, called the display poset of $S$, defined as follows.
Definition 15 (adapted from Curry et al. (2022))
Given a functor $S:P\rightarrow\operatorname*{Sets}$, with $P$ being a poset, the display poset of $S$ is defined as the set of couples $D_{S}:=\{(p,s)\mid p\in P,s\in S(p)\}$. This is a poset with $(p,s)\leq(p^{\prime},s^{\prime})$ if and only if $p\leq p^{\prime}$ and $S(p\leq p^{\prime})(s)=s^{\prime}$. For every display poset $D_{S}$ wee have a monotone increasing projection $h:D_{S}\rightarrow P$ defined by $(p,s)\mapsto p$.
Proposition 4
The map $h:D_{S}\rightarrow P$ induces a graph homomorphism (a function between vertices that also sends edges to edges) between $\mathcal{G}(D_{S})$ and $\mathcal{G}(P)$.
Proof  By construction, any time we have $(p,s)\leq(p^{\prime},s^{\prime})\leq(p^{\prime\prime},s^{\prime\prime})$, this also implies $p\leq p^{\prime}\leq p^{\prime\prime}$. Viceversa, any time we have $(p,s)\leq(p^{\prime\prime},s^{\prime\prime})$ and $p\leq p^{\prime}\leq p^{\prime\prime}$, by functoriality there exists also $(p^{\prime},s^{\prime})$ such that $(p,s)\leq(p^{\prime},s^{\prime})\leq(p^{\prime\prime},s^{\prime\prime})$.
As a consequence, we can define $\varphi_{\mathcal{G}(D_{S})}:=h\circ\varphi_{\mathcal{G}(P)}$.
5.3 Decorated Merge Trees: Function Valued Weights
We now use the previous construction to tackle a particular kind of persistent sets.
[…]
Definition 16 ((Patel, 2018))
A persistent set $S$ is called constructible if there is a finite collection of real numbers $\{t_{1}<t_{2}<\ldots<t_{n}\}$ such that:
  • $S(t<t^{\prime})=\emptyset$ for all $t<t_{1}$;
  • for $t,t^{\prime}\in(t_{i},t_{i+1})$ or $t,t^{\prime}>t_{n}$, with $t<t^{\prime}$, then $S(t<t^{\prime})$ is bijective.
The set $\{t_{1}<t_{2}<\ldots<t_{n}\}$ is called critical set and $t_{i}$ are called critical values. If $S(t)$ is always a finite set, then $S$ is a finite persistent set.
Note that we can always consider a minimal set of critical values (Pegoraro, 2024a), which constitute the critical poset of the persistent set. We also add a last requirement for the persistent sets we want to work with.
Definition 17 (Pegoraro (2024a))
A constructible persistent set $S$ is regular if for every critical value $t_{i}$ and for every $\varepsilon>0$ small enough, $S(t_{i}\leq t_{i}+\varepsilon)$ is a bijective map.
At this point we have the following immediate result, which we state without proof.
Proposition 5
Given two finite regular persistent sets $S$ and $S^{\prime}$, then $S\simeq S^{\prime}$ if and only if:
  • they have the same minimal set of critical values $C=\{t_{1},\ldots,t_{n}\}$
  • $S_{\mid C}\simeq S^{\prime}_{\mid C}$, with $S_{\mid C}:C\hookrightarrow\mathbb{R}\rightarrow\operatorname*{Sets}$.
For the interested reader, we have very close relationships between the graph $\mathcal{G}(D_{S_{\mid C}})$ and the merge tree of $S$. Note that the upcoming proposition also further justifies our idea of working up to order $2$ vertices
Proposition 6
Consider a finite regular persistence set $S$ and its minimal set of critical values $C$. If $S(t)$ has cardinality $1$ for $t$ big enough, then the graph $\mathcal{G}(D_{S_{\mid C}})$ is equivalent up to order $2$ vertices (as a directed graph) to the merge tree of $S$, upon removing from the merge tree the edge that goes to infinity. In particular, $\mathcal{G}(D_{S_{\mid C}})$ is a tree.
As a consequence, if we then use item 1. in Section 5.2 with $P$ equal to $C$, the set of critical values of $S$, we obtain an editable graph with weights in $\mathbb{R}_{\geq 0}$ is which is very closely related to merge trees. In (Pegoraro, 2024a) this approach is used to represent merge trees and to adapt $d_{E}$ to such objects, obtaining an edit distance $d_{E}$ for merge trees.
Moreover, putting together results in Pegoraro (2024a) it is shown that:
$d_{I}(T_{f},T_{g})\leq d_{E}(T_{f},T_{g})\leq 2\cdot\#T_{f}\cdot\#T_{g}\parallel f-g\parallel_{\infty}$ (1)
with:
  • $f,g$ being two tame (Chazal et al., 2016) real valued functions defined on the same path-connected topological space;
  • $T_{f}$ and $T_{g}$ are the merge trees associated to the sublevel set filtrations of $f$ and $g$;
  • $d_{I}$ is the interleaving distance between merge trees (Morozov et al., 2013);
  • $d_{E}$ is an adaptation of metric here induced to compare merge trees represented by means of editable graphs, essentially truncating the edge at infinity (with some additional details, see Pegoraro (2024a));
  • #$\#T_{f}$ being the cardinality of the set of vertices of $T_{f}$.
Note that Equation 1 represents a stability property in line with our expectations: since our metric is a summation of all the costs of the local modifications needed to align two trees we have that the to control the distance between trees with the sup norm between functions, we need a constant that scales linearly with the size of the trees. The interleaving distance, instead, is focused on determining the maximal local modifications to turn one tree into the other and it is well known to be universally stable, in the sense that is the biggest distance which is always bounded by the sup norm between the functions. Moreover, the $1$-Wasserstein distance between persistence diagrams satisfies and analogous relationship w.r.t. the bottleneck distances.
As a consequence, our metric is better at picking up global differences and not just local differences (analogously to Wasserstein distances w.r.t bottleneck distance for persistence diagrams). On the other hand, this also means that is it more sensitive to noise. For more details and examples see Pegoraro (2024a).
[…]
  1. 1.
    diagram or Betti numbers decorated merge trees, on the same line of thought of Curry et al. (2022, 2023b);
  2. 2.
    merge trees decorated with the measure of sublevel sets.
We illustrate this with two qualitative examples: first we consider the datum of a point cloud a then a function defined on a open subset of $\mathbb{R}^{n}$.
Let $Y^{k}\subset\mathbb{R}^{n}$ be a finite set with $k$ points. A widely employed filtration is:
$Y_{t}^{k}:=\cup_{p\in Y^{k}}B_{t}(p)$
$B_{t}(p)$ being the metric ball centred in $p$, which is also called the offset filtration. Any point $D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$ is a couple $(t,U)$, with $t\in\mathbb{R}$ and $U$ being a path connected component of $Y_{t}^{k}$. Then, one collect some topological information about $U$ and consider a function $f$ with values in $X=\mathbb{R}^{N}_{\geq 0}$, for some $N<n$, defined by:
$f((t,U))=(\beta_{0}(U),\beta_{1}(U),\ldots,\beta_{N}(U)),$
with $\beta_{i}(U)=\dim\text{H}_{i}(U,\mathbb{R})$ being the $i$-th Betti number of $U$ clearly $\beta_{0}(U)=1$. For an example see Pegoraro (2024b). An alternative and richer way to capture shape-related information about $U$ is to consider the point cloud: $U_{k}:=Y^{k}\cap U$ and then build $\text{PD}_{q}(U_{k})$, the persistence diagram of the offset filtration of $U_{k}$ using $q$-dimensional homology. Consequently, one can define $f$ with values in $X=D(\mathbb{R}^{2}_{\Delta},\Delta)$, the space of persistence diagrams with the $1$-Wasserstein distance, defined by:
$f((t,U))=\text{PD}_{q}(U_{k}).$
See Figure 4(a) and Figure 4(b) for a visual example. Observe how, at each value $t$, the merge tree gives a decomposition of the point cloud and the local shape of its clusters is captured by persistence diagrams.
Consider now a Lebesgue measurable function $g:W\rightarrow\mathbb{R}$, with $W$ open in $\mathbb{R}^{n}$. Build $Y_{t}=f^{-1}((-\infty,t])$, the sublevel set filtration of $f$. And consider $D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$. This time another shape-related information that we can estract from a couple $(t,U)\in D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$, is the Lebesgue measure of $U$. And thus define:
$f((t,U))=\mathcal{L}(U),$
with $\mathcal{L}(U)$ being the Lebesgue measure of $U$. For an example of situations that can be tackled with this approach see Figure 4(c) and Figure 4(d). Note that the representation of functional data obtained with the volume-decorated merge trees, lies in between information captured by $L_{p}$ metrics and purely topological information: with merge trees one would not be able to distinguish the two functions in Figure 4(c) and Figure 4(d) (they have the same merge trees), but with $L_{p}$ metrics one would not capture any of the shape similarities between the functions. Moreover, functional data analysis approaches capturing the shape of functions - e.g. aligning functions with re-parametrizations, see for instance the Special Section on Time Warpings and Phase Variation on the Electronic Journal of Statistics, Vol 8 (2) - are difficult to develop over $\mathbb{R}^{n}$, $n\geq 2$.
[…]
$d_{E}(T_{f},T_{g})\leq(K+\max\{\text{diam}Y,\text{diam}Y^{\prime}\})\cdot 4\cdot\#Y\cdot\#Y^{\prime}\cdot d_{G-H}(Y,Y^{\prime}),$
with $d_{H}(Y,Y^{\prime})$ being the Gromov-Hausdorff distance between compact metric spaces. We believe that the first two approaches have the potential to be powerful shape summaries to distinguish between point clouds, as also shown with several examples by Curry et al. (2022, 2023b), employing similar (though not metrically equivalent) tools. Instead, for more details and examples on the use of functional weights for editable graphs, see Pegoraro (2024b). Note that between the metric $d_{E}$, applied to this context, and the metric developed in Curry et al. (2022), hold similar considerations to the ones we made comparing $d_{E}$ and the interleaving distance between merge trees.
(a) A point cloud (left), with the thickening giving the off-set filtration at some particular value (dashed horizontal cyan line) and the merge tree of the filtration (right). At the chosen value the filtration has three path connected components, shown in different colours, each with different shapes, characterized by the (qualitative) associated persistence diagram.
(b) Same situation as in Figure 4(a), but at a different filtration value. At the chosen value - bigger than in Figure 4(a), the filtration has two path connected components, shown in different colours, each with different shapes, characterized by the (qualitative) associated persistence diagram. At this resolution, the brown cluster shows a more persistent ringed shape.
(c) A real valued function defined on a square with four distinct local minima, associated to downward peaks with different shapes. The associated decorated merge tree (right) accordingly shows four branches (black) and, in red, the (qualitative) volume of the associated downward peaks included in the sublevel sets, which increases with the value of the filtration.
(d) Same situation as in Figure 4(c), but for a different image. We clearly see that the four downward peaks have very different sizes, and that is reflected by the very different volume profiles attached to the branches of the merge tree. Note that the two images have the same (non-decorated) merge trees.
Figure 4: Plots related to Section 5.3.
5.4 Persistent Sets with Embedded Posets: Vector Valued Weights
Here we consider the case of a persistent set $F:P\hookrightarrow\mathbb{R}^{n}\rightarrow\operatorname*{Sets}$ and focus on the approach described by item 2. in Section 5.2. Where $P$ is a finite poset and the poset structure in $\mathbb{R}^{n}$ is induced by $r=(r_{1},\ldots,r_{n})\leq r^{\prime}=(r^{\prime}_{1},\ldots,r^{\prime}_{n})$ if $r_{i}\leq r^{\prime}_{i}$ for all $i$. This is the kind of situations that works studying multiparameter persistence care about, by making finiteness assumptions like finitely presentable persistence modules (Oudot and Scoccola, 2024) - that is, modules on $\mathbb{R}^{n}$ which are determined by their values on a finite sub-poset of $\mathbb{R}^{n}$. We associate to $F$ a unique editable graph built in this way:
  • we take $P$, build the associated DAG and take its transitive reduction $\mathcal{G}(P)$;
  • build the weight function $\varphi_{\mathcal{G}(P)}$ as in Section 5.2, item 2), so that each edge between $a,b\in\mathbb{R}^{n}_{\geq 0}$, with $a\leq b$ we associate the vector $b-a$;
  • finally we build the display poset of $F_{\mid P}$, which we call $D_{F}$, take its transitive reduction $\mathcal{G}(D_{F})$ and build the weight function $\varphi_{\mathcal{G}(D_{F})}$ by pushing forward $\varphi_{\mathcal{G}(P)}$. See Figure 5 for a detailed example.
We believe that the metric $d_{E}$ is a natural metric to employ to compare these kind of graph representations for multidimensional persistent sets. In fact, the metric exhibits some natural behaviour, for instance as in Figure 6 and, in case $n=1$, and has proven stability results - which we reported in previous sections.
(a) A finite sub-poset of $\mathbb{R}^{2}$.
(b) The transitive reduction of the poset in Figure 5(a).
(c) The DAG associated to a persistent set on the poset in Figure 5(a). Arrows are colored consistently with the ones in the poset in Figure 5(a) and represent the weights associated to the corresponding edges. The maps between sets are represented by colors and follow (row-wise and column-wise) the elder rule.
(d) The editable graph associared to the transitive reduction of the DAG in Figure 5. As in Figure 5, arrows represent weights in $\mathbb{R}^{2}_{\geq 0}$
Figure 5: Plots related to Section 5.4.
(a) Another DAG representing a persistent set over a poset in $\mathbb{R}^{2}$.
(b) Starting from the DAG in Figure 6(a), with two ghosting edits, we can replace the two paths around the square associated to the pink sets, with the two diagonal edge. Thus the DAG in Figure 6(a) and this one are identified by our edit distance.
Figure 6: Additional plots related to Section 5.4.
5.5 Problems to be Faced
In the previous sections we have laid out some ideas which suggest how different persistent sets can be represented via editable graphs with weight functions taking values in different editable spaces. As pointed out at the beginning of Section 5, these ideas present some shortcomings which need to be overcome and some formal results which need to be established in order for the ideas to be properly applied. In this section we present the major issues affecting the proposed approach: some of them are already tackled and solved in other works, while other are left open and serve as invitations for future investigations.
In the same way a point cloud is represented by the pairwise distances between its points only up to isometry, similarly, collecting distances or vectors as weights for the edges does not provide a map from persistent sets to graphs that can be inverted up to natural isomorphisms. Consider the case of a one dimensional persistent set $\pi_{0}\circ Y:\mathbb{R}\rightarrow\operatorname*{Sets}$ and define for every $h\in\mathbb{R}$ the filtration $Y_{h}(t):=Y(t+h)$. If we then proceed by building the display poset and obtaining the associated graph as discussed in Section 5.4, then the editable graphs associated to $\pi_{0}\circ Y$ and $\pi_{0}\circ Y_{h}$ are going to be isomorphic as editable graphs, as their posets are going to be isomorphic as posets and $r\mapsto r+h$ is an isometry. While the persistent sets $\pi_{0}\circ Y$ and $\pi_{0}\circ Y_{h}$ are not isomorphic as functors. And this is true also for $Y:\mathbb{R}^{n}\rightarrow\operatorname*{Sets}$ and $h\in\mathbb{R}^{n}$.
As already mentioned, the fiber of the map sending a constructible persistent set $S$ into the editable graph $\mathcal{G}(D_{S_{\mid C}})$ (up to isomorphism) contains a lot of persistent sets which are not isomorphic. Moreover, we believe that, for any representation of persistent set/persistence module, it is always advisable to produce some stability result, to ensure the interpretability of the pipeline.
In the particular case of $\pi_{0}\circ Y:\mathbb{R}\rightarrow\operatorname*{Sets}$, both the issues presented have been taken care in Pegoraro (2024a). For all other scenarios these problems still await formal investigations.
Additionally, we would like also to extend this framework to Reeb graphs, similarly to what was done for merge trees. This is motivated by the extent to which Reeb graphs (Biasotti et al., 2008; De Silva et al., 2016) and their combinatorial approximation, namely mapper graphs (Singh et al., 2007) are studied (Carriere et al., 2018; Carriere and Oudot, 2018) and used in applications (Nielson et al., 2015; Rizvi et al., 2017). They add another layer of difficulties as a metric being insensitive to topological changes as the ones depicted in Figure 6 it is not desirable. This could be solved by modeling the possibility of having multiple edges connecting two vertices. It would also be important also to develop an algorithm to compute or approximate $d_{E}$ which works for Reeb graphs.
6 Trees
In the remaining part of the manuscript we specialize our discussion on a particular kind of directed graphs: trees. Thanks to 6 we have seen that this scenario still includes a lot of interesting situations, covering all constructible (one dimensional) persistent sets whose filtration which definitively is path-connected. Moreover, the peculiar structure of trees allows us to prove some key results which greatly decrease the complexity of computing optimal mappings between two trees, resulting in a bottom-up algorithm which is implemented and tested. We leave to future works the investigation of algorithms for more complicated kinds of graphs.
Remark 5
From now on our point of view is predominantly computational, thus we always assume to have finite graphs.
Definition 18
A tree structure $T$ is given by a connected rooted acyclic graph $(V_{T},E_{T})$. We indicate the root of the tree with $r_{T}$. The order of a vertex $v\in V_{T}$ is the number of edges which have that vertex as one of the extremes, and is called $ord_{T}(v)$. Any vertex with an edge connecting it to the root is its child and the root is its father: this is the first step of a recursion which defines the father and children relationship for all vertices in $V_{T}.$ The vertices with no children are called leaves and are collected in the set $L_{T}$. The relation $child<father$ generates a partial order on $V_{T}$ which induces an orientation on the edges. The edges in $E_{T}$ are identified in the form of ordered couples $(a,b)$ meaning $a<b$ and $a\rightarrow b$. A subtree of a vertex $v$, called $sub_{T}(v)$, is the tree structure whose set of vertices is $\{x\in V_{T}\mid x\leq v\}$.
To avoid overlapping some terminology with already existing notions of trees and for ease of notation, we refer to an editable tree by calling it dendrogram, as formalized by the following definition.
Definition 19
A tree structure $T$ with a weight function $\varphi_{T}:E_{T}\rightarrow X-\{0_{X}\}$, with $X$ editable space, is called dendrogram. The space of $X$-valued dendrograms is indicated as $(\mathcal{T},X)$.
A key fact is that given a tree structure $T$, identifying an edge $(v,v^{\prime})$ with its lower vertex $v$, gives a bijection between $V_{T}-\{r_{T}\}$ and $E_{T}$, that is $E_{T}\cong V_{T}-\{r_{T}\}$ as sets. Given this bijection, we often use $E_{T}$ to indicate the vertices $v\in V_{T}-\{r_{T}\}$, to simplify the notation. Moreover, since we now have a canonical way to associate edges to vertices, we can state the definition of mappings in a more directed and handy way, associating an edit operation to every edge/vertex in a tree - a part from the root. Such association is not available in general for graphs. In this way the ghosting of an edge can be described via the removal of individual vertices, and similarly deletions and shrinkings can be referred to as applied on vertices instead of edges. With an abuse of notation we call mappings the following objects, which from now on replace the mappings defined in Section 4.2.
The letters "$D"$ and "$G"$ are used to indicate deletion and ghosting. A mapping between $T$ and $T^{\prime}$ is thus a set $M\subset(E_{T}\cup\{"D","G"\})\times(E_{T^{\prime}}\cup\{"D","G"\})$ satisfying:
  • (M1)
    consider the projection of the Cartesian product $(E_{T}\cup\{"D","G"\})\times(E_{T^{\prime}}\cup\{"D","G"\})\rightarrow(E_{T}\cup\{"D","G"\})$; we can restrict this map to $M$ obtaining $\pi_{T}:M\rightarrow(E_{T}\cup\{"D","G"\})$. The maps $\pi_{T}$ and $\pi_{T^{\prime}}$ are surjective on $E_{T}\subset(E_{T}\cup\{"D","G"\})$ and $E_{T^{\prime}}\subset(E_{T^{\prime}}\cup\{"D","G"\})$ respectively;
  • (M2)
    $\pi_{T}$ and $\pi_{T^{\prime}}$ are injective;
  • (M3)
    $M\cap(V_{T}\times V_{T^{\prime}})$ is such that, given $(a,b)$ and $(c,d)\in M\cap(V_{T}\times V_{T^{\prime}})$, $a>c$, if and only if $b>d$;
  • (M4)
    if $(a,"G")\in M$ (or analogously ("$G",a)$), then after applying all deletions of the form $(v,"D")\in M$, the vertex $a$ becomes an order $2$ vertex. In other words: let $child(a)=\{b_{1},..,b_{n}\}$. Then there is one and only one $i$ such that for all $j\neq i$, for all $v\in V_{sub(b_{j})}$, we have $(v,"D")\in M$; and there is one and only one $c$ such that $c=\max\{x<b_{i}\mid(x,y)\in M$ for any $y\in V_{T^{\prime}}\}$.
We highlight some points with the following remarks.
Remark 6
This definition is equivalent to the one given in Section 4.2. In fact: (M1) and (M2) are equivalent to (G1) plus the fact that vertices not appearing in the image of $\eta_{G}^{M}:\mathcal{E}_{M}\rightarrow E_{T}$ are deleted. Property (G2) is equivalent to (M4): if a segment $s$ which is not simply an edge, and so it induces some ghostings - is such that all the edges in its open star are either deleted or belong to $Im(s)$, this means that all the vertices in $s^{o}$ are of order $2$ after the deletions. Lastly, (M3) implies that the trees considered are isomorphic as directed graphs once all deletions and ghostings are applied on the starting and target trees.
Remark 7
In literature usually mappings are subsets of $E_{T}\times E_{T^{\prime}}$ and their properties are equivalent to (M2) and (M3) (see Hong et al. (2017); Sridharamurthy et al. (2020); Wetzels et al. (2022) and references therein). In this way, mappings coherently maps edges into edges, respecting the tree ordering, inducing in this way the needed relabeling operations - see also 3. The differences with our definition are caused by the introduction of the ghosting edits. Accordingly, the way in which mappings parametrize edit paths in the present work is completely novel and, for instance, it establishes some (partial) ordering between the edits. Note that the set of mappings we consider is bigger than in usual edit distances - thus computing the distance has a higher computational cost.
6.1 Computing the Edit Distance for Trees: Decomposition Properties
In these last sections we develop an algorithm to compute the edit distance for dendrograms.
Following ideas found in Hong et al. (2017) and exploiting the properties of editable spaces, we now prove some theoretical results which allow us to recursively split up the problem of finding an optimal mapping between two dendrograms. Then, in Section 6.2, we formulate each of the split-up calculations as binary optimization problems, which are then aggregated recursively in a bottom-up fashion in Section 6.3·
Name $T_{2}$ the only representative without order $2$ vertices inside the equivalence class of $T$. One can always suppose that a dendrogram is given without order $2$ vertices. Thus, for notational convenience, from now on we suppose $T=T_{2}$ and $T^{\prime}=T^{\prime}_{2}$. To help us in the calculations define a particular subset of mappings $M_{2}(T,T^{\prime})\subset\operatorname*{Mapp}(T,T^{\prime})$.
Definition 20
A mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ has maximal ghostings if $(v,"G")\in M$ if and only if $v$ is of order $2$ after the deletions in $T$ and, similarly ("$G",w)\in M$ if and only if $w$ is of order $2$ after the deletions in $T^{\prime}$. A mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ has minimal deletions if $(v,"D")\in M$ only if neither $v$ nor $father(v)$ are of order $2$ after applying all the other deletions in $T$ and, similarly, ("$D",w)\in M$ only if $w$ is not of order $2$ after applying all the other deletions in $T^{\prime}$.
We collect all mappings with maximal ghostings in the set $M_{\mathcal{G}}(T,T^{\prime})$ and paths with minimal deletions in the set $M_{\mathcal{D}}(T,T^{\prime})$. Lastly: $M_{2}(T,T^{\prime}):=M_{\mathcal{G}}(T,T^{\prime})\bigcap M_{\mathcal{D}}(T,T^{\prime})$.
In other words with mappings in $M_{2}(T,T^{\prime})$ we are always eliminating all the order $2$ vertices which arise from deletions and we are not deleting edges which we can shrink. The following lemma then applies.
Lemma 1
$\min\{cost(M)\mid M\in\operatorname*{Mapp}(T,T^{\prime})\}=\min\{cost(M)\mid M\in M_{2}(T,T^{\prime})\}$
In addition to that, we consider some particular subsets of $E_{T}\times E_{T^{\prime}}$ which play a fundamental role in what follows. Recall that, using $E_{T}\cong V_{T}-\{r_{T}\}$, we can induce $\pi_{T}:E_{T}\times E_{T^{\prime}}\rightarrow V_{T}$.
A set $M^{*}\subset E_{T}\times E_{T^{\prime}}$ is in $\mathcal{C}^{*}(T,T^{\prime})$ if:
  • (A1)
    the points in $\pi_{T}(M^{*})$ form antichains in $V_{T}$ (and the same for $\pi_{T^{\prime}}(M^{*})$ in $V_{T^{\prime}}$), with respect to the partial order given by $father>son$. This means that any two distinct vertices of $T$ (respectively of $T^{\prime}$) which appear in $M^{*}$ are incomparable with respect to “$>$”;
  • (A2)
    the projections $\pi_{T}:M^{*}\rightarrow V_{T}$ and $\pi_{T^{\prime}}:M^{*}\rightarrow V_{T^{\prime}}$ are injective.
Consider now $M^{*}\in\mathcal{C}^{*}(T,T^{\prime})$. Starting from such set of couples we build a set of edits which form a “partial” mapping between $T$ and $T^{\prime}$: each couple $(x,y)\in M^{*}$ means that we do not care of what lies below $x\in V_{T}$ and $y\in V_{T^{\prime}}$ and we need to define edits only for the other vertices. The vertices below $x$ and $y$ will be taken care separately. See also Figure 7 for a visual example.
Loosely speaking, $M^{*}$ is used as a “dimensionality reduction tool”: instead of considering the problem of finding directly the optimal mapping between $T$ and $T^{\prime}$, we split up the problem in smaller subproblems, and put the pieces together using $M^{*}$. To formally do that, some other pieces of notation are needed.
Let $v\in E_{T}$. One can walk on the undirected graph of the tree-structure $T$ going towards any other vertex. For any $v\in E_{T}$, $\zeta_{v}$ is the shortest graph-path connecting $v$ to $r_{T}$. Note that this is the ordered set $\zeta_{v}=\{v^{\prime}\in V_{T}\mid v^{\prime}>v\}$. Similarly, denote with $\zeta_{x}^{x^{\prime}}$ the shortest path on the graph of $T$ connecting $x$ and $x^{\prime}$. Define the least common ancestor of a set of vertices $A$ as $LCA(A):=\min\{v\in V_{T}\mid v\geq A\}=\min\bigcap_{a\in A}\zeta_{a}$. In particular $\min\zeta_{x}\cap\zeta_{x^{\prime}}$ is the least common ancestor between $x$ and $x^{\prime}$: $LCA(x,x^{\prime})=\min\zeta_{x}\cap\zeta_{x^{\prime}}$.
By Property (A1), given $x\in V_{T}\cap\pi_{T}(M^{*})$, there exist a unique $\Omega_{M^{*}}(x)\notin\pi_{T}(M^{*})$ such that:
$\Omega_{M^{*}}(x)=\min\{LCA(x,x^{\prime})\mid x^{\prime}\in\pi_{T}(M^{*})\text{ and }x\neq x^{\prime}\}$
And the same holds for $y\in V_{T^{\prime}}\cap\pi_{T^{\prime}}(M^{*})$. For ease of notation we will often avoid explicit reference to $M^{*}$ and write directly $\Omega(x)$. Note that $\zeta_{x}^{\Omega(x)}$ is a segment (see 9).
With these bits of notation, given $M^{*}\in\mathcal{C}^{*}(T,T^{\prime})$, we build the “partial” mapping $\alpha(M^{*})$: is a mapping that ignores all the vertices which lie below $x\in V_{T}$ and $y\in V_{T}$ if $(x,y)\in M^{*}$. Figure 7 may help in following the upcoming paragraph. Consider $v\in V_{T}$:
  1. 1.
    if $(v,w)\in M^{*}$, then $(v,w)\in\alpha(M^{*})$;
  2. 2.
    if there is not $x\in V_{T}$ such that $v<\Omega(x)$ or $v>\Omega(x)$, then $(v,"D")\in\alpha(M^{*})$;
  3. 3.
    if there is $x\in V_{T}$ such that $v>\Omega(x)$ then $(v,"D")\in\alpha(M^{*})$;
  4. 4.
    if there is $x\in V_{T}$ such that $v<\Omega(x)$:
    1. (a)
      if $v\in\zeta_{x}^{\Omega(x)}$ then $(v,"G")\in\alpha(M^{*})$
    2. (b)
      if $v<v_{i}$ for some $v_{i}\in\zeta_{x}^{\Omega(x)}=\{v_{0}<v_{1}<\ldots<v_{n}\}$ then $(v,"D")\in\alpha(M^{*})$;
    3. (c)
      if $v<x$ no edit is associated to $v$.
Remark 8
By Properties (A1) and (A2), the conditions used to build $\alpha(M^{*})$ are mutually exclusive. This means that each $v\in V_{T}$ satisfies one and only one of the above conditions and so $\alpha(M^{*})$ is well defined.
The idea behind $\alpha(M^{*})$ is that, for all couples $(x,y)\in M^{*}$, we want to make the ghostings to turn the segments $\zeta_{x}^{\Omega(x)}$ and $\zeta_{y}^{\Omega(y)}$ respectively into the edges $(x,\Omega(x))$ and $(y,\Omega(y))$, and then shrink one in the other. As we already anticipated, $\alpha(M^{*})$ takes care of all the vertices in $V_{T}$ and $V_{T^{\prime}}$, a part from the sets $\cup_{(x,y)\in M^{*}}\{x^{\prime}\in E_{T}\mid x^{\prime}<x\}$ and $\cup_{(x,y)\in M^{*}}\{y^{\prime}\in E_{T^{\prime}}\mid y^{\prime}<y\}$. For this reason we say that $\alpha(M^{*})$ is a partial mapping.
We state this formally with the next proposition.
Proposition 7
Consider $T$ and $T^{\prime}$ and $M^{*}\in\mathcal{C}^{*}(T,T^{\prime})$. We obtain from such dendrograms, respectively, the dendrograms $\widetilde{T}$ and $\widetilde{T}^{\prime}$ by deleting all the vertices $\bigcup_{(x,y)\in M^{*}}\{x^{\prime}\in E_{T}\mid x^{\prime}<x\}$ and $\bigcup_{(x,y)\in M^{*}}\{y^{\prime}\in E_{T^{\prime}}\mid y^{\prime}<y\}$. The set $\alpha(M^{*})$ is a mapping in $M_{2}(\widetilde{T},\widetilde{T}^{\prime})$.
Now we have all the pieces we need to obtain the following key result.
Theorem 3 (Decomposition)
Given $T$, $T^{\prime}$ dendrograms:
$d_{E}(T,T^{\prime})=\min_{M^{*}\in\mathcal{C}^{*}(T,T^{\prime})}\sum_{(x,y)\in M^{*}}d_{E}(sub_{T}(x),sub_{T^{\prime}}(y))+cost(\alpha(M^{*})).$ (2)
Figure 7: Given two weighted trees $T$ (left) and $T^{\prime}$ (right) - which are the same of Figure 3(a) - we consider $T_{x}=sub_{T}(x)$ and $G_{y}=sub_{T^{\prime}}(y)$ and we use 3 to compute $d_{E}(T_{x},T_{y})$, as in the algorithm in Section 6.2. The set $M^{*}=\{(v,w),(v^{\prime},w^{\prime}),(v^{\prime\prime},w^{\prime\prime})\}$ satisfies (A1),(A2). The set $\alpha(M^{*})$ is made by the deletions/insertions indicated by the red edges, the ghostings/splittings indicated by the yellow and the shrinkings given by edges of the same color, different from red. To obtain $\widetilde{T}_{x}$ and $\widetilde{T}^{\prime}_{y}$ as in 7 all the black vertices covered by shaded regions must be deleted.
6.2 Computing the Edit Distance for Trees: Dynamical Binary Linear Programming problems
We want to use 3 to write a dynamical, binary linear optimization algorithm to calculate $d_{E}$: by translating 3 into a Binary Linear Programming (BLP) problem, we obtain a single step in a bottom-up procedure.
6.2.1 Notation
We are given two dendrograms $T,T^{\prime}$ inside some dendrogram space $(\mathcal{T},X)$. Our objective is to write down Equation 2 as a function of some binary variables.
Consider $x\in V_{T}$ and $y\in V_{T^{\prime}}$. Along with keeping the notation defined in Section 6.1, define $T_{x}:=sub_{T}(x)$ and $T_{y}:=sub_{T^{\prime}}(y)$, $N_{x}:=\dim(T_{x})=\#E_{T}$ and $N_{y}:=\dim(T_{y})=\#E_{T^{\prime}}$. Moreover, given $v\in V_{T_{x}}$, the sequence $v_{0}=v<v_{1}<\ldots<r_{T}$ indicates the points in the segment $\zeta_{v}$. Thus $v_{i}$ will be a vertex $v_{i}>v$. The same with $w\in V_{T_{y}}$
6.2.2 Relaxing the Optimization Problem
We would like to find $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ minimizing Equation (2) for $T_{x}$ and $T_{y}$, but this is a difficult task. In fact, as evident in the construction of $\alpha(M^{*})$, a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ has the role of pairing segments: if $(v,w)\in M^{*}$, then the segments $\zeta_{v}^{\Omega(v)}$ and $\zeta_{w}^{\Omega(w)}$ are shrunk one on the other by $\alpha(M^{*})$. However, the points $\Omega(v)$ and $\Omega(w)$ depend on the whole set $M^{*}$, and not simply on the couple $(v,w)$. Modelling such global dependence gives rise to non-linear relationships between coupled points, and so leading to a non linear cost function, in terms of points interactions, to be minimized. For this reason we “weaken” the last term in Equation (2), allowing also mappings different from $\alpha(M^{*})$ to be built from $M^{*}$. In other words we minimize over $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ the following equation:
$\sum_{(v,w)\in M^{*}}d_{E}(sub_{T_{x}}(v),sub_{T_{y}}(w))+cost(\beta(M^{*}))$ (3)
where $\beta(M^{*})$ is such that:
  • $\beta(M^{*})\in M_{\mathcal{G}}(\widetilde{T}_{x},\widetilde{T}_{y})$ (using the notation established in 7, replacing $T$ and $T^{\prime}$ with $T_{x}$ and $T_{y}$ respectively);
  • the set of vertices coupled by $\beta(M^{*})$ is exactly $M^{*}$: $M^{*}=\beta(M^{*})\bigcap V_{T_{x}}\times V_{T_{y}}$.
Since, by construction $M^{*}=\alpha(M^{*})\bigcap V_{T}\times V_{T^{\prime}}$ and by 7, $\alpha(M^{*})\in M_{2}(\widetilde{T}_{x},\widetilde{T}_{y})$, minimizing Equation (2) or Equation (3) gives the same results.
6.2.3 Setup and Variables
Suppose we already have $W_{xy}$ which is a $N_{x}\times N_{y}$ matrix such that $(W_{xy})_{v,w}=d_{E}(T_{v},T_{w})$ for all $v\in E_{T_{x}}$ and $w\in E_{T_{y}}$. Note that:
  • if $x$ and $y$ are leaves, $W_{xy}=0$.
  • if $v,w$ are vertices of $T_{x}$, $T_{y}$, then $W_{vw}$ is a submatrix of $W_{xy}$.
The function to be optimized is defined on the following set of binary variables: for every $v\in E_{T_{x}}$ and $w\in E_{T_{y}}$, for $v_{i}\in\zeta_{v}$, $v_{i}<r_{T_{x}}$, and $w_{j}\in\zeta_{w}$, $w_{j}<r_{T_{y}}$, take a binary variable $\delta^{v,w}_{i,j}$. We use $\delta$ to indicate the matrix of variables $(\delta^{v,w}_{i,j})_{v,w,i,j}$.
The mapping $\beta(M^{*})$ is built according to the variables $\delta^{v,w}_{i,j}$ with value $1$: we write a constrained optimization problem such that having $\delta^{v,w}_{i,j}=1$ means pairing the segments $\zeta_{v}^{v_{i+1}}$ (that is, the segment which starts with $(v,v_{1})$ and ends with $(v_{i},v_{i+1})$) and $\zeta_{w}^{w_{j+1}}$, and shrinking one in the other in the induced mapping.
In order to pair and shrink the segments $\zeta_{v}^{v_{i+1}}=\{v=v_{0},v_{1},\ldots,v_{i+1}\}$ and $\zeta_{w}^{w_{j+1}}$ we need to collect some edits in the set $\beta(M^{*})$ adding the following edits:
  • all the points $v_{k}\in\zeta_{v}^{v_{i+1}}$ with $10<k<i+$ 1 are ghosted, that is $(v_{k},G)\in\beta(M^{*})$;
  • if $v^{\prime}<v_{k}$ for some $10<k<i+$ 1, then $(v^{\prime},D)\in\beta(M^{*})$;
  • if $v^{\prime}\geq v_{i+1}$ and $v^{\prime}\neq r_{T}$, then $(v^{\prime},D)\in\beta(M^{*})$
  • $(v,w)\in\beta(M^{*})$.
In the end, all edges which are not assigned an edit operation, are deleted. Note that analogous edits must be induced on vertices in $T_{y}$. Thus, the edit $(v,w)\in\beta(M^{*})$, along the edit paths induced by $\beta(M^{*})$, means: shrinking the edge $(v,v_{i+1})$ onto $(w,w_{j+1})$). Recall that, if $\delta^{v,w}_{i,j}=1$, we do not need to define edits for $sub_{T_{x}}(v)$ and $sub_{T_{y}}(w)$ since, by assumption, we already know $d_{E}(T_{v},T_{w})$.
6.2.4 Constraints
Clearly, the set of $\delta^{v,w}_{i,j}$ with value $1$ does not always produce $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$: for instance segments could be paired multiple times. To avoid such issues, we build a set of constraints for the variable $\delta$.
For each $v^{\prime}\in V_{T_{x}}$ we call $\Phi(v^{\prime}):=\{(v^{\prime\prime},i)\in V_{T}\times\mathbb{N}\mid v^{\prime}=v^{\prime\prime}_{i}\in\zeta_{v^{\prime\prime}}^{v^{\prime\prime}_{i+1}}\}$. In an analogous way we define $\Phi(w^{\prime})$ for $w^{\prime}\in V_{T_{y}}$. Call $\mathcal{K}$ the set of values of $\delta$ such that for each leaf $l$ in $V_{T_{x}}$:
$\sum_{v^{\prime}\in\zeta_{l}}\left(\sum_{(v^{\prime\prime},i)\in\Phi(v^{\prime})}\sum_{w,j}\delta^{v^{\prime\prime},w}_{i,j}\right)\leq 1$ (4)
and for each leaf $l^{\prime}$ in $V_{T_{y}}$:
$\sum_{w^{\prime}\in\zeta_{l^{\prime}}}\left(\sum_{(w^{\prime\prime},j)\in\Phi(w^{\prime})}\sum_{v,i}\delta^{v,w^{\prime\prime}}_{i,j}\right)\leq 1$ (5)
The following proposition clarifies the properties of any value of $\delta\in\mathcal{K}$.
Proposition 8
If $\delta\in\mathcal{K}$:
  • the couples $(v,w)$ such that $\delta^{v,w}_{i,j}=1$ define a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$;
  • the edits induced by all $\delta^{v,w}_{i,j}=1$ give a mapping $\beta(M^{*})$ in $M_{``G^{\prime\prime}}(\widetilde{T}_{x},\widetilde{T}_{y})$. With $\widetilde{T}_{x},\widetilde{T}_{y}$ being obtained from $T_{x}$ and $T_{y}$ as in 7.
Remark 9
If for every $\delta^{v,w}_{i,j}=1$ we have $v_{i+1}=\Omega(v)$, then $\beta(M^{*})=\alpha(M^{*})$.
6.2.5 Objective Function
Having built a mapping $\beta(M^{*})$ using the binary variables, we want to define a cost functions which computes the cost of such mapping depending on $\delta$.
As before, consider $v\in E_{T_{x}}$ and interpret $\delta^{v,w}_{i,j}=1$ as coupling the segments $\zeta_{v}^{v_{i+1}}$ and $\zeta_{w}^{w_{j+1}}$; then $v$ is coupled with some $w\in E_{T_{y}}$ if $C(v):=\sum_{i,w,j}\delta^{v,w}_{i,j}=1$ and is ghosted if $G(v):=\sum_{\{i,v^{\prime}\mid v\in\zeta_{v^{\prime}}^{v^{\prime}_{i}}\}}\sum_{w,j}\delta^{v^{\prime},w}_{i,j}=1$. The vertex $v$ is instead deleted if $D(v):=1-C(v)-G(v)=1$. We introduce also the following quantities, which correspond to the cost of shrinking $\zeta_{v}^{v_{i+1}}$ on $\zeta_{w}^{w_{j+1}}$:
$\Delta^{v,w}_{i,j}=d_{X}\left(\bigodot_{v^{\prime}\in\zeta_{v}^{v_{i}}}\varphi_{T_{x}}(v^{\prime}),\bigodot_{w^{\prime}\in\zeta_{w}^{w_{j}}}\varphi_{T_{y}}(w^{\prime})\right)$
Note that the above operations are taken inside the editable space $X$.
The function which computes the cost given by coupled points is therefore:
$F^{C}(\delta):=\sum_{v,w,i,j}\Delta^{v,w}_{i,j}\cdot\delta^{v,w}_{i,j}$
The contribution of deleted points is: $F^{D}(\delta)-F^{-}(\delta)$, where
$F^{D}(\delta):=\sum_{v\in T_{x}}D(v)\cdot d(\varphi_{T_{x}}(v),0)+\sum_{w\in T_{y}}D(w)\cdot d(\varphi_{T_{y}}(w),0)$
and
$F^{-}(\delta):=\sum_{v\in T_{x}}C(v)\cdot\mid\mid sub_{T_{x}}(v)\mid\mid+\sum_{w\in T_{y}}C(w)\cdot\mid\mid sub_{T_{y}}(w)\mid\mid$
where the “norm” of a tree $T$ is $\mid\mid T\mid\mid=\sum_{e\in E_{T}}d_{X}(\varphi(e),0)$.
Finally, one must take into account the values of $d_{E}(T_{v},T_{w})$, whenever $v$ and $w$ are coupled; this information is contained in $(W_{xy})_{v,w}$:
$F^{S}(\delta):=\sum_{v,w}(W_{xy})_{v,w}\cdot\left(\sum_{i,j}\delta^{v,w}_{i,j}\right)$
Proposition 9
With the notation previously introduced:
$d_{E}(T_{x},T_{y})=\min_{\delta\in\mathcal{K}}F^{C}(\delta)+F^{D}(\delta)-F^{-}(\delta)+F^{S}(\delta)$ (6)
Proof  The contribution of coupled points is $F^{C}(\delta)$ and the contribution of deleted points is $F^{D}(\delta)-F^{-}(\delta)$.
The cost of $\beta(M^{*})$ is: $F^{\beta}(\delta):=F^{C}(\delta)+F^{D}(\delta)-F^{-}(\delta)$. Lastly, $F^{S}(\delta)$ takes into account the value of $d_{E}(T_{v},T_{w})$, if $v$ and $w$ are coupled. By 3, combined with 8, the solution of the following optimization problem:
$\min_{\delta\in\mathcal{K}}F^{S}(\delta)+F^{\beta}(\delta)$ (7)
is equal to $d_{E}(T_{x},T_{y})$.
Remark 10
A solution to Problem Equation 6 exists because the minimization domain is finite and there are admissible values; it is not unique in general.
6.3 Computing the Edit Distance for Trees: Bottom-Up Algorithm
In this section the results obtained in Section 6.1 and the formulation established in Section 6.2 are used to obtain the algorithm implemented to compute the metric $d_{E}$ between dendrograms. Some last pieces of notation are introduced in order to describe the “bottom-up” nature of the algorithm.
Given $x\in V_{T}$, define $\operatorname*{len}(x)$ to be the number of vertices in $\zeta_{x}$ and $\operatorname*{len}(T)=\max_{v\in V_{T}}\operatorname*{len}(v)$. Therefore, $\operatorname*{lvl}(x)=\operatorname*{len}(T)-\operatorname*{len}(x)$. Lastly, $\operatorname*{lvl}_{T}(n)=\{v\in V_{T}\mid\operatorname*{lvl}(v)=n\}$
The key property is that: $\operatorname*{lvl}(x)>\operatorname*{lvl}(v)$ for any $v\in sub(x)$. Thus, if $W_{xy}$ is known for any $x\in\operatorname*{lvl}_{T}(n)$ and $y\in\operatorname*{lvl}_{T^{\prime}}(m)$, then for any $v$, $w$ in $V_{T}$, $V_{T^{\prime}}$ such that $\operatorname*{lvl}(v)<n$ and $\operatorname*{lvl}(w)<m$, $W_{vw}$ is known as well. With this notation we can write down Algorithm 1.
Result: $d_{E}(T,T^{\prime})$
initialization: $N=\operatorname*{len}(T)$, $M=\operatorname*{len}(T^{\prime})$, $n=m=0$;
while $n\leq N$ or $m\leq M$ do
       for $(x,y)\in V_{T}\times V_{T^{\prime}}$ such that $\operatorname*{lvl}(x)\leq n$ and $\operatorname*{lvl}(y)\leq m$ do
             Calculate $(W_{r_{T}r_{T^{\prime}}})_{x,y}$ solving Problem (6);
       end for
      $n=n+1$; $m=m+1$;
end while
return $(W_{r_{T}r_{T^{\prime}}})_{r_{T},r_{T^{\prime}}}$
Algorithm 1 Bottom-Up Algorithm.
An hands-on example of some iterations of Algorithm 1 can be found in Section A.
6.4 Computational Cost
We end up with a result to analyze the performances of Algorithm 1 in the case of dendrograms with binary tree structures.
Proposition 10 (Computational Cost)
Let $T$ and $T^{\prime}$ be two dendrograms with full binary tree structures with $\dim(T)=\#E_{T}=N$ and $\dim(T^{\prime})=M$.
Then $d_{E}(T,T^{\prime})$ can be computed by solving $O(N\cdot M)$ BLP problems with $O(N\cdot\log(N)\cdot M\cdot\log(M))$ variables and $O(\log_{2}(M)+\log_{2}(N))$ constraints.
Proof  In a full binary tree structure, at each level $l$ we have $2^{l}$ vertices. Let $L=\operatorname*{len}(T)$ and $L^{\prime}=len(T^{\prime})$. We have that, for any vertex $v\in V_{T}$ at level $l$, the cardinality of the path from $v$ to any of the leaves in $sub_{T}(v)$ is ${L-l}$ and the number of leaves in $sub_{T}(v)$ is $2^{L-l}$.
So, given $v\in V_{T}$ at level $l$ and $w\in V_{T^{\prime}}$ at level $l^{\prime}$, to calculate $d_{E}(sub_{T}(v),sub_{T^{\prime}}(w))$ (having already $W_{vw}$) we need to solve a binary linear problem with $2^{L-l}\cdot(L-l)\cdot 2^{L^{\prime}-l^{\prime}}\cdot(L^{\prime}-l^{\prime})$ variables and $2^{L-l}+2^{L^{\prime}-l^{\prime}}$ linear constraints.
Thus, to calculate $d_{E}(T,T^{\prime})$, we need to solve $(2^{L+1}-1)\cdot(2^{L^{\prime}+1}-1)$ linear binary optimization problems, each with equal or less than $2^{L}\cdot L\cdot 2^{L^{\prime}}\cdot L^{\prime}$ variables and equal or less that $2^{L}+2^{L^{\prime}}$ constraints. Substituting $L=\log_{2}(N)$ and $L^{\prime}=\log_{2}(M)$ in these equations gives the result.
Note that binary dendrograms are dense (with respect to $d_{E}$) in any dendrogram space as long as for any $\varepsilon>0$, there is $x\in(X,\odot,0_{X})$ such that $d_{X}(x,0)<\varepsilon$.
Remark 11
We end the manuscript with a remark about the computational cost of $d_{E}$ as described in 10. We found this result surprising to some extent. If we consider the latest implementation of the classical edit distance between unlabelled and unordered trees obtained with BLP (Hong et al., 2017), then the complexity of the two metrics is very similar: the classical edit distance can be computed by solving $O(N\cdot M)$ BLP problems with $O(N\cdot M)$ variables and $O(\log_{2}(M)+\log_{2}(N))$ constraints - $O(N+M)$ if we count also the constraints restricting the integer variables to $\{0,1\}$, as the authors of Hong et al. (2017) do. Thus, the introduction of the ghosting and splitting edits, with all the complexities they carry, just increases the cost by two $\log$ factors in the number of variables. Moreover, the condition used in Zhang (1996) to make the classical edit distance tractable (that is, item (2) in Section 3.1. of Zhang (1996)) is compatible with (M1)-(M4) and thus a constrained version of $d_{E}$ is well defined. We believe that it could be computed with a polynomial time algorithm. We leave this investigation to future works.
7 Conclusions
We develop a novel framework to work with graphs that arise as topological summaries in the context of topological data analysis. These kind of summaries are increasingly frequent and require some particular care to be properly handled by a metric. We define an edit distance employing some novel edit operations, proving that such definition induces a metric up to a certain equivalence relationship between graphs.
For situations in which the graphs possess the structure of a tree then, by exploiting the ordering properties of trees, we produce a binary linear programming algorithm to compute such metric. The modifications that we make w.r.t. the classical edit distance between unlabeled trees make it only marginally more expensive, namely by some log factors. For this reason we argue that polynomial time approximations should be available also for $d_{E}$. Still, the computational procedure we describe here is used in separate applied works, as mentioned in the introduction. These additional modifications are essential for the stability properties assessed in Pegoraro (2024a), justifying the higher computational cost w.r.t. other edit distances for merge trees, which, instead, are unstable.
The generality of the work opens up many possible research directions. Some of which are already studied in other works, while others still await for investigations:
  • we think that the properties of the editable spaces can be relaxed; however, the algorithm presented in this manuscript may need to be adapted to the properties of the chosen weight space;
  • Reeb Graphs can be investigated theoretically and algorithmically, perhaps following the decomposition presented in Stefanou (2020);
  • likewise, computational procedures and theoretic investigations should be considered for particular finite sub-posets for $\mathbb{R}^{n}$;
  • as motivated by 11 we believe that polynomial time approximations can be obtained for computing $d_{E}$ between dendrograms;
  • further applications and case studies exploiting the generality of the weights we consider on the edges should be considered and explored.
Acknowledgments
This work was carried out as part of my PhD Thesis, under the supervision of Professor Piercesare Secchi.
Outline of the Appendix
Section A contains an hands-on example of some iterations of the algorithm presented in Section 6.3. Lastly, Section B contains the proofs of the results in the paper.
A Algorithmic Example
Here we present in details the first steps of the Algorithm 1, used to calculate the distance between two merge trees.
We consider the following couple of weighted trees. Let $(T,w_{T})$ be the tree given by: $V_{T}=\{a,b,c,d,r_{T}\}$, $E_{T}=\{(a,d),(b,d),(d,r_{T}),(c,r_{T})\}$ and $w_{T}(a)=w_{T}(b)=w_{T}(d)=1$, $w_{T}(c)=5$; the tree $(T^{\prime},w_{T^{\prime}})$ instead, is defined by: $V_{T^{\prime}}=\{a^{\prime},b^{\prime},c^{\prime},d^{\prime},r_{T^{\prime}}\}$, $E_{T^{\prime}}=\{(a^{\prime},d^{\prime}),(b^{\prime},d^{\prime}),(d^{\prime},r_{T^{\prime}}),(c^{\prime},r_{T^{\prime}})\}$ and $w_{T^{\prime}}(a)=1$, $w_{T^{\prime}}(b)=w_{T^{\prime}}(c)=2$ and $w_{T^{\prime}}(d)=3$. We employ the notation $w_{T}$ for the weight function instead of $\varphi_{T}$ to indicate the common situation of a tree with positive real weights.
Step: $n=m=0$
This step is trivial since we only have couples between leaves, like $(a,a^{\prime})$, which have trivial subtrees and thus $d_{E}(sub_{T}(a),sub_{T^{\prime}}(a^{\prime}))=0$.
Step: $n=m=1$
The points $x\in V_{T}$ with $\operatorname*{lvl}_{T}(x)\leq 1$ are $\{a,b,c,d\}$ and the points $y\in V_{T^{\prime}}$ with $\operatorname*{lvl}_{T^{\prime}}(y)\leq 1$ are $\{a^{\prime},b^{\prime},c^{\prime},d^{\prime}\}$. Thus the couples $(x,y)$ which are considered are: $(d,d^{\prime})$, $(d,a^{\prime})$, $(d,b^{\prime})$, $(d,c^{\prime})$ and $(a,d^{\prime})$, $(b,d^{\prime})$, $(c,d^{\prime})$. The couples between leaves, like $(a,a^{\prime})$ have already been considered.
Couple: $(d,d^{\prime})$
Let $T_{d}=sub_{T}(d)$ and $T_{d^{\prime}}=sub_{T^{\prime}}(d^{\prime})$. The set of internal vertices are respectively $E_{T_{d}}=\{a,b\}$ and $E_{T_{d^{\prime}}}=\{a^{\prime},b^{\prime}\}$. For each vertex $v<\text{root}$ in each subtree, where “root” stands for $d$ or $d^{\prime}$, roots of $T_{d}$ and $T_{d^{\prime}}$ respectively, we have $\zeta_{v}=\{v_{0}=v,v_{1}=\text{root}\}$. Thus, the binary variables we need to consider, are the following: $\delta_{0,0}^{a,a^{\prime}}$, $\delta_{0,0}^{a,b^{\prime}}$, $\delta_{0,0}^{b,a^{\prime}}$ and $\delta_{0,0}^{b,b^{\prime}}$. The quantities $\Delta_{i,j}^{v,w}$ are given by: $\Delta_{0,0}^{a,a^{\prime}}=0$, $\Delta_{0,0}^{a,b^{\prime}}=1$, $\Delta_{0,0}^{b,a^{\prime}}=0$ and $\Delta_{0,0}^{b,b^{\prime}}=1$. Thus:
$F^{C}(\delta)=0\cdot\delta_{0,0}^{a,a^{\prime}}+\delta_{0,0}^{a,b^{\prime}}+0\cdot\delta_{0,0}^{b,a^{\prime}}+\delta_{0,0}^{b,b^{\prime}}$
While:
$F^{D}(\delta)=(1-\delta_{0,0}^{a,a^{\prime}}-\delta_{0,0}^{a,b^{\prime}})\cdot 1+(1-\delta_{0,0}^{b,a^{\prime}}-\delta_{0,0}^{b,b^{\prime}})\cdot 1+(1-\delta_{0,0}^{a,a^{\prime}}-\delta_{0,0}^{b,a^{\prime}})\cdot 1+(1-\delta_{0,0}^{a,b^{\prime}}-\delta_{0,0}^{b,b^{\prime}})\cdot 2$
and:
$F^{-}(\delta)=(\delta_{0,0}^{a,a^{\prime}}+\delta_{0,0}^{a,b^{\prime}})\cdot 0+(\delta_{0,0}^{b,a^{\prime}}+\delta_{0,0}^{b,b^{\prime}})\cdot 0+(\delta_{0,0}^{a,a^{\prime}}+\delta_{0,0}^{b,a^{\prime}})\cdot 0+(\delta_{0,0}^{a,b^{\prime}}+\delta_{0,0}^{b,b^{\prime}})\cdot 0$
and:
$F^{S}(\delta)=\delta_{0,0}^{a,a^{\prime}}\cdot 0+\delta_{0,0}^{a,b^{\prime}}\cdot 0+\delta_{0,0}^{b,a^{\prime}}\cdot 0+\delta_{0,0}^{b,b^{\prime}}\cdot 0$
Lastly the constraints are:
$\delta_{0,0}^{a,a^{\prime}}+\delta_{0,0}^{a,b^{\prime}}\leq 1;\text{ }\delta_{0,0}^{b,a^{\prime}}+\delta_{0,0}^{b,b^{\prime}}\leq 1;\text{ }\delta_{0,0}^{a,a^{\prime}}+\delta_{0,0}^{b,a^{\prime}}\leq 1;\text{ }\delta_{0,0}^{a,b^{\prime}}+\delta_{0,0}^{b,b^{\prime}}\leq 1$
A solution is given by $\delta_{0,0}^{a,a^{\prime}}=\delta_{0,0}^{b,b^{\prime}}=1$ and $\delta_{0,0}^{a,b^{\prime}}=\delta_{0,0}^{b,a^{\prime}}=0$, which entails $F^{C}(\delta)=1$, $F^{D}(\delta)=0$, $F^{-}(\delta)=0$ and $F^{S}(\delta)=0$ and $d_{E}(T_{d},T_{d^{\prime}})=1$.
Couple: $(d,a^{\prime})$
Obviously: $d_{E}(sub_{T}(d),sub_{T^{\prime}}(a^{\prime}))=\mid\mid sub_{T}(d)\mid\mid$. All the couples featuring a leaf and an internal vertex (that is, a vertex which is not a leaf), such as $(d,b^{\prime})$, $(a,d^{\prime})$ etc. behave similarly.
Step: $n=m=2$
The points $x\in V_{T}$ with $\operatorname*{lvl}_{T}(x)\leq 2$ are $\{a,b,c,d,r_{T}\}$ and the points $y\in V_{T^{\prime}}$ with $\operatorname*{lvl}_{T^{\prime}}(y)\leq 2$ are $\{a^{\prime},b^{\prime},c^{\prime},d^{\prime},r_{T^{\prime}}\}$. Thus the couples $(x,y)$ which are considered are $(d,r_{T^{\prime}})$, $(r_{T},d^{\prime})$, $(r_{T},r_{T^{\prime}})$ and then the trivial ones: $(r_{T},a^{\prime})$, $(r_{T},b^{\prime})$, $(r_{T},c^{\prime})$ and $(a,r_{T^{\prime}})$, $(b,r_{T^{\prime}})$, $(c,r_{T^{\prime}})$. Some couples have already been considered and thus are not repeated.
Couple: $(d,r_{T^{\prime}})$
Let $T_{d}=sub_{T}(d)$ and $T^{\prime}=sub_{T^{\prime}}(r_{T^{\prime}})$. The set of internal vertices are respectively $E_{T_{d}}=\{a,b\}$ and $E_{T_{d^{\prime}}}=\{a^{\prime},b^{\prime},c^{\prime},d^{\prime}\}$. Thus, the binary variables we need to consider, are the following: $\delta_{0,0}^{a,a^{\prime}}$, $\delta_{0,1}^{a,a^{\prime}}$, $\delta_{0,0}^{a,b^{\prime}}$, $\delta_{0,1}^{a,b^{\prime}}$, $\delta_{0,0}^{a,c^{\prime}}$, $\delta_{0,0}^{a,d^{\prime}}$, $\delta_{0,0}^{b,a^{\prime}}$, $\delta_{0,1}^{b,a^{\prime}}$, $\delta_{0,0}^{b,b^{\prime}}$, $\delta_{0,1}^{b,b^{\prime}}$ $\delta_{0,0}^{b,c^{\prime}}$, and $\delta_{0,0}^{b,d^{\prime}}$.
The quantities $\Delta_{i,j}^{v,w}$ are given by: $\Delta_{0,0}^{a,a^{\prime}}=0$, $\Delta_{0,1}^{a,a^{\prime}}=3$, $\Delta_{0,0}^{a,b^{\prime}}=1$, $\Delta_{0,1}^{a,b^{\prime}}=4$, $\Delta_{0,0}^{a,c^{\prime}}=1$, $\Delta_{0,0}^{a,d^{\prime}}=2$, $\Delta_{0,0}^{b,a^{\prime}}=0$, $\Delta_{0,1}^{b,a^{\prime}}=3$, $\Delta_{0,0}^{b,b^{\prime}}=1$, $\Delta_{0,1}^{b,b^{\prime}}=4$, $\Delta_{0,0}^{b,c^{\prime}}=1$ and $\Delta_{0,0}^{b,d^{\prime}}=2$. The function $F^{C}(\delta)$ is easily obtained by summing over $\delta_{i,j}^{v,w}\cdot\Delta_{i,j}^{v,w}$.
While:
$F^{D}(\delta)=(1-\delta_{0,0}^{a,a^{\prime}}-\delta_{0,1}^{a,a^{\prime}}-\delta_{0,0}^{a,b^{\prime}}-\delta_{0,1}^{a,b^{\prime}}-\delta_{0,0}^{a,c^{\prime}}-\delta_{0,0}^{a,d^{\prime}})\cdot 1+\ldots+(1-\delta_{0,0}^{a,d^{\prime}}-\delta_{0,0}^{b,d^{\prime}})\cdot 3$
and:
$F^{-}(\delta)=(\delta_{0,0}^{a,a^{\prime}}+\delta_{0,1}^{a,a^{\prime}}+\delta_{0,0}^{a,b^{\prime}}+\delta_{0,1}^{a,b^{\prime}}+\delta_{0,0}^{a,c^{\prime}}+\delta_{0,0}^{a,d^{\prime}})\cdot 0+\ldots+(\delta_{0,0}^{a,d^{\prime}}+\delta_{0,0}^{b,d^{\prime}})\cdot 3$
and:
$F^{S}(\delta)=(\delta_{0,0}^{a,a^{\prime}}+\delta_{0,1}^{a,a^{\prime}})\cdot 0+(\delta_{0,0}^{a,b^{\prime}}+\delta_{0,1}^{a,b^{\prime}})\cdot 0+\ldots+\delta_{0,0}^{a,d^{\prime}}\cdot 3+\delta_{0,0}^{b,d^{\prime}}\cdot 3$
Lastly the constraints are:
$\delta_{0,0}^{a,a^{\prime}}+\delta_{0,1}^{a,a^{\prime}}+\delta_{0,0}^{a,b^{\prime}}+\delta_{0,1}^{a,b^{\prime}}+\delta_{0,0}^{a,c^{\prime}}+\delta_{0,0}^{a,d^{\prime}}\leq 1$
$\delta_{0,0}^{b,a^{\prime}}+\delta_{0,1}^{b,a^{\prime}}+\delta_{0,0}^{b,b^{\prime}}+\delta_{0,1}^{b,b^{\prime}}+\delta_{0,0}^{b,c^{\prime}}+\delta_{0,0}^{b,d^{\prime}}\leq 1$
$\delta_{0,0}^{a,a^{\prime}}+\delta_{0,1}^{a,a^{\prime}}+\delta_{0,0}^{b,a^{\prime}}+\delta_{0,1}^{b,a^{\prime}}+\delta_{0,0}^{a,d^{\prime}}+\delta_{0,0}^{b,d^{\prime}}\leq 1$
$\delta_{0,0}^{a,b^{\prime}}+\delta_{0,1}^{a,b^{\prime}}+\delta_{0,0}^{b,b^{\prime}}+\delta_{0,1}^{b,b^{\prime}}+\delta_{0,0}^{a,d^{\prime}}+\delta_{0,0}^{b,d^{\prime}}\leq 1$
$\delta_{0,0}^{a,c^{\prime}}+\delta_{0,0}^{b,c^{\prime}}\leq 1$
In this case there are many minimizing solutions. One is given by: $\delta_{0,1}^{a,a^{\prime}}=\delta_{0,0}^{b,c^{\prime}}=1$ and all other variables equal to 00. This value of $\delta$ is feasible since the variables $\delta_{0,1}^{a,a^{\prime}}$ and $\delta_{0,0}^{b,c^{\prime}}$ never appear in the same constraint. This value of $\delta$ entails $F^{C}(\delta)=3+1$, $F^{D}(\delta)=2$, $F^{-}(\delta)=0$ and $F^{S}(\delta)=0$, and thus $d_{E}(T_{d},T^{\prime})=6$.
Another solution can be obtained with: $\delta_{0,0}^{a,d^{\prime}}=\delta_{0,0}^{b,c^{\prime}}=1$ and all other variables equal to 00. Also this value of $\delta$ is feasible since the variables $\delta_{0,0}^{a,d^{\prime}}$ and $\delta_{0,0}^{b,c^{\prime}}$ never appear in the same constraint. This value of $\delta$ entails $F^{C}(\delta)=2+1$, $F^{D}(\delta)=w_{T^{\prime}}(a^{\prime})+w_{T^{\prime}}(b^{\prime})=1+2$, $F^{-}(\delta)=\mid\mid sub_{T^{\prime}}(d^{\prime})\mid\mid=3$ and $F^{S}(\delta)=d_{E}(sub_{T}(a),sub_{T^{\prime}}(d^{\prime}))=\mid\mid sub_{T^{\prime}}(d^{\prime})\mid\mid=3$, and thus $d_{E}(T_{d},T^{\prime})=3+3-3+3=6$.
Couple: $(r_{T},d^{\prime})$
This and the other couples are left to the reader.
B Proofs
Proof of 2.
We indicate with $\text{Lip}(\mathbb{R}^{2}_{\Delta})$ the set of functions $k:\mathbb{R}^{2}_{\Delta}\rightarrow\mathbb{R}$ which are Lipschitz on $\mathbb{R}^{2}_{\Delta}$ endued with the $1$-norm and such that $k(\Delta)=0$.
By Bubenik and Elchesen (2022) we have:
$W_{1}(D,D^{\prime})=W_{1}(\sum_{x\in D}\{x\},\sum_{y\in D^{\prime}}\{y\})=\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in D}k(x)-\sum_{y\in D^{\prime}}k(y)).$
At this point $(P3)$ and $(P4)$ follow easily by, respectively:
  • replacing $D^{\prime}$ with the empty diagram and noting:
    $\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in D}k(x))=\sum_{x\in D}\min_{y\in\Delta}\parallel x-y\parallel_{1};$
  • observing that when we have $W_{1}(D+D^{\prime\prime},D^{\prime}+D^{\prime\prime})$ the contributions by the elements in $D^{\prime\prime}$ when computing:
    $\sum_{x\in D+D^{\prime\prime}}k(x)-\sum_{y\in D^{\prime}+D^{\prime\prime}}k(y)$
    appear in both sums, erasing themselves out.
$\blacksquare$
Proof of 3.
We prove $s+s^{\prime}\in\mathcal{S}(G)$ and that it is well defined on the equivalence classes of $\mathcal{S}(G)$.
Because of the intersection condition then $s+s^{\prime}$ is injective. By construction it is also continuous. Moreover it is automatically an homeomorphism onto its image in a small enough neighbourhood of any point except for $1/2$. Let $(a,b)$ be an open neighbourhood of $1/2$ in $[0,1]$. We know that $s+s^{\prime}$ restricted to $(a,1/2]$ and to $[1/2,b)$ is in both cases an embedding. Which implies that $s+s^{\prime}(1/2)$ is going to be an internal point in $(s+s^{\prime})((a,b))$, which is then an open set. Thus $s+s^{\prime}$ is an homeomorphism also if restricted to $(a,b)$. Associativity follows easily.
In case $G$ is a directed graph we are done as considering $s^{\prime\prime}\sim s$ and $s^{\prime\prime\prime}\sim s^{\prime}$ would still give:
  • $s(1)=s^{\prime}(0)$;
  • $Im(s)\cap Im(s^{\prime})=s(1)$.
In case $G$ is an undirected graph, then $Im(s^{\prime\prime})\cap Im(s^{\prime\prime\prime})=s^{\prime\prime}(1)=s(1)$ holds in any case. However we may have $s^{\prime\prime}(1)\neq s^{\prime\prime\prime}(0)$. But, if that is the case, it is enough to change $s^{\prime\prime}(t)$ with $s^{\prime\prime}(1-t)$ and/or $s^{\prime\prime\prime}(t)$ with $s^{\prime\prime\prime}(1-t)$.
$\blacksquare$
Proof of 1.
To lighten the notation we use the following symbols:
  • the symbols $e_{d}$ and $e^{\prime-1}_{d}$ respectively mean that the edge $e$ is deleted and $e^{\prime}$ is inserted;
  • if instead we write $s_{g}$ and $s^{\prime-1}_{g}$ we imply that the segment $s$ is involved in a ghosting operation and the segment $s^{\prime}$ is the result of the splitting of an edge;
  • lastly, the shrinking induced by $(e,e^{\prime})$ is called $e_{w,w^{\prime}}$ with $w$ being the original weight function, and $w^{\prime}$ the weight function after the shrinking.
Note that all edits are defined on a single edge a part from ghostings and splittings, which in general are associated to a segment.
Suppose that $\gamma$ is a finite edit path. This means that $\gamma$ is the composition of a finite set of edits. We indicate such ordered composition with $\gamma=\prod_{i=0}^{N}e_{i}$ with $e_{i}$ edit operation. We would like to change the order of the edit operations without raising the cost and changing the extremes of the edit path. This is not always possible. However we can work it around in the useful cases using properties (P1)-(P4). In particular, we would like to know when we can commute a generic edit $e_{i}$ in the following situations:
  • $e_{d}\circ e_{i}$ and $e_{i}\circ e^{-1}_{d}$
  • $s_{g}\circ e_{i}$ and $e_{i}\circ e^{-1}_{g}$.
Moreover we want to reduce the edit path to max one shrinking/deletion/insertion for any edge of $G$ and $G^{\prime}$.
We divide the upcoming part of the proof in subsections, each devoted to different combinations of edits.
Inverse Operations
We point out the following fact which will be used in the proof. Suppose we have a couple of edits $e_{i}\circ e_{j}$ that can be replaced with the couple $e^{\prime}_{i}\circ e^{\prime}_{j}$. Now consider $e^{-1}_{j}\circ e^{-1}_{i}$:
$e^{-1}_{j}\circ e^{-1}_{i}=(e_{i}\circ e_{j})^{-1}=(e^{\prime}_{i}\circ e^{\prime}_{j})^{-1}=e^{\prime-1}_{j}\circ e^{\prime-1}_{i}.$ (8)
$e_{d}$ and $e_{d}^{-1}$
Note that, when we delete or insert one edge $e$, we are modifying the graph only at the level of its open star $\mathcal{N}(e)$. Consider then the following scenarios:
  • $e_{d}\circ s_{g}$ so that the ghostings induced by $s$ result in the edge $e$, which is then deleted. This composition of edits can be safely replaced with the deletions of all the edges in $s$, conserving the length of the path by (P3). If the ghostings defined by $s$ do not involve $e$ then we can safely commute the operations.
  • $e_{d}\circ s_{g}^{-1}$, with $e$ being a new edge originated by the splitting induced by $s_{g}^{-1}$. This composition of edits can clearly be replaced with a shrinking of the original edge and then a suitably defined splitting (which avoids the edge $e$) without changing the path length - by (P3) and (P4).
  • $e_{d}\circ e_{w,w^{\prime}}$ can be replaced by $e_{d}$ potentially diminishing the length of the path, but surely not raising it by (P1).
  • $e_{w,w^{\prime}}\circ e_{d}^{-1}$ can be replaced with an insertion directly with weight $w(e)$, possibly shortening the path (P1).
  • $s^{\prime}_{g}\circ e_{d}^{-1}$follows from Equation 8.
  • $s^{\prime-1}_{g}\circ e_{d}^{-1}$ follows from Equation 8.
$s_{g}$ and $s_{g}^{-1}$
Like in the previous case, we only take into account transformations regarding edges in the open star of the ghostings/splittings.
  • $s_{g}\circ s^{\prime}_{g}$, with $s$ and $s^{\prime}$ being segments whose interiors have non-empty intersection, can be merged in a unique ghosting by (P2). Note that, in case $G$ and $G^{\prime}$ are directed graphs, then the directions of the two segments are compatible. Otherwise the intersection would be empty. We avoid repeating this same comment in the upcoming points.
  • $s_{g}\circ s^{\prime-1}_{g}$, with $s$ and $s^{\prime}$ being segments whose interiors have non-empty intersection, can commute provided we define carefully the splittings $s^{\prime-1}_{g}$ by (P2), - avoiding the splittings which lie in the intersection.
  • $s_{g}\circ e_{w,w^{\prime}}$ with $e\subset Im(s)$. This means that we are shrinking an edge before merging it with others via some ghostings. However, we can achieve the same result, without increasing the path length, by making the ghostings at first, and then shrinking the resulting edge - by (P1)-(P4).
  • $e_{w,w^{\prime}}\circ s_{g}^{-1}$ follows from Equation 8.
$e_{w,w^{\prime}}$
  • $v_{w^{\prime},w^{\prime\prime}}\circ e_{w,w^{\prime}}$ can be replaced by $e_{w,w^{\prime\prime}}$ which is either conserving or shortening the path (P1).
  • note that $v_{w,w^{\prime}}^{-1}=v_{w^{\prime},w}$.
Thanks to these properties we can take a given path $\gamma=\prod_{i=0,...,N}e_{i}$ and modify the edit operations in order to obtain the following situation:
  • the first operations are all in the form $e_{d}$; this can be achieved because $e_{d}\circ-$ can be always rearranged so that the deletions are always done first, potentially shortening the path. Of course there can be only one deletion for each edge of $G$;
  • then we have all the edits in the form $s_{g}$; since $s_{g}\circ-$ is exchangeable when we do not have $s_{g}$ pre-composed by some deletion, this is not a problem. Observe that all ghostings (and so all splittings) involving segments whose interiors have non empty intersection can be joined in a unique ghosting operation;
  • in the same way we can put last all the paths in the form $e_{d}^{-1}$ and before them the ones in the form $s_{g}^{-1}$;
  • in the middle we are left with the shrinking paths, which by construction are on edges which were segments in the original graphs. Since we can substitute $e_{w,w^{\prime}}\circ v_{w^{\prime},w^{\prime\prime}}$ with $e_{w,w^{\prime\prime}}$, we can obtain just one single shrinking for each edge and this means that the shrinkings are pairing uniquely segments in the original graphs.
Thus the edit path
$\overline{\gamma}=(\gamma_{d}^{G^{\prime}})^{-1}\circ(\gamma_{g}^{G^{\prime}})^{-1}\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G},$
with:
  • $\gamma_{d}^{G}=\prod e_{d}$
  • $\gamma_{g}^{G}=\prod s_{g}$
  • $\gamma_{s}^{G}=\prod e_{w,w^{\prime}}$
  • $(\gamma_{g}^{G^{\prime}})^{-1}=\prod s_{g}^{-1}$
  • $(\gamma_{d}^{G^{\prime}})^{-1}=\prod e_{d}^{-1}$
is such that $\gamma(T)=\overline{\gamma}(G)=G^{\prime}$ and $cost(\overline{\gamma})\leq cost(\gamma)$. The key point is that $\overline{\gamma}$ can be easily realized as a mapping whose couples are given by the segments paired by the shrinkings.
$\blacksquare$
Proof of 6.
For the construction of the merge tree associated to a persistent set see Pegoraro (2024b).
Given a persistent set $S$ as in the hypotheses of the statement. To simplify the notation we make a slight abuse of notation and call $S$ also the restriction of $S$ to its critical set $C$.
Let $D_{S}$ be its display poset and $(T_{S},f_{S}:V\rightarrow\mathbb{R})$ with $T_{S}=(V,E)$ its merge tree. By construction $V-\{r_{T_{S}}\}$ with $r_{T_{S}}$ being the root of the merge tree and $f_{S}(r_{T_{S}})=+\infty$ can be sent into $D_{S}$: every vertex $v$ in $V-\{r_{T_{S}}\}$ is identified by a critical value $t_{i}=f_{S}(v)$ and an element $v\in S(t_{i})$. Thus we have $V-\{r_{T_{S}}\}\hookrightarrow D_{S}$. Moreover, this embedding is an embedding of posets if we consider $T_{S}$ with the relationship generated by $child<father$.
Consider now $(t_{j},p)\in D_{S}-V-\{r_{T_{S}}\}$: this happens if there are $t_{i}<t_{j}$ and $t_{k}>t_{j}$ such that for every $\varepsilon_{i}>0$ such that $t_{j}-\varepsilon_{i}>t_{i}$ and for every $\varepsilon_{k}>0$ such that $t_{j}+\varepsilon_{k}<t_{k}$, we have #$\#S(t_{j}\leq t_{j}+\varepsilon_{k})^{-1}(S(t_{j}\leq t_{j}+\varepsilon_{k})(p))=1$ and #$\#S(t_{j}-\varepsilon_{i}\leq t_{j})^{-1}(p)=1$. Moreover we can choose $t_{j}$ and $t_{k}$ so that:
  • $\max\{(t,s)\in D_{S}\mid(t,s)<(t_{j},p)\}=(t_{i},S(t_{i}\leq t_{j})^{-1}(p))$;
  • $\min\{(t,s)\in D_{S}\mid(t,s)>(t_{j},p)\}=(t_{k},S(t_{i}\leq t_{k})(p))$
This implies that for any edge $e$ connecting $(t,s)\rightarrow(t_{j},p)$ in the DAG representing $D_{S}$, we have that if $(t,s)\neq(t_{i},S(t_{i}\leq t_{j})^{-1}(p))$, then we also have an edge $(t,s)\rightarrow(t_{i},S(t_{i}\leq t_{j})^{-1}(p))$. But this implies that we can remove $e$ from the DAG without changing its transitive closure. That is, $e$ does not appear in $\mathcal{G}(D_{S})$. The same for any edge $(t_{j},p)\rightarrow(t,s)$ with $(t,s)\neq(t_{k},S(t_{i}\leq t_{k})(p))$. Thus $(t_{j},p)$ induces an order $2$ vertex in $\mathcal{G}(D_{S})$.
To conclude the proof we just need to prove that $\mathcal{G}(D_{S})$ is a tree. In fact, if this is the case, then we have a poset map $V-\{r_{T}\}\hookrightarrow V_{\mathcal{G}(D_{S})}$ so that all the vertices not in the image of this map are of order $2$. I.e. the map is an isomorphism of posets onto the image, which is the poset obtained by removing all order $2$ vertices from $\mathcal{G}(D_{S})$.
We know that $\mathcal{G}(D_{S})$ is a DAG which has, by construction one maximal element. The only thing we need to prove is that for every $x=(t_{i},p)\in V_{\mathcal{G}(D_{S})}$ we only have one directed edge of the form $e=(x,y)$. But this is easily seen: $y=S(t_{i}\leq t_{i+1})(p)$. By functoriality, for every other $z=(t,s)\geq z$, we also have $z\geq y$. Thus we are done.
$\blacksquare$
Proof of 1.
Any order $2$ vertex which is not ghosted is paired with another order $2$ vertex. Ghosting both of them does not increase the cost of the mapping.
$\blacksquare$
Proof of 7.
Condition (M2) coincide with condition (A2). Condition (M3) is clearly satisfied because of the antichain condition (A1). Consider a vertex $v\in E_{T}$. The only case in which $v$ is not edited is when $v<x$ with $x\in v_{T}\cap\pi_{T}(M^{*})$. However, in this case, $v$ does not appear in $\widetilde{T}$, and thus (M1) is satisfied. Moreover, all and only order $2$ vertices, after the deletions, are ghosted, and (M4) follows .
$\blacksquare$
Proof of 3.
Let $M\in M_{2}(T,T^{\prime})$ such that $d_{E}(T,T^{\prime})=cost(M)$.
We note that $father>son$ induces a partial order relationship also on the pairs given by coupled points in $M$: $(x,y)>(v,w)$ if $x>v$ and $y>w$. In fact, by property (M3), $x>v$ if and only if $y>w$. So we can select $(x_{i},y_{i})$, the maxima with respect to this partial order relationship. Thus, we obtain $(x_{0},y_{0})$,…,$(x_{n},y_{n})$ which form an antichain (both in $V_{T}$ and $V_{T^{\prime}}$).
Clearly $M^{*}=\{(x_{0},y_{0})$,…,$(x_{n},y_{n})\}\in\mathcal{C}^{*}(T,T^{\prime})$. Now we build $\alpha(M^{*})$ and compare the cost of its edits with the ones in $M$. Let $\bar{x}=LCA(x_{i},x_{j})$. Since $\bar{x}>x_{i},x_{j}$, it is not coupled in $M$. Since $x_{i}$ and $x_{j}$ are coupled, $\bar{x}$ cannot be ghosted, so it is deleted in $M$. Any point $x$ above $\bar{x}$ is deleted for the same reasons. So the edits above $\bar{x}$ are shared between $\alpha(M^{*})$ and $M$.
In $\alpha(M^{*})$ we ghost any point between $\bar{x}$ and $x_{i}$ (and the same for $x_{j}$) and this is not certain to happen in $M$ (some points could be deleted). Nevertheless, even in the worst case, these ghostings are guaranteed not to increase the distance. For instance, suppose $x_{i}<x<\bar{x}$ is deleted in $M$ and ghosted by $\alpha(M^{*})$, then:
$d(x_{i}\odot x,y_{i})\leq d(x_{i}\odot x,y_{i}\odot x)+d(y_{i}\odot x,y_{i})=d(x_{i},y_{i})+d(x,0)$
by properties (P1)-(P4). Since $\alpha(M^{*})\in M_{2}(\widetilde{T},\widetilde{T}^{\prime})$ by 7, we have:
$\sum_{(x,y)\in M^{*}}d_{E}(sub_{T}(x),sub_{T^{\prime}}(y))+cost(\alpha(M^{*}))\leq cost(M)$
Now we prove the other inequality.
Consider $M^{*}$ which realizes the minimum of the right side of Equation (2), and $M_{i}$ which realizes $d_{E}(sub(x_{i}),sub(y_{i}))$ with $(x_{i},y_{i})\in M^{*}$. We build a mapping $M$ collecting edits in the following way: for every $x^{\prime}\in E_{T}$ if $x^{\prime}\in sub(x_{i})$, we take the edit associated to it from $M_{i}$, otherwise we know that it is edited by $\alpha(M^{*})$, and we take it from there; the set of these assignments gives $M\in M_{2}(T,T^{\prime})$ whose cost is exactly $\sum_{(x_{i},y_{i})\in M^{*}}cost(M_{i})+cost(\alpha(M^{*}))$. This gives the second inequality.
$\blacksquare$
Proof of 8.
Having fixed a leaf $l$, the constraint in Equation (4) allows for at most one path $\zeta_{v}^{v_{i+1}}\subset\zeta_{l}$ to be kept after the edits induced by all the variables equal to $1$. Moreover if $(v^{\prime\prime},i)\in\Phi(v)\cap\Phi(v^{\prime})$, then $v=v^{\prime\prime}_{i}=v^{\prime}$. Thus, variables are added at most one time in Equation 4 and Equation 5. Which means that for any $a\in V_{T_{x}}$, we are forcing that the vertex $a$ can be an internal vertex or lower extreme of at most one path $\zeta_{v}^{v_{i+1}}$ such that $\delta_{i,j}^{v,w}=1$. In other words if two “kept” segments $\zeta_{v}^{v_{i+1}}$ and $\zeta_{v^{\prime}}^{v^{\prime}_{i^{\prime}+1}}$ (i.e. with $\delta_{i,j}^{v,w}=\delta_{i^{\prime},j}^{v^{\prime},w}=1$) intersect each other, it means that they just share the upper extreme $v_{i+1}=v^{\prime}_{i^{\prime}+1}$. These facts together imply that (if the constraints are satisfied) the edits induced on $T_{x}$ by $\delta^{v,w}_{i,j}=1$ and $\delta^{v^{\prime},w^{\prime}}_{i^{\prime},j^{\prime}}=1$ are always compatible: if $v^{\prime\prime}\in V_{sub_{T}(v_{i})}$ then it is not touched by (the edits induced by) $\delta^{v^{\prime},w^{\prime}}_{i^{\prime},j^{\prime}}=1$ (and the same exchanging the role of $v^{\prime}$ and $v$), if $v^{\prime\prime}$ is equal or above $v_{i+1}$ and/or $v^{\prime}_{i^{\prime}+1}$, then it is deleted in any case. Lastly, by noticing that if $\delta^{v,w}_{i,j}\in\Phi(v^{\prime})$ then $\delta^{v,w^{\prime}}_{i,j^{\prime}}\in\Phi(v^{\prime})$ for all other possible $w^{\prime}$ and $j^{\prime}$, we see that every segment $\zeta_{v}^{v_{i}}$ is paired with at most one segment $\zeta_{w}^{w_{j}}$, and viceversa.
As a consequence, for any vertex $v^{\prime}$ in any of the tree structures, at most one point on the segment $\zeta_{v^{\prime}}$ is coupled in $M^{*}$, guaranteeing the antichain condition. Moreover, any point of $T_{x}$ which is in $\pi_{T_{x}}(M^{*})$ is assigned to one and only one point of $T^{\prime}_{y}$ and viceversa. The edits induced by $\delta=1$ clearly satisfy properties (M2)-(M4). Passing to $\widetilde{T}_{x}$ and $\widetilde{T}_{y}$, also (M1) is satisfied.
$\blacksquare$

### v4

@@
-A Persistence-Driven Edit Distance for Graphs with Abstract Weights
+A Persistence-Driven Edit Distance for Trees with Abstract Weights
 Abstract
-In this work we define a novel edit distance for graphs considered with some weights on the edges.
-The metric is driven by the idea of considering graphs as topological summaries in the context of persistence and topological data analysis.
+In this work we define a novel edit distance for trees considered with some abstract weights on the edges.
+The metric is driven by the idea of considering trees as topological summaries in the context of persistence and topological data analysis.
 Several examples related to persistent sets are presented.
-In case the graphs are trees, the metric can be computed with a dynamical binary linear programming approach.
+The metric can be computed with a dynamical binary linear programming approach.
 This framework is applied and further studied in other works focused on merge trees, where the problems of stability and merge trees estimation are also assessed.
@@
 1 Introduction
-The use of graphs is almost ubiquitous in science: any time some system needs to be discretised or some relationships between points/statistical units/objects formalized or represented, graphs are always involved.
-For this reason a huge number of mathematical tools have been defined to work with graphs: differential calculus, distances between graphs, algebraic structures to solve graph-related problems and all other sorts of results have been developed throughout the years.
-In this paper we consider graphs assuming the point of view of topological data analysis (TDA) (Edelsbrunner and Harer, 2008), i.e.
-we consider them as representing topological information, and develop an edit distance to analyse populations of graph-shaped topological summaries.
-TDA is a quite recent type of data analysis which resorts to algebraic topology (Hatcher, 2000) to extract from data information regarding each statistical unit’s shape.
-A typical TDA pipeline starts from a topological space and then induces a family of nested subspaces, called filtration, from which some topological information is extracted via the use of some functors, for instance simplicial homology functors (Hatcher, 2000).
-Generalizations of this standard procedure include using family of subspaces indexed by partially ordered sets (posets), like $\mathbb{R}^{n}$ ($n\geq 2$) (Botnan and Lesnick, 2022) and the use of Laplacians (Mémoli et al., 2022), path-connected components (Curry et al., 2022, 2021; Pegoraro, 2024a; Cavinato et al., 2022), zig-zag homology (Carlsson and De Silva, 2010), cosheaves (De Silva et al., 2016), families of functors (Turner et al., 2014) etc..
-Our investigation starts from objects of the form $F:(P,\leq)\rightarrow\operatorname*{Sets}$ (with $(P,\leq)$ being a poset), which generalize persistent sets as defined in Carlsson and Mémoli (2013), and arise from very general situations like $(P,\leq)\rightarrow\operatorname*{Top}\xrightarrow{\pi_{0}}\operatorname*{Sets}$ (with $\operatorname*{Top}$ being the category of topological spaces and $\pi_{0}$ being the functor of path-connected components).
-For instance, if $(P,\leq)$ is given by the ordered set $\{f^{-1}((-\infty,t])\}_{t\in\mathbb{R}}$ for some $f:X\rightarrow\mathbb{R}$ we recover merge trees of real valued functions (Morozov et al., 2013; Curry et al., 2021; Pegoraro, 2024a), while if $(P,\leq)$ is the poset of open intervals in $\mathbb{R}$ we recover Reeb cosheaves (De Silva et al., 2016).
-Relationships between posets and graphs have been extensively studied: for instance it is well-known that any poset can be represented with a directed acyclic graph (DAG) (Aho et al., 1972).
+Graphs are widely used across scientific disciplines.
+Whenever a system is discretized, or when relations among points, statistical units, or abstract objects need to be formalized, graphs naturally emerge as a fundamental modeling tool.
+This ubiquity has led to the development of a broad array of mathematical techniques for analyzing and comparing graphs, ranging from differential calculus on discrete structures and graph metrics, to specialized algebraic frameworks.
+Over time, this has given rise to a rich and diverse landscape of graph-theoretic methods.
+Within this broader context, trees occupy a particularly important role.
+Their structural simplicity, combined with considerable expressive power, offers an appealing balance between interpretability and computational tractability.
+As a result, trees have become central in many areas of science.
+Two especially prominent examples are phylogenetic trees Billera et al.
+(2001); Felsenstein and Felenstein (2004); Garba et al.
+(2021); Lueg et al.
+(2024) and clustering dendrograms Murtagh and Contreras (2017); Xu and Tian (2015), which are closely related structures.
+Both describe how a set of elements evolves or merges under some notion of similarity or agglomeration, and both illustrate the versatility of trees as a means of encoding hierarchical information.
+In this paper, we approach trees from the perspective of topological data analysis (TDA) (Edelsbrunner and Harer, 2008), interpreting them as carriers of topological information.
+Our goal is to develop an edit distance tailored to analyzing populations of tree-shaped topological summaries.
+TDA is a relatively recent framework for data analysis that leverages tools from algebraic topology (Hatcher, 2000) to extract shape-related features from data at the level of individual statistical units.
+A typical TDA pipeline begins with a topological space and constructs a nested sequence of subspaces—known as a filtration—from which topological information is extracted using functors such as simplicial homology (Hatcher, 2000).
+This standard approach has been generalized in several directions, including filtrations indexed by more general partially ordered sets (e.g., $\mathbb{R}^{n}$ for $n\geq 2$) (Botnan and Lesnick, 2022), and the use of other mathematical objects such as Laplacians (Mémoli et al., 2022), path-connected components (Curry et al., 2022, 2021; Pegoraro, 2024a; Cavinato et al., 2022), zigzag homology (Carlsson and De Silva, 2010), cosheaves (De Silva et al., 2016) among others.
+Previous Works on Edit Distances for Trees in TDA
+The topological perspective we assume in this work stems from persistent sets, as defined in Carlsson and Mémoli (2013), which are functors of the form $F:(P,\leq)\rightarrow\operatorname*{Top}\xrightarrow{\pi_{0}}\operatorname*{Sets}$ (with $(P,\leq)$ being a poset, $\operatorname*{Top}$ being the category of topological spaces and $\pi_{0}$ being the functor of path-connected components).
+The most typical example being $(P,\leq)$ equal to $\{f^{-1}((-\infty,t])\}_{t\in\mathbb{R}}$ for some $f:X\rightarrow\mathbb{R}$, which amounts to using merge trees of real valued functions (Morozov et al., 2013; Curry et al., 2021), while, if $(P,\leq)$ is the poset of open intervals in $\mathbb{R}$, up to some gluing conditions, we recover Reeb cosheaves (De Silva et al., 2016).
+Relations between posets and graphs have been extensively studied: for instance it is well-known that any poset can be represented with a directed acyclic graph (DAG) (Aho et al., 1972).
 Thus, it is no surprise that a common way to represent instances of $F:(P,\leq)\rightarrow\operatorname*{Sets}$ is via some kind of graphs, like merge trees or Reeb graphs.
-Metrics to compare graphs are not a novel research topic (Billera et al., 2001; Gao et al., 2010; Bille, 2005; Feragen and Nye, 2020), but recent advances in stratified spaces, optimal transport and geometric data analysis have fostered some novel techniques to deal with such objects (Mémoli, 2011; Calissano et al., 2023).
-Similarly, the use of graphs (possibly labeled Yan et al.
+Accordingly, the use of graphs (possibly labeled Yan et al.
 (2019)) as summaries of persistent sets has produced a series of original definitions and results (Bauer et al., 2014; De Silva et al., 2016; Di Fabio and Landi, 2016; Bauer et al., 2020; Morozov et al., 2013; Gasparovic et al., 2019; Pont et al., 2022; Sridharamurthy et al., 2020; Wetzels et al., 2022; Sridharamurthy and Natarajan, 2021; Cardona et al., 2021; Touli, 2020; Curry et al., 2022) which are all driven by these novel scenarios.
-To differentiate this work from the previous ones, we quickly mention that the metric here presented, when used (up to some minor modifications) to compare merge trees, satisfies some stability properties (contrary to Pont et al.
+A significant number of these works focus on defining suitable metric structures for merge trees and Reeb graphs through variants of edit distances.
+In particular, (Di Fabio and Landi, 2016; Bauer et al., 2020) are focused on sup-norm kinds of distances, where the difference between (Reeb) graphs is determined by the cost of the biggest modification needed to turn one graph into another.
+While Pont et al.
 (2022); Sridharamurthy et al.
 (2020); Wetzels et al.
-(2022); Sridharamurthy and Natarajan (2021)), which are analogous to the ones of the $1$-Wasserstein distance between persistence diagrams (Edelsbrunner and Harer, 2022).
-In particular our metric adds all the local modifications needed to turn one graph into the other, while the other stable metrics look for the biggest modification that needs to be made, making those metrics more stable, but also less sensitive to differences on a global scale.
-Exactly as for the bottleneck distance and the Wasserstein distances between persistence diagrams.
-Formal details are given in Pegoraro (2024a).
-Lastly, differently from the other stable metrics, we provide, implement and employ an exact algorithm to compute the metric for trees.
+(2022); Sridharamurthy and Natarajan (2021) consider edit distances for merge trees which are more closely related to classical tree edit distances Tai (1979), where the objective is to minimize the total cost of a sequence of modifications converting one merge tree into another.
+In this work, we propose an edit distance for weighted trees that aligns more closely with the second family of approaches mentioned above, but it differs from previous methods in two key ways:
+  • Stability Properties: in Pegoraro (2024a) and Pegoraro and Secchi (2024) it is shown that, when adapted to merge trees, the metric we define here produces an edit distance $d_{E}$ satisfying the following stability relation :
+    $d_{I}(T,T^{\prime})\leq d_{E}(T,T^{\prime})\leq 2(\text{size}(T)+\text{size}(T^{\prime}))d_{I}(T,T^{\prime})$
+    where $d_{I}$ is interleaving distance between merge trees (Morozov et al., 2013) and $\text{size}(T)$ is the number of edges in $T$.
+Which is analogous to the relation between the bottleneck distance ($d_{B}$) and the $1$-Wasserstein metric ($W_{1}$) between persistence diagrams (PDs):
+    $d_{B}(D,D^{\prime})\leq W_{1}(D,D^{\prime})\leq(\#D+\#D^{\prime})d_{B}(D,D^{\prime}).$
+    See Edelsbrunner and Harer (2010).
+Among the edit distances presented in Pont et al.
+(2022); Sridharamurthy et al.
+(2020); Wetzels et al.
+(2022); Sridharamurthy and Natarajan (2021), only the one in Wetzels et al.
+(2022) was shown—in Pegoraro (2024a)—to posses some stability properties.
+This stability result also clarifies the theoretical distinction between our approach and those in Bauer et al.
+(2014); Di Fabio and Landi (2016); Bauer et al.
+(2020), all of which yield distances that are bounded from above by $d_{I}$.
+Since $d_{B}$ is the universal metric between PDs and $d_{I}$ is the universal one between merge trees, by analogy, one can further characterize the difference between $d_{E}$ and Di Fabio and Landi (2016); Bauer et al.
+(2020), via the easier and well understood difference between $d_{B}$ and $W_{1}$.
+  • Abstract Weights: we investigate generalizations in which trees are enriched not only with edge lengths, but with weights taking values in more general metric spaces.
+This allows topological summaries to capture and represent a wider variety of information.
 Main Contributions
-In this paper we define a novel edit distance apt to compare graphs which are interpreted as topological summaries.
-Since, increasingly often, such objects are enriched with some kind of additional information which is, to some extent, attached to vertices and edges of the graphs (Ushizima et al., 2012; Curry et al., 2022; Pegoraro, 2024b; Curry et al., 2023a), we keep our dissertation as general as possible, allowing for some abstract weights to be associated to the edges in the graph.
-In particular, we can articulate the contributions of this paper in two distinct points:
+As anticipated, the main contribution of this work is the introduction of a novel edit distance designed to compare trees interpreted as topological summaries.
+Increasingly, such trees are enriched with additional information, typically associated with the vertices or edges of the graph, as seen in various recent works (Ushizima et al., 2012; Curry et al., 2022; Pegoraro, 2024b; Curry et al., 2023a).
+To remain as general as possible, we allow abstract weights to be assigned to the edges of the trees.
+Given the wide range of settings that emerge from this enriched perspective (as discussed in Section 5 and reflected in the diverse works built on the framework we establish) this paper focuses on the following core contributions:
   1. 1.
-    the main contributions are given by the definition of a distance between weighted graphs, with weights lying in certain metric spaces, and a computational framework to exactly compute the distance with binary linear optimization when the graph is a tree;
+    Defining a general edit distance for trees with abstract weights;
   2. 2.
-    we start exploring the possibilities offered by the different weights we can attach to the edges, discussing the potential applications for these abstract weights and how can they be used to tackle different data analysis scenarios (some of which have already been investigated in other works).
-The edit distance we define stems from classical edit distances (Levenshtein et al., 1966; Tai, 1979; Bille, 2005; Gao et al., 2010) but add some deep modifications, which can be intuitively motivated by the following example, represented in Figure 1.
+    Formalizing the abstract properties that edge weights must satisfy for the metric to be well-defined;
+  3. 3.
+    Developing a computational framework for efficiently computing the metric;
+  4. 4.
+    Beginning to explore the variety of edge weights that can be used, discussing their potential applications and how can they support diverse data analysis scenarios.
+We leave a deeper investigation into the interpretability of the metric in specific scenarios and the study of its stability properties to other works.
+The edit distance we define stems from classical edit distances (Levenshtein et al., 1966; Tai, 1979; Bille, 2005; Gao et al., 2010; Lerouge et al., 2016), but add some deep modifications, which can be intuitively motivated by the following example, represented in Figure 1.
 (a) A function (orange) with its associated merge tree.
@@
 Suppose that we were to remove from the (merge) tree the edge $(a,b)$, we would obtain the red (merge) tree in Figure 1(c).
-Vertex $a$ is now an order $2$ vertex which represents a merging between path connected components (i.e.
+Vertex $a$ is now a degree $2$ vertex which represents a merging between path connected components (i.e.
 a local maximum in the function), which is no more happening.
@@
 Interconnected Works
-As already pointed out, the framework we introduce here lies the foundations for a series of developments which are carried out in separate works.
-Some of these developments, in turn, contain results which retrospectively validate the approach we pursue in this manuscript.
-In particular, we have the following group of interconnected works:
-  • the present work, where the main theoretic and computational framework is established, with the aim of providing a tool to work with graphs with general weights;
-  • when the weights on the edges are just positive numbers, this general framework is used by Pegoraro (2024a) to define a metric for merge trees, which can be computed with the algorithm we present here.
+The framework introduced in this work lays the foundation for a series of developments explored in separate studies: the edit distance we define is applied across different contexts, each accompanied by dedicated theoretical results that both justify the specific application and reinforce the foundational framework developed in this manuscript.
+We summarize such relations as follows:
+  • All the works we mention in the following points rely heavily on Section 4.2 and 2 for developing theoretic results, and on Algorithm 1 for applications and case studies.
+  • Merge trees: this general framework is used by Pegoraro (2024a) to define a metric for merge trees, which can be computed with Algorithm 1.
 With the help of some results contained in Pegoraro and Secchi (2024), which is more focused on the statistical aspects of such metric, Pegoraro (2024a) shows that the approach we pursue here induces a metric with stability properties which are analogous to the ones of the $1$-Wasserstein distance between persistence diagrams.
-In Pegoraro and Secchi (2024) the metric is also tested (using the algorithm provided in this work) on a classification pipeline for a data set which is benchmark in the field of functional data analysis; ;
-  • Pegoraro (2024b), instead, exploits the present work to explore the case with weights being functions defined on the edges of the trees.
-The stability result obtained in Pegoraro (2024b) stems from the one proved in Pegoraro and Secchi (2024);
-  • finally, Cavinato et al.
-(2022) considers hierarchical clustering dendrograms and carries out an analysis of a radiomic data set, putting to work the framework (theoretic and computational) we define here on real data.
-Therein, a stability result in terms of Gromov-Hausdorff distance between point clouds is also obtained (which relies again on the one in Pegoraro and Secchi (2024)).
+Pegoraro and Secchi (2024) and Cavinato et al.
+(2022) present real data applications to show the practical effectiveness of the metric.
+  • Functions defined on merge trees: Pegoraro (2024b) exploits the present work, exploring the case where weights are functions defined on the edges of the trees, obtaining another stability result.
 Outline
 Section 2 introduces the idea of editable spaces, which are the spaces we want the weights of our graphs to live in.
-Their properties are exploited in Section 3 to define editable graphs.
-Section 4 develops the novel edit distance and proves that it is well defined for graphs considered up to order $2$ vertices.
-Section 5 presents a series of general examples connecting persistent sets, and their generalizations, to editable graphs.
-Section 6 specialises the previously introduced tools to the case where the graph is a tree and obtains an algorithm to compute the distance via linear binary optimization.
-Section 7 concludes the paper by highlighting pros and cons of the approach along with possible further developments.
+Their properties are exploited in Section 3 to define dendrograms.
+Section 4 develops the novel edit distance and proves that it is well defined for trees considered up to degree $2$ vertices.
+Section 5 presents a series of general examples regarding persistent sets.
+Section 6 obtains an algorithm to compute the distance via linear binary optimization.
+Section 7 concludes the paper by highlighting the pros and cons of the approach along with possible further developments, with a brief comparison with graph edit distances.
 2 Editable Spaces
-As already stated, the aim of the manuscript is to present a novel metric structure to work with weighted (and possibly directed) graphs, where the word weighted is intended in a broad sense: weights are given by a map associating the edges of a graph to elements in a metric space.
-In order to define the metric, we need these spaces of weights to satisfy some properties which are fundamental to ensure some desired behaviours when modifying graphs.
+As outlined in the introduction, the goal of this manuscript is to introduce a novel metric structure for working with weighted trees, where the term weighted is interpreted broadly: edge weights are defined via a map that assigns each edge a value in a metric space.
+To define this metric meaningfully, we require the metric spaces to satisfy certain properties.
+These conditions are essential to guarantee well-behaved transformations when modifying trees.
 Definition 1
@@
   • (P2)
-    $(X,\odot,0_{X})$ is a monoid (that is $X$ has an associative operation $\odot$ with zero element $0_{X}$)
+    $(X,\odot,0_{X})$ is a commutative monoid (that is $X$ has an associative operation $\odot$ with zero element $0_{X}$)
   • (P3)
@@
 Moreover (P3)-(P4) imply that the points $0_{X}$, $x$, $y$ and $x\odot y$ form a rectangle which can be isometrically embedded in an Euclidean plane with the Manhattan geometry (that is, with the norm $\mid\mid\cdot\mid\mid_{1}$): $d_{X}(x,x\odot y)=d_{X}(0_{X},y)$, $d_{X}(y,x\odot y)=d_{X}(0_{X},x)$ and $d_{X}(x\odot y,0_{X})=d_{X}(0_{X},x)+d(0_{X},y)$.
-Exploiting the monoidal structure of an editable space we can define a partial order which we will use in later sections.
+Exploiting the monoidal structure of an editable space, we can turn it into a partially ordered set.
 Proposition 1
@@
   • Antisymmetry: suppose $x\leq y$ and $x\geq y$. Then $x\odot r=y$ and $y\odot q=x$ for some $r,q\in X$. But then:
-    $d_{X}(y,0)=d_{X}(x\odot r,0)=d_{X}(x,0)+d_{X}(r,0)=d_{X}(y\odot q,0)+d_{X}(r,0)=d_{X}(y,0)+d_{X}(q,0)+d_{X}(r,0),$
+    $d_{X}(y,0)=$ $d_{X}(x\odot r,0)=d_{X}(x,0)+d_{X}(r,0)$
+    $=$ $d_{X}(y\odot q,0)+d_{X}(r,0)=d_{X}(y,0)+d_{X}(q,0)+d_{X}(r,0),$
     which implies $r=q=0$ and so $x=y$.
   • Transitivity follows from associativity of $\odot$.
+If the editable space $X$ is contained in a group $V$, the partial order can be further characterized.
 Corollary 1
@@
 $\mathbb{R}^{2}_{\Delta}=\{(b,d)\in R^{2}\mid 0\leq b\leq d\}.$
-Similarly $D(\Delta)$ is the set of formal sums of elements in $\Delta=\{(b,b)\in R^{2}\mid 0\leq b\}$.
+Similarly $D(\Delta)$ is the set of formal sums of elements in $\Delta=\{(b,b)\in\mathbb{R}^{2}\mid 0\leq b\}$.
 In both sets of formal sums we have an operation given by the sum of formal sums, which is commutative and has the empty formal sum as zero element.
-Given $p,q\in D(\mathbb{R}^{2}_{\Delta})$ we say that $s\sim q$ if there exist $r,t\in D(\Delta)$ such that $p+r=q+t$.
-We collect the equivalence classes of such relationships in $D(\mathbb{R}^{2}_{\Delta},\Delta)$ and call them persistence diagrams.
+Given $D_{1},D_{2}\in D(\mathbb{R}^{2}_{\Delta})$ we say that $D_{1}\sim D_{2}$ if there exist $D_{3},D_{4}\in D(\Delta)$ such that $D_{1}+D_{3}=D_{2}+D_{4}$.
+We collect the equivalence classes of such relations in $D(\mathbb{R}^{2}_{\Delta},\Delta)$ and call them persistence diagrams, and indicate such an equivalence class with the notation $[D]$.
 The sum operation and the zero element are coherently defined on the quotient.
-As a consequence of 2 we can write every diagram $D$ as a sum of points in $\mathbb{R}^{2}_{\Delta}$ each with coefficient $1$.
+As a consequence of 2 we can write every diagram $[D]$ as a sum of points in $\mathbb{R}^{2}_{\Delta}$ each with coefficient $1$.
 This “decomposition” is clearly non-unique as it varies between elements in the same equivalence class.
-With an abuse of notation we may say that $x\in D$, meaning that $x=(b,d)$ is an off-diagonal point of $\mathbb{R}^{2}_{\Delta}$ (i.e.
+With an abuse of notation we may say that $x\in[D]$, meaning that $x=(b,d)$ is an off-diagonal point of $\mathbb{R}^{2}_{\Delta}$ (i.e.
 $d>b$) appearing in such decomposition of an element (in fact, in any element) in the equivalence class of $D$.
@@
 The $1$-Wasserstein metric between persistence diagrams is:
-$W_{1}(D,D^{\prime})=\inf_{\gamma:D\rightarrow D^{\prime}}\parallel x-\gamma(x)\parallel_{1},$
-where $\gamma$ is any bijection between an element in the equivalence class of $D$ and an element in the equivalence class $D^{\prime}$.
-The infimum in the definition of $W_{1}(D,D^{\prime})$ is always attained by some $\gamma$, and thus, it is always a minimum.
+$W_{1}([D],[D^{\prime}])=\inf_{\gamma:\widetilde{D}\rightarrow\widetilde{D}^{\prime}}\parallel x-\gamma(x)\parallel_{1},$
+where $\gamma$ is any bijection between some $\widetilde{D}$ such that $\widetilde{D}\sim D$, and some $\widetilde{D}^{\prime}$ such that $\widetilde{D}^{\prime}\sim D^{\prime}$.
+The infimum in the definition of $W_{1}([D],[D^{\prime}])$ is always attained by some $\gamma$, and thus, it is always a minimum.
 Proposition 2
 The space $(D(\mathbb{R}^{2}_{\Delta},\Delta),W_{1},+,0)$ is editable.
+Remark 1
+We assumed 3 as it makes 2 easier to prove, compared to other formulations.
+However, for the sake of simplicity, we drop the notation $[D]$ in the remaining of the manuscript referring to a diagram simply as $D$, chosing the representative with no points of the form $(b,b)$.
 2.1.3 Functions in Editable Spaces
-Consider an editable space $(X,d_{X},\odot_{X},0_{X})$ and a measure space $(M,\mu)$.
-Then, the space of functions $L_{1}(M,\mu,X):=\{f:M\rightarrow X\mid\int_{M}d_{X}(0_{X},f(p))d\mu(p)<\infty\}/\sim$ with $\sim$ being the equivalence relationship of functions up to zero measure subsets and $\odot_{W}$ being defined pointwise - is editable.
+Consider an editable space $(X,d_{X},\odot_{X},0_{X})$ and a measure space $(M,\mu)$. Then, the space of functions:
+$L_{1}(M,\mu,X):=\{f:M\rightarrow X\mid\int_{M}d_{X}(0_{X},f(p))d\mu(p)<\infty\}/\sim,$
+with $\sim$ being the equivalence relation of functions up to zero measure subsets and $\odot_{W}$ being defined pointwise - is editable.
 In fact, the function $p\mapsto d_{X}(0,f(p))$ is always non negative, so if properties (P3) and (P4) hold for a fixed $p\in M$, then they hold also for integrals.
@@
 For instance, $\mathbb{R}^{n}_{\geq 0}$ with the $1$-norm is an editable space.
-3 Editable Graphs
-Now we introduce the objects we will work with. That is, graphs, possibly directed, with a function associating to each edge a weight in a fixed editable space.
+3 Dendrograms
+Now we introduce the objects we will work with: trees, with a function associating to each edge a weight in a fixed editable space.
 Definition 4
-A directed graph is made by a couple $G=(V_{G},E_{G})$ with $V$ being the set of its vertices and $E_{G}\subseteq V\times V$ the set of its edges.
-An undirected graph instead is a couple $G=(V_{G},E_{G})$ with the set of edges being a subset of $V\times V/\operatorname*{Sym}(2)$ where the quotient is over the action of $\operatorname*{Sym}(n)$, the symmetric group over $n$ elements.
-Note that any directed graph $(V,E)$ can be turned into an undirected graph by considering its the image of the following map: $E\hookrightarrow V\times V\rightarrow V\times V/\operatorname*{Sym}(2)$.
-We use the generic word graph when we don’t want to distinguish between directed and undirected.
-This means implied that the word “graphs” can be replaced with “directed graphs” or “directed graphs” and any edge can be considered either as $e\in V\times V/\operatorname*{Sym}(2)$ or as $e\in V\times V$.
-We will point out if there is something noteworthy distinguishing the directed and undirected scenarios.
-The upcoming definitions should clarify how the reader should interpret this simplification process.
+A tree structure $T$ is given by a connected rooted acyclic graph $(V_{T},E_{T})$.
+We indicate the root of the tree with $r_{T}$.
+The degree of a vertex $v\in V_{T}$ is the number of edges which have that vertex as one of the endpoints.
+Any vertex with an edge connecting it to the root is its child and the root is its parent: this is the first step of a recursion which defines the parent and children relation for all vertices in $V_{T}.$ The vertices with no children are called leaves and are collected in the set $L_{T}$.
+The relation $child<parent$ generates a partial order on $V_{T}$ which induces an orientation on the edges.
+The edges in $E_{T}$ are identified in the form of ordered pairs $(a,b)$ meaning $a<b$ and $a\rightarrow b$.
+A subtree of a vertex $v$, called $sub_{T}(v)$, is the tree structure whose set of vertices is $\{x\in V_{T}\mid x\leq v\}$.
+A key fact is that given a tree structure $T$, identifying an edge $(v,v^{\prime})$ with its lower vertex $v$, gives a bijection between $V_{T}-\{r_{T}\}$ and $E_{T}$, that is $E_{T}\cong V_{T}-\{r_{T}\}$ as sets.
+Given this bijection, we often use $E_{T}$ to indicate the vertices $v\in V_{T}-\{r_{T}\}$, to simplify the notation.
+To avoid overlapping some terminology with already existing notions of trees and for ease of notation, we refer to a weighted tree, with the weight function taking values in an editable space, by calling it dendrogram, as formalized by the following definition.
 Definition 5
-An editable (undirected/directed) graph with weights in $X$ is a (undirected/directed) graph $G=(V,E)$ such that: there is a weight function defined on the edges $\varphi_{G}:E\rightarrow X$ with $X$ being an editable space and $0\notin Im(\varphi_{G})$.
-Two editable (undirected/directed) graphs $G=(V_{G},E_{G}),G^{\prime}=(V_{G^{\prime}},E_{G^{\prime}})$ are isomorphic if there is a bijection between the vertices $\eta:V_{G}\rightarrow V_{G^{\prime}}$ giving an isomorphism of (undirected/directed) graphs such that $\varphi_{G}(e)=\varphi_{G^{\prime}}(\eta(e))$, with $\eta((v_{1},v_{2})):=(\eta(v_{1}),\eta(v_{2}))$.
-Example 1
-Positively weighted undirected graphs and positively weighted directed graphs $G=(V,E)$ induce editable graphs with values in $X=\mathbb{R}_{\geq 0}$.
+A tree structure $T$ with a weight function $\varphi_{T}:E_{T}\rightarrow X-\{0_{X}\}$, with $X$ editable space, is called dendrogram.
+The space of $X$-valued dendrograms is indicated as $(\mathcal{T},X)$.
+3.1 Finite Posets, DAGs and Transitive Reductions
+In this section, we justify our focus on tree-shaped objects by outlining, at a high level, the various contexts in which such graphs arise in TDA.
+This also yields a conceptual blueprint for potential extensions of the present work.
+We start from the definition of a directed graph.
 Definition 6
-An editable graphs space $(\mathcal{G},X)$ is the set of editable graphs with weight values in $X$ up to editable graphs isomorphisms.
-Definition 7
-An editable graph is finite if $V_{G}$ and $E_{G}$ are finite sets.
-Given a graph $G=(V,E)$ we indicate with $\mathcal{CW}(G)$ the CW-complex built by considering as many 00-cells as vertices in $V$ and a $1$-cell for every edge in $E$ with the gluing function induced by the extremes of the edge.
-See Figure 2(a) and Figure 2(b).
-Note that $\mathcal{CW}(G)$ is well defined (up to homeomorphism) also for undirected graphs: exchanging the extremes of the gluing maps gives an homeomorphic $\mathcal{CW}(G)$.
-We have a canonical map $V\hookrightarrow\mathcal{CW}(G)$.
-With an abuse of notation we will often write $v\in\mathcal{CW}(G)$ meaning the image of $v\in V$ via $V\hookrightarrow\mathcal{CW}(G)$.
-Similarly there is a map $E\hookrightarrow\mathbf{C}(\mathcal{CW}(G))$ (with $\mathbf{C}(U)$ being the set of closed sets of a topological space $U$), induced by the image of the $1$-cells.
-We need to establish the following notation.
-Definition 8 (Schelp (1972))
-A partial semigroup (PSG) is a set $A$ with a partial binary operation: $+:A^{\prime}\rightarrow A$, with $A^{\prime}\subset A\times A$, which is associative, i.e.:
-  1. 1.
-    $(x,y+z)\in A^{\prime}$ if and only if $(x+y,z)\in A^{\prime}$
-  2. 2.
-    if $(x,y+z)\in A^{\prime}$ then $(x,y+z)=(x,y+z)$.
-We want to work with particular paths defined on the $CW$-complexes $\mathcal{CW}(G)$.
-To avoid creating ambiguous notation related to the word “paths” and in analogy with the literature on real trees (Evans, 2008), we give the following definition.
-Definition 9
-Consider a graph $G=(V,E)$.
-A segment in $G$ is an embedding $s:[0,1]\rightarrow\mathcal{CW}(G)$ such that $s(0),s(1)\in V\hookrightarrow\mathcal{CW}(G)$ and there is a path on $G$, $\gamma_{s}=\{(v_{0},v_{1}),(v_{1},v_{2}),\ldots,(v_{n-1},v_{n})\}\subset E$ such that: $v_{0}=s(0)$, $v_{n}=s(1)$ and $Im(\gamma_{s}\hookrightarrow\mathbf{C}(\mathcal{CW}(G)))=Im(s)$.
-For two segments $s,s^{\prime}$ we have $s\sim s^{\prime}$ if $Im(s)=Im(s^{\prime})$.
-The internal points of a segment are defined as $s^{o}:=Im(s)-\{s(0),s(1)\}$.
-We collect all equivalence classes of segments in $G$ in the set $\mathcal{S}(G)$.
-Remark 1
-We have a map $E\hookrightarrow\mathcal{S}(G)$ induced by the attachment maps of the $1$-cells of the $CW$-complex: $e\mapsto s_{e}\in\mathcal{S}(G)$ so that $s_{e}([0,1])=e$.
-Note that in case $G$ is an undirected graph, $s(t)\sim s(1-t)$.
-Proposition 3
-$\mathcal{S}(G)$ can be given a PSG structure via paths composition. Consider $s,s^{\prime}\in\mathcal{S}(G)$ such that:
-  • $s(1)=s^{\prime}(0)$;
-  • $Im(s)\cap Im(s^{\prime})=s(1)$;
-then $s+s^{\prime}:=s^{\prime\prime}:[0,1]\rightarrow\mathcal{CW}(G)$ with $s^{\prime\prime}(t)=s(2t)$ for $t\in[0,1/2]$ and $s^{\prime\prime}(t)=s^{\prime}(2t-1)$ for $t\in[1/2,1]$.
-(a) A directed graph.
-(b) The CW-complex associated to the graph in Figure 2(a).
-(c) The sum of segments by paths composition.
-(d) Two different segments $s,s^{\prime}$ with their open stars $\mathcal{N}(s^{o})$ and $\mathcal{N}(s^{\prime o})$. We have $\mathcal{N}(s^{\prime o})=s^{\prime o}$.
-Figure 2: Plots related to Section 3.
-See also Figure 2(c).
-Note that, as a consequence of 3, the set $\mathcal{S}(G)$ is almost a directed structure on the space $\mathcal{CW}(G)$ (as defined in Fajstrup et al.
-(2016)), missing just the constant paths.
-We state a couple of observations in the following remark.
-Remark 2
-$\mathcal{S}(G)$ is generated by $s_{e}\in\mathcal{S}(G)$: every $s\in\mathcal{S}(G)$ can be uniquely written as $\sum_{e\in Im(s)}s_{e}$.
-Thus, the weight function of any editable (directed) graph $\varphi_{G}$ can be extended to a PSG map $\varphi_{G}:\mathcal{S}(G)\rightarrow X$.
-Lastly we define the open star of a subset of the $CW$-complex associated to a graph.
-This definition is standard for simplicial complexes and we extend it roughly by looking at the open star of the graph projected onto the $CW$-complex.
-Definition 10
-Given $U\subset\mathcal{CW}(G)$, the open star of $U$, $\mathcal{N}(U)\subset\mathcal{CW}(G)$, is the image under $V\hookrightarrow\mathcal{CW}(G)$ of all the vertices $v\in U$, united with $s_{e}^{o}$ for all the edges $e=(v_{1},v_{2})$ such that $v_{1}\in U$ or $v_{2}\in U$.
-Particularly important segments are the ones whose open star in $\mathcal{CW}(G)$ is equal to the (open) segment itself. See also Figure 2(d).
-Definition 11
-We define $\mathcal{O}(G)\subset\mathcal{S}(G)$ as the set of all segments $s\in\mathcal{S}(G)$ such that $\mathcal{N}(s^{o})=s^{o}$.
-4 Edits of Graphs
-Now we present the main contribution of the paper.
-The approach we follow is to define a distance inspired by Edit Distances (Tai, 1979), but with substantial differences in the edit operations.
-The philosophy of these distances is to allow certain modifications of the base object, called edits, each being associated to a cost, and to define the distance between two objects as the minimal cost that is needed to transform the first object into the second with a finite sequence of edits.
-In this way, up to properly setting up a set of edits, one can formalize the deformation of a graph modifying the local information induced by the weight function defined on the edges and the graph structure itself.
-On top of that, edit distances frequently enjoy some decomposition properties which simplify the calculations (Hong et al., 2017), which are notoriously very heavy (Hein et al., 1995).
-We also point out that, in literature, graphs are often considered with weights being real numbers and avoid modelling the weight space as we do with the definition of editable spaces.
-Given an editable (directed) graphs space $(\mathcal{G},X)$, with $(X,\odot,0_{X})$ editable space, we can define our edits.
-  • We call shrinking of an edge a change of the weight function.
-The new weight function must be equal to the previous one on all edges, apart from the “shrunk” one.
-In other words, for an edge $e$, this means changing the value $\varphi(e)$ with another non zero value in $X$.
-  • A deletion is an edit with which an edge $e$ is removed from the graph.
-If $e=(v_{1},v_{2})$ is the only edge between $v_{1}$ and $v_{2}$, then we also remove either $v_{1}$ or $v_{2}$.
-The result is well defined up to (directed) graph isomorphism.
-The inverse of the deletion is the insertion of an edge along with, potentially, the addition of a novel vertex.
-We can insert an edge between two existing vertices or at a vertex $v$ specifying the name of the new neighbour of $v$, the neighbours of the newly added vertex (that can be either none, or any portion of the neighbours of $v$).
-Clearly we always need to define the value of the weight function on the new edge.
-  • Lastly, consider $s=\sum s_{e}\in\mathcal{O}(G)$.
-We can remove all edges and vertices in $s^{o}$ and add to $E_{G}$ the edge $(s[0],s[1])$ with, eventually, the direction coherently induced.
-The weight of the new edge is already encoded in $w:\mathcal{S}(G)\rightarrow X$ and we have $w(s)=\bigodot_{e\in Im(s)}w(e)$.
-This transformation is called the ghosting of the vertices in $Im(s)$.
-Its inverse transformation is called the splitting of an edge $e=(v_{1},v_{2})$ and replaces $e$ with some novel vertices $v^{\prime}_{1},\ldots,v^{\prime}_{n}$ and edges $(v_{1},v^{\prime}_{1}),(v^{\prime}_{1},v^{\prime}_{2})\ldots,(v^{\prime}_{n},v_{2})$ which form a segment in the new graph $G^{\prime}$ such that $s\in\mathcal{O}(G^{\prime})$ with $s(0)=v_{1}$, $s(1)=v_{2}$.
-Note that any $s_{e}\in\mathcal{O}(G)$ induces a trivial ghosting.
-(a) Starting and target weighted graphs.
-(b) Deletions in red.
-(c) Ghostings in yellow.
-(d) Shrinkings in green.
-(e) Splittings in yellow.
-(f) Insertions in red.
-Figure 3: (b)$\rightarrow$(e) form an edit path made from the left graph in Figure 3(a) to the right one.
-In each figure the edges involved in the editing are highlighted with different colors.
-In the following plot such vertices return black.
-This edit path can clearly be represented with a mapping - Section 4.2 - which first deletes all the red vertices in Figure 3(b), then ghosts all the yellow segments in Figure 3(c), then shrinks the green edges in Figure 3(d), split edges by means of the yellow vertices in Figure 3(c) and inserts all the red edges in Figure 3(f).
-An editable graph $G$ can be edited to obtain another graph, on which one can apply a new edit to obtain a third graph and so on.
-One can think of this as composing two edits $e_{0}$, $e_{1}$ which are not defined on the same graph, since the second edit is defined on the already edited graph.
-This is what we mean by composition of edits.
-Any finite composition of edits is referred to as an edit path.
-The notations we use are functional notations, even if the edits are not operators, since an edit is not defined on the whole space of graph but on a single graph.
-For example $e_{1}\circ e_{0}(G)$ means that $G$ is edited with $e_{0}$, and then $e_{0}(G)$ with $e_{1}$.
-See Figure 3 for an example of an edit path between two graphs whose weights are edges lenghts - Section 2.1.1.
-Remark 3
-Even if shrinking and deletions are classical edits, with shrinking being usually referred to as relabeling, the ghosting edit is completely unusual and we are not aware of any previous work employing it.
-Its definition is largely driven by the TDA perspective presented in the Introduction and it is fundamental for the stability results contained in Pegoraro and Secchi (2024) and Pegoraro (2024a) - dealing with the case of $G$ being a tree.
-We also give the following definition.
-Definition 12
-Graphs are equal up to order $2$ vertices if they become isomorphic after applying a finite number of ghostings or splittings. We write $(G,\varphi_{G})\cong_{2}(G^{\prime},\varphi_{G^{\prime}})$.
-12 induces an equivalence relationship which identifies the set of graphs inside $(\mathcal{G},X)$ that we want to treat as equal.
-We call $(\mathcal{G}_{2},X)$ the space of equivalence classes of graphs in $(\mathcal{G},X)$, equal up to order $2$ vertices.
-4.1 Costs of Edit Operations
-Now we associate to every edit a cost, that is, a length measure in the space $(\mathcal{G},X)$. The costs of the edit operations are defined as follows:
-  • if, via shrinking, an edge goes from weight $x$ to weight $y$, then the cost of such operation is $d_{X}(x,y)$;
-  • for any deletion/insertion of an edge with weight $x$, the cost is equal to $d_{X}(x,0_{X})$;
-  • the cost of ghosting/splitting operations is always zero.
-This is coherent with the properties of editable spaces: consider a ghosting merging two edges separated by an order two vertex, with $x$ being the weight of one edge, and $y$ the cost of the other.
-The total weight of the two edges is the same before and after the ghosting: $\mid d_{X}(x\odot y,0_{X})-d_{X}(x,0_{X})-d_{X}(y,0_{X})\mid=0$.
-Definition 13
-Given two graphs $G$ and $G^{\prime}$ in $(\mathcal{G},X)$, define:
-  • $\Gamma(G,G^{\prime})$ as the set of all finite edit paths between $G$ and $G^{\prime}$;
-  • $cost(\gamma)$ as the sum of the costs of the edits for any $\gamma\in\Gamma(G,G^{\prime})$;
-  • the graph edit distance as:
-    $d_{E}(G,G^{\prime})=\inf_{\gamma\in\Gamma(G,G^{\prime})}cost(\gamma)$
-By definition the triangle inequality and symmetry must hold, but, up to now, this edit distance is intractable; one would have to search for all the possible finite edit paths which connect two graphs in order to find the minimal ones.
-On top of that, having an edit which is completely “for free”, it is not even obvious that $d_{E}(G,G^{\prime})>0$ for some graphs.
-However, it is clear that $d_{E}$ induces a pseudo-metric on classes of graphs up to order two vertices.
-4.2 Mappings
-A mapping between the editable graphs $G=(V_{G},E_{G})$ and $G^{\prime}=(V_{G^{\prime}},E_{G^{\prime}})$ is a set $M\subset\mathcal{S}(G)\times\mathcal{S}(G^{\prime})$ satisfying some constraints.
-Before listing such properties, we introduce the following notation: (1) given two sets $A,B$, $A-B:=\{a\in A\mid a\notin B\}$ (2) for any $M\subset\mathcal{S}(G)\times\mathcal{S}(G^{\prime})$:
-$\mathcal{E}_{M}:=\coprod_{e\subset Im(s),s\in\pi_{G}(M)}e$
-with $\pi_{G}:\mathcal{S}(G)\times\mathcal{S}(G^{\prime})\rightarrow\mathcal{S}(G)$ being the projection of the cartesian product.
-We then have the induced maps $\eta_{G}^{M}:\mathcal{E}_{M}\rightarrow E_{G}$.
-Lastly, $M$ is a mapping if:
-  • (G1)
-    the maps $\eta_{G}^{M}$ and $\eta_{G^{\prime}}^{M}$ are injective;
-  • (G2)
-    if $s\in\pi_{G}(M)$ such that $s\in\mathcal{S}(G)-E_{G}$, then $\mathcal{N}(s^{o})-s^{o}\subset G-\pi_{G}(M)$;
-  • (G3)
-    if we delete all edges in $E_{G}-Im(\eta_{G}^{M})$ and do all the ghostings induced by $s\in\pi_{G}(M)-E_{G}$ (and analogously for $G^{\prime}$), then $M$ induces a graph isomorphism between the resulting subgraphs of $G$ and $G^{\prime}$.
-Let’s try to unveil the meaning of (G1)-(G3).
-The main idea is that any couple in a mapping $M$ identifies a segment of $G$ and a segment of $G^{\prime}$ which are paired and turned one in the other by the mapping.
-Condition (G1) is asking that every edge in $G$ and $G^{\prime}$ is at most within one segment which is coupled with a segment of the other graph.
-We refer to such edges as coupled edges.
-The edges which are not coupled will be deleted.
-Condition (G2) ensures that once we remove all edges which are not coupled, and so which are deleted, then we are in the position to make all the ghostings to turn all the segments $s\in\pi_{G}(M)-E_{G}$ (and $s^{\prime}\in\pi_{G^{\prime}}(M)-E_{G^{\prime}}$), in edges of a graph.
-Lastly, (G3) forces the couples in $M$ to induce a graph isomorphism between the resulting graphs.
-Remark 4
-Properties (G1)-(G3) are not canonical properties of mappings, and descend from the fact that we have introduced the ghosting operation, which is a novelty of this work.
-To better compare properties (G1)-(G3) with more classical edit distances we refer to Section 6, where we report (G1)-(G3) for the special case in which the graph is indeed a tree.
-Mappings in fact are often used to take advantage of the properties of trees to reduce the computational cost of matching two trees.
-In what follows we exploit (G1)-(G3) to parametrize a set of edit paths in the editable graphs space starting from $G$ and ending in $G^{\prime}$, which are collected under the name $\gamma_{M}$.
-We call:
-  • $\gamma_{d}^{G}$ a path made by the deletions to be done in $G$, that is, the edges $E_{G}-Im(\eta_{G}^{M})$, executed in any order.
-So we obtain the graph $G^{M}_{d}=\gamma_{d}^{G}(G)$, which, instead, is well defined and not depending on the order of the deletions.
-  • One then proceeds with ghosting all the segments $s\in\pi_{G}(M)-E_{G}$, in any order, getting a path $\gamma^{G}_{g}$ and the graph $G_{M}:=\gamma^{G}_{g}\circ\gamma^{G}_{d}(G)$.
-  • By property (G3), $M$ induces an isomorphism between the two graphs $G^{\prime}_{M}$ (defined in the same way as $G_{M}$, but starting from $G^{\prime}$) and $G_{M}$.
-So one can shrink the edges of $G_{M}$ onto the edges of $G^{\prime}_{M}$ as described by $M$.
-The composition of all the shrinkings, executed in any order is an edit path $\gamma_{s}^{G}$.
-By construction:
-$\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G}(G)=G^{\prime}_{M},$
-and:
-$(\gamma_{d}^{G^{\prime}})^{-1}\circ(\gamma_{g}^{G^{\prime}})^{-1}\circ\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G}(G)=G^{\prime}.$
-where the inverse of an edit path is thought as the composition of the inverses of the single edit operations, taken in the inverse order.
-Lastly, we call $\gamma_{M}$ the set of all possible edit paths:
-$(\gamma_{d}^{G^{\prime}})^{-1}\circ(\gamma_{g}^{G^{\prime}})^{-1}\circ\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G},$
-obtained by changing the order in which the edit operations are executed inside $\gamma_{d}$, $\gamma_{g}$ and $\gamma_{s}$.
-Observe that, even if $\gamma_{M}$ is a set of paths, its cost is well defined:
-$cost(M):=cost(\gamma_{M})=cost(\gamma_{d}^{G})+cost(\gamma_{s}^{G})+cost(\gamma_{d}^{G^{\prime}}).$
-See Figure 3 for an example of a mapping between weighted graphs.
-[…]
-Theorem 1 (Main Theorem)
-Given two editable graphs $G$ and $G^{\prime}$, for every finite edit path $\gamma$, there is a mapping $M\in\operatorname*{Mapp}(G,G^{\prime})$ such that $cost(M)\leq cost(\gamma)$.
-A first corollary immediately follows.
-Corollary 2
-For any two finite editable graphs $G$ and $G^{\prime}$, the set $\operatorname*{Mapp}(G,G^{\prime})$ is finite. Thus we have the following well defined pseudo-metric between finite editable graphs:
-$d_{E}(G,G^{\prime})=\inf\{cost(\gamma)\mid\gamma\in\Gamma(G,G^{\prime})\}=\min\{cost(M)\mid M\in\operatorname*{Mapp}(G,G^{\prime})\}$
-which we will refer to as the edit distance between $G$ and $G^{\prime}$.
-A second corollary is obtained observing that, if a mapping has cost equal to zero, then it must contain only ghostings.
-Corollary 3
-Given $G$ and $G^{\prime}$ editable graphs, $d_{E}(G,G^{\prime})=0$ if and only if $G$ and $G^{\prime}$ are equal up to order $2$ vertices.
-In other words $d_{E}$ is a metric for graphs considered up to order $2$ vertices.
-5 A Discussion on the Use of Graphs and Different Editable Spaces
-In this section we report a discussion featuring some key examples which regard editable graphs with values in different editable spaces induced from persistent sets (Carlsson and Mémoli, 2013; Curry, 2018; Curry et al., 2022).
-Persistent sets are objects of great interest in TDA and represent a (non-equivalent) alternative to persistence modules/diagrams when working with information collected from path connected components of filtrations.
-We stress again that this is to be intended as a roadmap, the outlining of a general idea and not as a formal and coherent description on how one can treat persistent sets via directed editable graphs.
-In fact, even in the simplest case, i.e.
-merge trees, there are non-trivial details which need to be taken care of when bridging between persistent sets and editable graphs (see Section 5.5 and Pegoraro (2024a)).
-Coherently, we explicitly point out which questions and problems should be assessed by future works.
-5.1 Finite Posets, DAGs and Transitive Reductions
-In this section, following previous works on graphs and posets, we obtain a graph representation of a poset, which we than employ to represent poset-indexed functors.
+A directed graph is made by a pair $G=(V,E)$ with $V$ being the set of its vertices and $E\subseteq V\times V$ the set of its directed edges.
 Consider a finite partially ordered set $(P,\leq)$.
@@
 Moreover, the graph $G_{P}$ can be simplified without losing ordering information.
-Definition 14 (adapted from Aho et al. (1972))
-A directed graph $G=(V,E)$ is called transitive if for every segment $s\in\mathcal{S}(G)$, $(s(0),s(1))\in E$.
-For every directed graph $G=(V,E)$ we define its transitive closure as the graph $G^{T}:=(V,E_{G^{T}})$, with $E_{G^{T}}=\bigcap\{E^{\prime}\subset V\times V\mid E\subset E^{\prime},(V,E^{\prime})\text{ transitive graph}\}$.
+We know formalize these constructions.
+Definition 7 (Aho et al. (1972))
+A directed graph $G=(V,E)$ is called transitive if $(v,v^{\prime})\in E$, whenever there is a directed path from $v$ to $v^{\prime}$.
+For every directed graph $G=(V,E)$ we define its transitive closure as the graph $G^{T}:=(V,E_{G^{T}})$, with:
+$E_{G^{T}}=\bigcap_{\{E^{\prime}\subset V\times V\mid E\subset E^{\prime},(V,E^{\prime})\text{ transitive graph}\}}E^{\prime}.$
 Theorem 1 (adapted from Aho et al. (1972))
@@
 Such graph is called the transitive reduction of $G$ and it is given by:
-$E_{G^{t}}:=\bigcap\{E^{\prime}\subset V\times V\mid(V,E^{\prime})^{T}=G^{T}\}.$
-Given a finite poset $P$, we call $\mathcal{G}(P)$ the transitive reduction of the associated DAG. See Figure 5 for some examples.
-5.2 From Persistent Sets to Editable Graphs
-Now, first we associate an editable graph structure to a poset $P$.
-And then we use this to work with functors of the form $S:P\rightarrow\operatorname*{Sets}$, which we call persistent sets, generalising terminology previously introduced by (Carlsson and Mémoli, 2013; Curry, 2018).
-If $P$ is finite, we propose two general ways in which we can associate weights to $\mathcal{G}(P)$, relying just on the structure of $P$:
-  1. 1.
-    suppose now that the poset $P$ is also endowed with a pseudo-metric (see Burago et al.
-(2022) for a definition) $d_{P}:P\times P\rightarrow\mathbb{R}_{\geq 0}$.
-Suppose moreover that any time $a<b$, then $d_{P}(a,b)>0$.
-Then we can associate to $P$ an editable graph $(\mathcal{G}(P),E_{\mathcal{G}(P)},\varphi_{\mathcal{G}(P)})$, with $\varphi_{\mathcal{G}(P)}:E_{\mathcal{G}(P)}\rightarrow\mathbb{R}_{>0}$ defined as $\varphi_{\mathcal{G}(P)}((a,b))=d_{P}(a,b)$.
-Note that for any poset $P$ with a monotone increasing function $f:P\rightarrow\mathbb{R}$ we can define $d_{P}(a,b):=\mid f(a)-f(b)\mid$ and this is a pseudo metric which satisfies our requirements since, if $a<b$, then $f(a)<f(b)$;
-  2. 2.
-    instead, suppose there is a map of posets $f:P\rightarrow X$, with $X$ editable space, and the poset structure on $X$ is as defined in 1.
-Suppose moreover we are in the situation to apply 1.
-Then, for every edge $(a,b)\in\mathcal{G}(P)$ there is one and only one $r\in X$ with $f(a)+r=f(b)$.
-Thus we can define $\varphi_{\mathcal{G}(P)}((a,b))=r$.
-As further step consider the case of functors of the form $S:P\rightarrow\operatorname*{Sets}$, with $P$ being a finite poset.
-We might refer to $P$ as the indexing poset.
-The functor $S$ can be used to push forward onto its image the partial order we have on $P$.
-We do so associating to $S$ a poset $D_{S}$, called the display poset of $S$, defined as follows.
+$E_{G^{t}}:=\bigcap_{\{E^{\prime}\subset V\times V\mid(V,E^{\prime})^{T}=G^{T}\}}E^{\prime}.$
+Given a finite poset $P$, we call $\mathcal{G}(P)$ the transitive reduction of the associated DAG.
+3.2 From Persistent Sets to Dendrograms
+We now establish a connection between functors, posets, and tree.
+To this end, we introduce a general procedure that applies beyond the specific case of functors $S:(\mathbb{R},\leq)\rightarrow\operatorname*{Sets}$, called persistent sets (Carlsson and Mémoli, 2013; Curry, 2018; Curry et al., 2022), which constitutes the primary focus of this work.
+Such objects are in fact of great interest in TDA and represent a (non-equivalent) alternative to persistence modules/diagrams when working with information collected from path-connected components of filtrations.
+A more thorough exploration of the broader scenarios is left for future research.
+The objects we introduce in the following are represented in Figure 2.
 Definition 8 (adapted from Curry et al. (2022))
-Given a functor $S:P\rightarrow\operatorname*{Sets}$, with $P$ being a poset, the display poset of $S$ is defined as the set of couples $D_{S}:=\{(p,s)\mid p\in P,s\in S(p)\}$.
+Given a functor $S:P\rightarrow\operatorname*{Sets}$, with $P$ being a poset, the display poset of $S$ is defined as the set of pairs $D_{S}:=\{(p,s)\mid p\in P,s\in S(p)\}$.
 This is a poset with $(p,s)\leq(p^{\prime},s^{\prime})$ if and only if $p\leq p^{\prime}$ and $S(p\leq p^{\prime})(s)=s^{\prime}$.
-For every display poset $D_{S}$ wee have a monotone increasing projection $h:D_{S}\rightarrow P$ defined by $(p,s)\mapsto p$.
+For every display poset $D_{S}$ we have a monotone increasing projection $h:D_{S}\rightarrow P$ defined by $(p,s)\mapsto p$.
+The poset structure on $P$ and the one on $D_{S}$ are consistent with their respective graph representations, as made precise in the following proposition.
 Proposition 3
@@
 Viceversa, any time we have $(p,s)\leq(p^{\prime\prime},s^{\prime\prime})$ and $p\leq p^{\prime}\leq p^{\prime\prime}$, by functoriality there exists also $(p^{\prime},s^{\prime})$ such that $(p,s)\leq(p^{\prime},s^{\prime})\leq(p^{\prime\prime},s^{\prime\prime})$.
-As a consequence, we can define $\varphi_{\mathcal{G}(D_{S})}:=h\circ\varphi_{\mathcal{G}(P)}$.
-5.3 Decorated Merge Trees: Function Valued Weights
-We now use the previous construction to tackle a particular kind of persistent sets.
-[…]
+Now we restrict our attention to persistent sets.
 Definition 9 ((Patel, 2018))
-A persistent set $S$ is called constructible if there is a finite collection of real numbers $\{t_{1}<t_{2}<\ldots<t_{n}\}$ such that:
+A constructible persistent set is a functor $S:(\mathbb{R},\leq)\rightarrow\operatorname*{Sets}$ such that there is a finite collection of real numbers $\{t_{1}<t_{2}<\ldots<t_{n}\}$ such that:
   • $S(t<t^{\prime})=\emptyset$ for all $t<t_{1}$;
@@
 Definition 10 (Pegoraro (2024a))
-A constructible persistent set $S$ is regular if for every critical value $t_{i}$ and for every $\varepsilon>0$ small enough, $S(t_{i}\leq t_{i}+\varepsilon)$ is a bijective map.
+A regular persistent set $S$ is a finite constructible persistent set such that, for every critical value $t_{i}$ and for every $\varepsilon>0$ small enough, $S(t_{i}\leq t_{i}+\varepsilon)$ is a bijective map.
 At this point we have the following immediate result, which we state without proof.
 Proposition 4
-Given two finite regular persistent sets $S$ and $S^{\prime}$, then $S\simeq S^{\prime}$ if and only if:
+Given two regular persistent sets $S$ and $S^{\prime}$, then $S\simeq S^{\prime}$ if and only if:
   • they have the same minimal set of critical values $C=\{t_{1},\ldots,t_{n}\}$
   • $S_{\mid C}\simeq S^{\prime}_{\mid C}$, with $S_{\mid C}:C\hookrightarrow\mathbb{R}\rightarrow\operatorname*{Sets}$.
-For the interested reader, we have very close relationships between the graph $\mathcal{G}(D_{S_{\mid C}})$ and the merge tree of $S$.
-Note that the upcoming proposition also further justifies our idea of working up to order $2$ vertices
+Without going into the details of the construction of a merge tree, we just show that tree-structures are the natural discrete tool to work with persistent sets. See also 2.
 Proposition 5
-Consider a finite regular persistence set $S$ and its minimal set of critical values $C$.
-If $S(t)$ has cardinality $1$ for $t$ big enough, then the graph $\mathcal{G}(D_{S_{\mid C}})$ is equivalent up to order $2$ vertices (as a directed graph) to the merge tree of $S$, upon removing from the merge tree the edge that goes to infinity.
-In particular, $\mathcal{G}(D_{S_{\mid C}})$ is a tree.
-As a consequence, if we then use item 1.
-in Section 5.2 with $P$ equal to $C$, the set of critical values of $S$, we obtain an editable graph with weights in $\mathbb{R}_{\geq 0}$ is which is very closely related to merge trees.
-In (Pegoraro, 2024a) this approach is used to represent merge trees and to adapt $d_{E}$ to such objects, obtaining an edit distance $d_{E}$ for merge trees.
-Moreover, putting together results in Pegoraro (2024a) it is shown that:
-$d_{I}(T_{f},T_{g})\leq d_{E}(T_{f},T_{g})\leq 2\cdot\#T_{f}\cdot\#T_{g}\parallel f-g\parallel_{\infty}$ (1)
-with:
-  • $f,g$ being two tame (Chazal et al., 2016) real valued functions defined on the same path-connected topological space;
-  • $T_{f}$ and $T_{g}$ are the merge trees associated to the sublevel set filtrations of $f$ and $g$;
-  • $d_{I}$ is the interleaving distance between merge trees (Morozov et al., 2013);
-  • $d_{E}$ is an adaptation of metric here induced to compare merge trees represented by means of editable graphs, essentially truncating the edge at infinity (with some additional details, see Pegoraro (2024a));
-  • #$\#T_{f}$ being the cardinality of the set of vertices of $T_{f}$.
-Note that Equation 1 represents a stability property in line with our expectations: since our metric is a summation of all the costs of the local modifications needed to align two trees we have that the to control the distance between trees with the sup norm between functions, we need a constant that scales linearly with the size of the trees.
-The interleaving distance, instead, is focused on determining the maximal local modifications to turn one tree into the other and it is well known to be universally stable, in the sense that is the biggest distance which is always bounded by the sup norm between the functions.
-Moreover, the $1$-Wasserstein distance between persistence diagrams satisfies and analogous relationship w.r.t.
-the bottleneck distances.
-As a consequence, our metric is better at picking up global differences and not just local differences (analogously to Wasserstein distances w.r.t bottleneck distance for persistence diagrams).
-On the other hand, this also means that is it more sensitive to noise.
-For more details and examples see Pegoraro (2024a).
+Consider a finite regular persistence set $S$ and a set of critical values $C$. If $S(t)$ has cardinality $1$ for $t$ big enough, $\mathcal{G}(D_{S_{\mid C}})$ is a tree.
+Remark 2
+In the proof of 5, we actually show that the graph $\mathcal{G}(D_{S_{\mid C}})$ is equivalent up to degree $2$ vertices (as a directed graph) to the merge tree of $S$, upon removing from the merge tree the edge that goes to infinity.
+In fact, proving the actual statement of 5 reduces to observing that, within the setting considered, $\mathcal{G}(D_{S_{\mid C}})$ is a rooted directed acyclic graph, with the $parent>child$ relationship being induced by the poset relationship on $D_{S_{\mid C}}$.
+Note that this proposition further justifies our idea of working up to degree $2$ vertices.
+[…]
+(b) The persistent set $\pi_{0}(X_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})$.
+(c) The display poset $D_{\pi_{0}(X_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$.
+[…]
+Figure 2: An example of a filtration along with the persistent set of its path-connected components, the associated display poset and its graph representation.
+The colors are used throughout the plots to highlight the relations between the different objects.
+4 Edits of Dendrograms
+Now we present the main contributions of the paper.
+The approach we follow is to define a distance inspired by Edit Distances (Tai, 1979), but with substantial differences in the edit operations.
+The philosophy of these distances is to allow certain modifications of the base object, called edits, each being associated to a cost, and to define the distance between two objects as the minimal cost that is needed to transform the first object into the second with a finite sequence of edits.
+In this way, up to properly setting up a set of edits, one can formalize the deformation of a dendrogram modifying the local information induced by the weight function defined on the edges and the tree structure itself.
+On top of that, edit distances frequently enjoy some decomposition properties which simplify the calculations (Hong et al., 2017), which are notoriously very heavy (Hein et al., 1995).
+We also point out that, in literature, tree are often considered with weights being real numbers and avoid modeling the weight space as we do with the definition of editable spaces.
+Given a space of dendrograms $(\mathcal{T},X)$, with $(X,\odot,0_{X})$ editable space, and given a dendrogram $(T,\varphi)$, we define our edits as operators $\{(T,\varphi)\}\rightarrow(\mathcal{T},X)$.
+So, an edit operation is not defined on the whole $(\mathcal{T},X)$, but only on $(T,\varphi)$, and its image can be defined up to isomorphism.
+Moreover, given two edits $e_{i}:\{(T,\varphi)\}\rightarrow(\mathcal{T},X)$ and $e_{j}:e_{i}(\{(T,\varphi)\})\rightarrow(\mathcal{T},X)$, we can consider their composition $e_{i}\circ e_{j}$.
+Any finite composition of edits is referred to as an edit path.
+The distance $d_{E}$ we introduce differs substantially from previously defined edit distances, as it is tailored specifically for comparing topological summaries—roughly meaning that topologically irrelevant points can be removed from a tree at no cost.
+Recall that, via the identification $E_{T}\cong V_{T}-\{r_{T}\}$, the upcoming transformations can be described equivalently in terms of either edges or vertices.
+  • Shrinking an edge means changing the weight value of the edge with a non zero element in $X$. The inverse of this transformation is the shrinking which restores the original edge weight.
+  • Deleting an edge $(v_{1},v_{2})$ results into a new tree, with the same vertices apart from $v_{1}$ (the lower one), and with the parent of the deleted vertex which gains all of its children.
+    The inverse of deletion is the insertion of an edge along with its child vertex.
+We can insert an edge at a vertex $v$ specifying the child of $v$ and its children (that can be either none or any portion of the children of $v$) and the weight of the edge.
+  • Lastly, we can eliminate a degree two vertex $v$, that is, a parent with an only child, connecting the two adjacent edges which arrive and depart from $v.$ The weight of the resulting edge is the sum of the weights of the joined edges.
+This transformation is the ghosting of the vertex $v$.
+Its inverse transformation is called the splitting of an edge.
+(a) Starting and target weighted trees, with highlighted the vertices involved in the edit path.
+(b) Deletions of $v_{1},v_{2},v_{3}$, in red.
+(c) Ghostings of $v_{4},v_{5}$, in yellow.
+(d) Shrinkings of $v_{6}$ and $v_{7}$ to match the weight, respectively, of $w_{6}$ and $w_{7}$, in green.
+(e) Splitting edges via $w_{4}$ and $w_{5}$, in yellow.
+(f) Insertions of $w_{1},w_{2},w_{3}$, in red.
+Figure 3: (b)$\rightarrow$(e) form an edit path made from the left weighted tree in Figure 3(a) to the right one.
+The edit path can be represented with a mapping - Section 4.2 - consisting of the pairs $(v_{i},\mathfrak{D})$ for all the red vertices in Figure 3(b), $(v_{i},\mathfrak{G})$ for all the yellow vertices in Figure 3(c), $(v_{i},w_{i})$ for all the vertices associated via the green color in Figure 3(d), $(\mathfrak{G},w_{i})$ for all the yellow vertices in Figure 3(c) and $(\mathfrak{D},w_{i})$ for all the red vertices in Figure 3(f).
+See Figure 3 for an example of an edit path between two graphs whose weights are edges lenghts - Section 2.1.1.
+Remark 3
+Shrinking and deletions are classical edits for trees, with shrinking being usually referred to as relabeling.
+The ghosting edit (and its inverse) is completely unusual and we are not aware of any previous work employing it.
+Its definition is largely driven by the TDA perspective presented in the Introduction and it is fundamental for the stability results contained in Pegoraro and Secchi (2024) and Pegoraro (2024a).
+Even in the context of graph edit distances (GEDs) (Ambauen et al., 2003; Gao et al., 2010; Zeng et al., 2009; Lerouge et al., 2016; Serratosa, 2021), we are not aware of any existing work that adopts definitions similar to ours.
+While vertex splitting and merging in GEDs bear some resemblance to our notions of insertion and deletion, there no correspondence for splittings and ghostings as we define them.
+We also give the following definition.
+Definition 11
+Dendrograms are equal up to degree $2$ vertices if they become isomorphic after applying a finite number of ghostings or splittings.
+We write $(T,\varphi_{T})\cong_{2}(T^{\prime},\varphi_{T^{\prime}})$.
+11 induces an equivalence relation which identifies the set of dendrograms inside $(\mathcal{T},X)$ that we want to treat as equal.
+We call $(\mathcal{T}_{2},X)$ the space of equivalence classes of dendrograms in $(\mathcal{T},X)$, equal up to degree $2$ vertices.
+4.1 Costs of Edit Operations
+Now we associate to every edit a cost, that is, a length measure in the space $(\mathcal{T},X)$.
+In light of this interpretation, we will often use the words “length” and “cost” interchangeably when referring to an edit path.
+The costs of the edit operations are defined as follows:
+  • if, via shrinking, an edge goes from weight $x$ to weight $y$, then the cost of such operation is $d_{X}(x,y)$;
+  • for any deletion/insertion of an edge with weight $x$, the cost is equal to $d_{X}(x,0_{X})$;
+  • the cost of ghosting/splitting operations is always zero.
+This is coherent with the properties of editable spaces: consider a ghosting merging two adjacent edges separated by a degree two vertex, with $x$ being the weight of one edge, and $y$ the cost of the other.
+The total weight of the two edges is the same before and after the ghosting: $\mid d_{X}(x\odot y,0_{X})-d_{X}(x,0_{X})-d_{X}(y,0_{X})\mid=0$.
+Definition 12
+Given two dendrograms $T$ and $T^{\prime}$ in $(\mathcal{T},X)$, define:
+  • $\Gamma(T,T^{\prime})$ as the set of all finite edit paths between $T$ and $T^{\prime}$;
+  • $cost(\gamma)$ as the sum of the costs of the edits for any $\gamma\in\Gamma(T,T^{\prime})$;
+  • the multigraph edit distance as:
+    $d_{E}(T,T^{\prime})=\inf_{\gamma\in\Gamma(T,T^{\prime})}cost(\gamma)$
+By definition the triangle inequality and symmetry must hold, but, up to now, this edit distance is intractable; one would have to search for all the possible finite edit paths which connect two dendrograms in order to find the minimal ones.
+On top of that, having an edit which is completely “for free”, it is not even obvious that $d_{E}(T,T^{\prime})>0$ for some dendrograms.
+However, it is clear that $d_{E}$ induces a pseudo-metric on classes of dendrograms up to degree two vertices.
+4.2 Mappings
+We introduce a combinatorial tool, called a mapping, which plays a central role in developing the theoretical properties of $d_{E}$ and in bringing it into the realm of computability.
+In particular, we prove that mappings can be used to parametrize a finite set of edit paths, with distinct and interpretable characteristics, which is guaranteed to contain at least an optimal edit path between two dendrograms.
+Definition 13
+A mapping between two dendrograms $T$ and $T^{\prime}$ is a set $M\subset(E_{T}\cup\{\mathfrak{D},\mathfrak{G}\})\times(E_{T^{\prime}}\cup\{\mathfrak{D},\mathfrak{G}\})$ satisfying:
+  • (M1)
+    consider the projection of the Cartesian product $(E_{T}\cup\{\mathfrak{D},\mathfrak{G}\})\times(E_{T^{\prime}}\cup\{\mathfrak{D},\mathfrak{G}\})\rightarrow(E_{T}\cup\{\mathfrak{D},\mathfrak{G}\})$; we can restrict this map to $M$ obtaining $\pi_{T}:M\rightarrow(E_{T}\cup\{\mathfrak{D},\mathfrak{G}\})$.
+The maps $\pi_{T}$ and $\pi_{T^{\prime}}$ are surjective on $E_{T}$ and $E_{T^{\prime}}$, i.e.
+$E_{T}\subset\operatorname*{Im}(\pi_{T})$ and $E_{T^{\prime}}\subset\operatorname*{Im}(\pi_{T^{\prime}})$;
+  • (M2)
+    $\pi_{T}$ and $\pi_{T^{\prime}}$ are injective on $M\cap(E_{T}\times E_{T^{\prime}})$;
+  • (M3)
+    given $(a,b)$ and $(c,d)\in M\cap(V_{T}\times V_{T^{\prime}})$, $a>c$, if and only if $b>d$;
+  • (M4)
+    if $(a,\mathfrak{G})\in M$ (or analogously $(\mathfrak{G},a)$), then after applying all deletions of the form $(v,\mathfrak{D})\in M$, the vertex $a$ becomes a degree $2$ vertex.
+In other words: let $child(a)=\{b_{1},..,b_{n}\}$.
+Then there is exactly one $i$ such that for all $j\neq i$, for all $v\in V_{sub(b_{j})}$, we have $(v,\mathfrak{D})\in M$; and there is one and only one $c$ such that $c=\max\{x<b_{i}\mid(x,y)\in M$ for any $y\in V_{T^{\prime}}\}$.
+We call $\operatorname*{Mapp}(T,T^{\prime})$ the set of all mappings between $T$ and $T^{\prime}$.
+We can always have a mapping made by:
+$\{(e,\mathfrak{D})\mid e\in E_{T}\}\cup\{(\mathfrak{D},e)\mid e\in E_{T^{\prime}}\}.$
+In the following, we will see that such a mapping corresponds to the edit path which deletes all the edges of one dendrogram and inserts all the edges of the second.
+We may refer to edges which appear in the pairs in $M\cap(V_{T}\times V_{T^{\prime}})$ as the paired or matched edges/vertices.
+Remark 4
+In literature usually mappings are subsets of $E_{T}\times E_{T^{\prime}}$ and their properties are equivalent to (M2) and (M3) (see Hong et al.
+(2017); Sridharamurthy et al.
+(2020) and references therein).
+In this way, mappings coherently maps edges into edges, respecting the tree ordering, inducing in this way the needed relabeling operations - see also 3.
+The differences with our definition are caused by the introduction of the ghosting edits.
+Accordingly, the way in which mappings parametrize edit paths in the present work is completely novel and, for instance, it establishes some (partial) ordering between the edits.
+Note that the set of mappings we consider is bigger than in usual edit distances - thus computing the distance has a higher computational cost.
+Every $M\in\operatorname*{Mapp}(T,T^{\prime})$ parametrizes a set of edit paths, with identical cost, as follows:
+  • $\gamma_{d}^{T}$ is made by the deletions to be done on $T$, that is, the pairs $(v,\mathfrak{D})$, executed in any order.
+So we obtain $T^{M}_{d}=\gamma_{d}^{T}(T)$, which is well defined and does not depend on the order of the deletions.
+Similarly, we define $\gamma_{d}^{T^{\prime}}$ as a path made by the deletions to be done on $T^{\prime}$, that is, the pairs $(\mathfrak{D},w)$, executed in any order, and obtain $T^{\prime M}_{d}=\gamma_{d}^{T^{\prime}}(T^{\prime})$.
+  • One then proceeds editing $T^{M}_{d}$ by ghosting all the vertices $(v,\mathfrak{G})$ in $M$, in any order, getting a path $\gamma^{T}_{g}$ and the dendrogram $T_{M}:=\gamma^{T}_{g}\circ\gamma^{T}_{d}(T)$.
+As before, we can do an analogous procedure on $T^{\prime M}_{d}$, ghosting all the vertices $(\mathfrak{G},w)$ in $M$, in any order, and building a path $\gamma^{T^{\prime}}_{g}$, along with the dendrogram $T^{\prime}_{M}:=\gamma^{T^{\prime}}_{g}\circ\gamma^{T^{\prime}}_{d}(T^{\prime})$.
+  • Since all the remaining points in $M$ are paired, the two dendrograms $T^{\prime}_{M}$ and $T_{M}$ must be isomorphic as tree structures.
+This is guaranteed by the properties of $M$.
+So one can shrink $T_{M}$ onto $T^{\prime}_{M}$, and the composition of the shrinkings, executed in any order, gives an edit path $\gamma_{s}^{T}$.
+By construction $\gamma_{s}^{T}\circ\gamma_{g}^{T}\circ\gamma_{d}^{T}(T)=T^{\prime}_{M}$, and $(\gamma_{d}^{T^{\prime}})^{-1}\circ(\gamma_{g}^{T^{\prime}})^{-1}\circ\gamma_{s}^{T}\circ\gamma_{g}^{T}\circ\gamma_{d}^{T}(T)=T^{\prime}$.
+Where the inverse of an edit path is thought as the composition of the inverses of the single edit operations, taken in the inverse order.
+We call $\gamma_{M}$ the set of all possible edit paths of the form $(\gamma_{d}^{T^{\prime}})^{-1}\circ(\gamma_{g}^{T^{\prime}})^{-1}\circ\gamma_{s}^{T}\circ\gamma_{g}^{T}\circ\gamma_{d}^{T}$, obtained by changing the order in which the edit operations are executed inside $\gamma_{d}$, $\gamma_{g}$ and $\gamma_{s}$.
+Even if $\gamma_{M}$ is a set of paths, its cost is well defined:
+$cost(M):=cost(\gamma_{M})=cost(\gamma_{d}^{T})+cost(\gamma_{s}^{T})+cost(\gamma_{d}^{T^{\prime}}).$
+We prove that there exists always a mapping $M$ such that the paths in $\gamma_{M}$ are optimal.
+Theorem 2 (Main Theorem)
+Given two dendrograms $T$ and $T^{\prime}$, for every finite edit path $\gamma$, there is a mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ such that $cost(M)\leq cost(\gamma)$.
+Two corollaries follow immediately from the facts that $\operatorname*{Mapp}(T,T\textquoteright)$ ’ ) is finite and that any mapping with cost zero must consist entirely of ghostings and splittings.
+Corollary 2
+For any two dendrograms $T$ and $T^{\prime}$, the set $\operatorname*{Mapp}(T,T^{\prime})$ is finite. Thus, we have the following well defined pseudo-metric between dendrograms:
+$d_{E}(T,T^{\prime})=\inf\{cost(\gamma)\mid\gamma\in\Gamma(T,T^{\prime})\}=\min\{cost(M)\mid M\in\operatorname*{Mapp}(T,T^{\prime})\}$
+which we will refer to as the edit distance between $T$ and $T^{\prime}$.
+Corollary 3
+Given $T$ and $T^{\prime}$ dendrograms, $d_{E}(T,T^{\prime})=0$ if and only if $T$ and $T^{\prime}$ are equal up to degree $2$ vertices.
+In other words $d_{E}$ is a metric for dendrograms considered up to degree $2$ vertices.
+To conclude this section, we define a particular family of mappings which help us in further restricting the search space for optimal edit paths.
+Definition 14
+A mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ has maximal ghostings if the following hold: $(v,\mathfrak{G})\in M$ if, and only if, $v$ is of degree $2$ after the deletions in $T$ and, similarly, $(\mathfrak{G},w)\in M$ if, and only if, $w$ is of degree $2$ after the deletions in $T^{\prime}$.
+A mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ has minimal deletions if the following hold: $(v,\mathfrak{D})\in M$ implies that neither $v$ nor $parent(v)$ are of degree $2$ after applying all the other deletions in $T$ and, similarly, $(\mathfrak{D},w)\in M$ implies that neither $w$ nor $parent(w)$ are of degree $2$ after applying all the other deletions in $T^{\prime}$.
+We collect all mappings with maximal ghostings in the set $M_{\mathcal{G}}(T,T^{\prime})$ and paths with minimal deletions in the set $M_{\mathcal{D}}(T,T^{\prime})$.
+Lastly, we set: $M_{2}(T,T^{\prime}):=M_{\mathcal{G}}(T,T^{\prime})\bigcap M_{\mathcal{D}}(T,T^{\prime})$.
+In other words with mappings in $M_{2}(T,T^{\prime})$ we are always eliminating all the degree $2$ vertices which arise from deletions and we are not deleting edges which we can shrink.
+The following lemma then applies.
+Lemma 1
+$\min\{cost(M)\mid M\in\operatorname*{Mapp}(T,T^{\prime})\}=\min\{cost(M)\mid M\in M_{2}(T,T^{\prime})\}.$
+5 A Discussion on the Use of Different Editable Spaces
+In this section we discuss some key examples which regard dendrograms with values in different editable spaces.
+We stress again that this is to be intended as a roadmap, the outlining of a general idea and not as a formal and coherent description on how one can treat persistent sets via dendrograms.
+Coherently, we explicitly point out which questions and problems should be assessed by future works.
+5.1 Function and Vector Valued Weights
+[…]
 […]
   1. 1.
-    diagram or Betti numbers decorated merge trees, on the same line of thought of Curry et al. (2022, 2023b);
+    diagram or Betti numbers enriched persistent sets, on the same line of thought of Curry et al. (2022, 2023b);
   2. 2.
-    merge trees decorated with the measure of sublevel sets.
-We illustrate this with two qualitative examples: first we consider the datum of a point cloud a then a function defined on a open subset of $\mathbb{R}^{n}$.
+    persistent sets enriched with the measure of sublevel sets.
+[…]
+We illustrate this with two examples: first we consider the datum of a point cloud a then a function defined on a open subset of $\mathbb{R}^{n}$.
 Let $Y^{k}\subset\mathbb{R}^{n}$ be a finite set with $k$ points. A widely employed filtration is:
-$Y_{t}^{k}:=\cup_{p\in Y^{k}}B_{t}(p)$
+$Y_{t}^{k}:=\bigcup_{p\in Y^{k}}B_{t}(p)$
 $B_{t}(p)$ being the metric ball centred in $p$, which is also called the offset filtration.
-Any point $D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$ is a couple $(t,U)$, with $t\in\mathbb{R}$ and $U$ being a path connected component of $Y_{t}^{k}$.
-Then, one collect some topological information about $U$ and consider a function $f$ with values in $X=\mathbb{R}^{N}_{\geq 0}$, for some $N<n$, defined by:
-$f((t,U))=(\beta_{0}(U),\beta_{1}(U),\ldots,\beta_{N}(U)),$
+Any point $D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$ is a pair $(t,U)$, with $t\in\mathbb{R}$ and $U$ being a path connected component of $Y_{t}^{k}$.
+Then, one collect some topological information about $U$ and consider a vector-valued function $f$ with values in $X=\mathbb{R}^{N}_{\geq 0}$, for some $N<n$, defined by:
+$f((t,U))=(\beta_{0}(U),\beta_{1}(U),\ldots,\beta_{N}(U)),$ (1)
 with $\beta_{i}(U)=\dim\text{H}_{i}(U,\mathbb{R})$ being the $i$-th Betti number of $U$ clearly $\beta_{0}(U)=1$.
 For an example see Pegoraro (2024b).
-An alternative and richer way to capture shape-related information about $U$ is to consider the point cloud: $U_{k}:=Y^{k}\cap U$ and then build $\text{PD}_{q}(U_{k})$, the persistence diagram of the offset filtration of $U_{k}$ using $q$-dimensional homology.
-Consequently, one can define $f$ with values in $X=D(\mathbb{R}^{2}_{\Delta},\Delta)$, the space of persistence diagrams with the $1$-Wasserstein distance, defined by:
-$f((t,U))=\text{PD}_{q}(U_{k}).$
+An alternative and richer way to capture shape-related information about $U$ is to consider the point cloud: $U_{k}:=Y^{k}\cap U$ and then build $\text{PD}_{i}(U_{k})$, the persistence diagram of the offset filtration of $U_{k}$ using $i$-dimensional homology.
+Consequently, one can define $f$ with values in $X=D(\mathbb{R}^{2}_{\Delta},\Delta)\times\ldots D(\mathbb{R}^{2}_{\Delta},\Delta)$, the space of persistence diagrams with the $1$-Wasserstein distance, defined by:
+$f((t,U))=(\text{PD}_{0}(U_{k}),\ldots,\text{PD}_{N}(U_{k})).$
 See Figure 4(a) and Figure 4(b) for a visual example.
@@
 And consider $D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$.
-This time another shape-related information that we can estract from a couple $(t,U)\in D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$, is the Lebesgue measure of $U$.
+This time another shape-related information that we can extract from a pair $(t,U)\in D_{\pi_{0}(Y_{\operatorname*{\scalerel*{\cdot}{\bigodot}}})}$, is the Lebesgue measure of $U$.
 And thus define:
@@
 For an example of situations that can be tackled with this approach see Figure 4(c) and Figure 4(d).
-Note that the representation of functional data obtained with the volume-decorated merge trees, lies in between information captured by $L_{p}$ metrics and purely topological information: with merge trees one would not be able to distinguish the two functions in Figure 4(c) and Figure 4(d) (they have the same merge trees), but with $L_{p}$ metrics one would not capture any of the shape similarities between the functions.
-Moreover, functional data analysis approaches capturing the shape of functions - e.g.
-aligning functions with re-parametrizations, see for instance the Special Section on Time Warpings and Phase Variation on the Electronic Journal of Statistics, Vol 8 (2) - are difficult to develop over $\mathbb{R}^{n}$, $n\geq 2$.
+Note that the representation of functional data obtained with the volume-enriched persistent sets, lies in between information captured by $L_{p}$ metrics and purely topological information: with persistent sets one would not be able to distinguish the two functions in Figure 4(c) and Figure 4(d) (they have isomorphic persistent sets), but with $L_{p}$ metrics one would not capture any of the shape similarities between the functions.
+We therefore believe that volume-enriched persistent sets can be a useful tool to employ when there are alignment issues with functional data i.e.
+all those situations where one may want to resort to some reparametrization of the domain - see, for instance, the Special Section on Time Warpings and Phase Variation on the Electronic Journal of Statistics, Vol 8 (2).
+One such example can be found in (Pegoraro, 2024b, Appendix,).
 […]
-[…]
-$d_{E}(T_{f},T_{g})\leq(K+\max\{\text{diam}Y,\text{diam}Y^{\prime}\})\cdot 4\cdot\#Y\cdot\#Y^{\prime}\cdot d_{G-H}(Y,Y^{\prime}),$
-with $d_{H}(Y,Y^{\prime})$ being the Gromov-Hausdorff distance between compact metric spaces.
-We believe that the first two approaches have the potential to be powerful shape summaries to distinguish between point clouds, as also shown with several examples by Curry et al.
-(2022, 2023b), employing similar (though not metrically equivalent) tools.
-Instead, for more details and examples on the use of functional weights for editable graphs, see Pegoraro (2024b).
-Note that between the metric $d_{E}$, applied to this context, and the metric developed in Curry et al.
-(2022), hold similar considerations to the ones we made comparing $d_{E}$ and the interleaving distance between merge trees.
-(a) A point cloud (left), with the thickening giving the off-set filtration at some particular value (dashed horizontal cyan line) and the merge tree of the filtration (right).
+(a) A point cloud (left), with the thickening giving the off-set filtration at some particular value (dashed horizontal cyan line) and the display poset of (the persistent set of) the filtration (right).
 At the chosen value the filtration has three path connected components, shown in different colours, each with different shapes, characterized by the (qualitative) associated persistence diagram.
@@
 (c) A real valued function defined on a square with four distinct local minima, associated to downward peaks with different shapes.
-The associated decorated merge tree (right) accordingly shows four branches (black) and, in red, the (qualitative) volume of the associated downward peaks included in the sublevel sets, which increases with the value of the filtration.
+The associated display poset (of the persistent set) accordingly shows four branches (black) and, in red, the (qualitative) volume of the associated downward peaks included in the sublevel sets, which increases with the value of the filtration.
 (d) Same situation as in Figure 4(c), but for a different image.
 We clearly see that the four downward peaks have very different sizes, and that is reflected by the very different volume profiles attached to the branches of the merge tree.
-Note that the two images have the same (non-decorated) merge trees.
+Note that the two images have the same merge trees.
 Figure 4: Plots related to Section 5.1.
-5.4 Persistent Sets with Embedded Posets: Vector Valued Weights
-Here we consider the case of a persistent set $F:P\hookrightarrow\mathbb{R}^{n}\rightarrow\operatorname*{Sets}$ and focus on the approach described by item 2.
-in Section 5.2.
-Where $P$ is a finite poset and the poset structure in $\mathbb{R}^{n}$ is induced by $r=(r_{1},\ldots,r_{n})\leq r^{\prime}=(r^{\prime}_{1},\ldots,r^{\prime}_{n})$ if $r_{i}\leq r^{\prime}_{i}$ for all $i$.
-This is the kind of situations that works studying multiparameter persistence care about, by making finiteness assumptions like finitely presentable persistence modules (Oudot and Scoccola, 2024) - that is, modules on $\mathbb{R}^{n}$ which are determined by their values on a finite sub-poset of $\mathbb{R}^{n}$.
-We associate to $F$ a unique editable graph built in this way:
-  • we take $P$, build the associated DAG and take its transitive reduction $\mathcal{G}(P)$;
-  • build the weight function $\varphi_{\mathcal{G}(P)}$ as in Section 5.2, item 2), so that each edge between $a,b\in\mathbb{R}^{n}_{\geq 0}$, with $a\leq b$ we associate the vector $b-a$;
-  • finally we build the display poset of $F_{\mid P}$, which we call $D_{F}$, take its transitive reduction $\mathcal{G}(D_{F})$ and build the weight function $\varphi_{\mathcal{G}(D_{F})}$ by pushing forward $\varphi_{\mathcal{G}(P)}$.
-See Figure 5 for a detailed example.
-We believe that the metric $d_{E}$ is a natural metric to employ to compare these kind of graph representations for multidimensional persistent sets.
-In fact, the metric exhibits some natural behaviour, for instance as in Figure 6 and, in case $n=1$, and has proven stability results - which we reported in previous sections.
-(a) A finite sub-poset of $\mathbb{R}^{2}$.
-(b) The transitive reduction of the poset in Figure 5(a).
-(c) The DAG associated to a persistent set on the poset in Figure 5(a).
-Arrows are colored consistently with the ones in the poset in Figure 5(a) and represent the weights associated to the corresponding edges.
-The maps between sets are represented by colors and follow (row-wise and column-wise) the elder rule.
-(d) The editable graph associared to the transitive reduction of the DAG in Figure 5. As in Figure 5, arrows represent weights in $\mathbb{R}^{2}_{\geq 0}$
-Figure 5: Plots related to Section 5.4.
-(a) Another DAG representing a persistent set over a poset in $\mathbb{R}^{2}$.
-(b) Starting from the DAG in Figure 6(a), with two ghosting edits, we can replace the two paths around the square associated to the pink sets, with the two diagonal edge.
-Thus the DAG in Figure 6(a) and this one are identified by our edit distance.
-Figure 6: Additional plots related to Section 5.4.
 5.2 Problems to be Faced
-In the previous sections we have laid out some ideas which suggest how different persistent sets can be represented via editable graphs with weight functions taking values in different editable spaces.
-As pointed out at the beginning of Section 5, these ideas present some shortcomings which need to be overcome and some formal results which need to be established in order for the ideas to be properly applied.
-In this section we present the major issues affecting the proposed approach: some of them are already tackled and solved in other works, while other are left open and serve as invitations for future investigations.
-In the same way a point cloud is represented by the pairwise distances between its points only up to isometry, similarly, collecting distances or vectors as weights for the edges does not provide a map from persistent sets to graphs that can be inverted up to natural isomorphisms.
-Consider the case of a one dimensional persistent set $\pi_{0}\circ Y:\mathbb{R}\rightarrow\operatorname*{Sets}$ and define for every $h\in\mathbb{R}$ the filtration $Y_{h}(t):=Y(t+h)$.
-If we then proceed by building the display poset and obtaining the associated graph as discussed in Section 5.4, then the editable graphs associated to $\pi_{0}\circ Y$ and $\pi_{0}\circ Y_{h}$ are going to be isomorphic as editable graphs, as their posets are going to be isomorphic as posets and $r\mapsto r+h$ is an isometry.
-While the persistent sets $\pi_{0}\circ Y$ and $\pi_{0}\circ Y_{h}$ are not isomorphic as functors.
-And this is true also for $Y:\mathbb{R}^{n}\rightarrow\operatorname*{Sets}$ and $h\in\mathbb{R}^{n}$.
-As already mentioned, the fiber of the map sending a constructible persistent set $S$ into the editable graph $\mathcal{G}(D_{S_{\mid C}})$ (up to isomorphism) contains a lot of persistent sets which are not isomorphic.
-Moreover, we believe that, for any representation of persistent set/persistence module, it is always advisable to produce some stability result, to ensure the interpretability of the pipeline.
-In the particular case of $\pi_{0}\circ Y:\mathbb{R}\rightarrow\operatorname*{Sets}$, both the issues presented have been taken care in Pegoraro (2024a).
-For all other scenarios these problems still await formal investigations.
-Additionally, we would like also to extend this framework to Reeb graphs, similarly to what was done for merge trees.
-This is motivated by the extent to which Reeb graphs (Biasotti et al., 2008; De Silva et al., 2016) and their combinatorial approximation, namely mapper graphs (Singh et al., 2007) are studied (Carriere et al., 2018; Carriere and Oudot, 2018) and used in applications (Nielson et al., 2015; Rizvi et al., 2017).
-They add another layer of difficulties as a metric being insensitive to topological changes as the ones depicted in Figure 6 it is not desirable.
-This could be solved by modeling the possibility of having multiple edges connecting two vertices.
-It would also be important also to develop an algorithm to compute or approximate $d_{E}$ which works for Reeb graphs.
-6 Trees
-In the remaining part of the manuscript we specialize our discussion on a particular kind of directed graphs: trees.
-Thanks to 6 we have seen that this scenario still includes a lot of interesting situations, covering all constructible (one dimensional) persistent sets whose filtration which definitively is path-connected.
-Moreover, the peculiar structure of trees allows us to prove some key results which greatly decrease the complexity of computing optimal mappings between two trees, resulting in a bottom-up algorithm which is implemented and tested.
-We leave to future works the investigation of algorithms for more complicated kinds of graphs.
-Remark 5
-From now on our point of view is predominantly computational, thus we always assume to have finite graphs.
-Definition 18
-A tree structure $T$ is given by a connected rooted acyclic graph $(V_{T},E_{T})$.
-We indicate the root of the tree with $r_{T}$.
-The order of a vertex $v\in V_{T}$ is the number of edges which have that vertex as one of the extremes, and is called $ord_{T}(v)$.
-Any vertex with an edge connecting it to the root is its child and the root is its father: this is the first step of a recursion which defines the father and children relationship for all vertices in $V_{T}.$ The vertices with no children are called leaves and are collected in the set $L_{T}$.
-The relation $child<father$ generates a partial order on $V_{T}$ which induces an orientation on the edges.
-The edges in $E_{T}$ are identified in the form of ordered couples $(a,b)$ meaning $a<b$ and $a\rightarrow b$.
-A subtree of a vertex $v$, called $sub_{T}(v)$, is the tree structure whose set of vertices is $\{x\in V_{T}\mid x\leq v\}$.
-To avoid overlapping some terminology with already existing notions of trees and for ease of notation, we refer to an editable tree by calling it dendrogram, as formalized by the following definition.
-Definition 19
-A tree structure $T$ with a weight function $\varphi_{T}:E_{T}\rightarrow X-\{0_{X}\}$, with $X$ editable space, is called dendrogram.
-The space of $X$-valued dendrograms is indicated as $(\mathcal{T},X)$.
-A key fact is that given a tree structure $T$, identifying an edge $(v,v^{\prime})$ with its lower vertex $v$, gives a bijection between $V_{T}-\{r_{T}\}$ and $E_{T}$, that is $E_{T}\cong V_{T}-\{r_{T}\}$ as sets.
-Given this bijection, we often use $E_{T}$ to indicate the vertices $v\in V_{T}-\{r_{T}\}$, to simplify the notation.
-Moreover, since we now have a canonical way to associate edges to vertices, we can state the definition of mappings in a more directed and handy way, associating an edit operation to every edge/vertex in a tree - a part from the root.
-Such association is not available in general for graphs.
-In this way the ghosting of an edge can be described via the removal of individual vertices, and similarly deletions and shrinkings can be referred to as applied on vertices instead of edges.
-With an abuse of notation we call mappings the following objects, which from now on replace the mappings defined in Section 4.2.
-The letters "$D"$ and "$G"$ are used to indicate deletion and ghosting.
-A mapping between $T$ and $T^{\prime}$ is thus a set $M\subset(E_{T}\cup\{"D","G"\})\times(E_{T^{\prime}}\cup\{"D","G"\})$ satisfying:
-  • (M1)
-    consider the projection of the Cartesian product $(E_{T}\cup\{"D","G"\})\times(E_{T^{\prime}}\cup\{"D","G"\})\rightarrow(E_{T}\cup\{"D","G"\})$; we can restrict this map to $M$ obtaining $\pi_{T}:M\rightarrow(E_{T}\cup\{"D","G"\})$.
-The maps $\pi_{T}$ and $\pi_{T^{\prime}}$ are surjective on $E_{T}\subset(E_{T}\cup\{"D","G"\})$ and $E_{T^{\prime}}\subset(E_{T^{\prime}}\cup\{"D","G"\})$ respectively;
-  • (M2)
-    $\pi_{T}$ and $\pi_{T^{\prime}}$ are injective;
-  • (M3)
-    $M\cap(V_{T}\times V_{T^{\prime}})$ is such that, given $(a,b)$ and $(c,d)\in M\cap(V_{T}\times V_{T^{\prime}})$, $a>c$, if and only if $b>d$;
-  • (M4)
-    if $(a,"G")\in M$ (or analogously ("$G",a)$), then after applying all deletions of the form $(v,"D")\in M$, the vertex $a$ becomes an order $2$ vertex.
-In other words: let $child(a)=\{b_{1},..,b_{n}\}$.
-Then there is one and only one $i$ such that for all $j\neq i$, for all $v\in V_{sub(b_{j})}$, we have $(v,"D")\in M$; and there is one and only one $c$ such that $c=\max\{x<b_{i}\mid(x,y)\in M$ for any $y\in V_{T^{\prime}}\}$.
-We highlight some points with the following remarks.
-Remark 6
-This definition is equivalent to the one given in Section 4.2.
-In fact: (M1) and (M2) are equivalent to (G1) plus the fact that vertices not appearing in the image of $\eta_{G}^{M}:\mathcal{E}_{M}\rightarrow E_{T}$ are deleted.
-Property (G2) is equivalent to (M4): if a segment $s$ which is not simply an edge, and so it induces some ghostings - is such that all the edges in its open star are either deleted or belong to $Im(s)$, this means that all the vertices in $s^{o}$ are of order $2$ after the deletions.
-Lastly, (M3) implies that the trees considered are isomorphic as directed graphs once all deletions and ghostings are applied on the starting and target trees.
-Remark 7
-In literature usually mappings are subsets of $E_{T}\times E_{T^{\prime}}$ and their properties are equivalent to (M2) and (M3) (see Hong et al.
-(2017); Sridharamurthy et al.
-(2020); Wetzels et al.
-(2022) and references therein).
-In this way, mappings coherently maps edges into edges, respecting the tree ordering, inducing in this way the needed relabeling operations - see also 3.
-The differences with our definition are caused by the introduction of the ghosting edits.
-Accordingly, the way in which mappings parametrize edit paths in the present work is completely novel and, for instance, it establishes some (partial) ordering between the edits.
-Note that the set of mappings we consider is bigger than in usual edit distances - thus computing the distance has a higher computational cost.
-6.1 Computing the Edit Distance for Trees: Decomposition Properties
-In these last sections we develop an algorithm to compute the edit distance for dendrograms.
+In the previous section, we outlined ideas suggesting that different persistent sets can be represented by dendrograms equipped with weight functions taking values in various editable spaces.
+As noted at the beginning of Section 5, these ideas require formal validation before they can be reliably applied.
+In this section, we briefly discuss the main challenges associated with this proposed framework.
+Some of these challenges have been addressed in separate work, while others remain open and point to promising directions for future research.
+There are three main issues that needs to be taken care of:
+  1. 1.
+    Discretization: Section 3.2 presents a procedure for turning a persistent set into a tree structure.
+This process depends crucially on the set of critical values $C$, and must be paired with the definition of an appropriate weight function, in order to obtain a dendrogram.
+Each of these steps—the discretization and the weight assignment—must be analyzed both individually and in combination.
+For example, one must verify whether the mapping from a persistent set to a dendrogram defines an injective transformation.
+This is not the case for the merge tree representation in Wetzels et al.
+(2022), while injectivity is ensured in Pegoraro (2024a, b).
+  2. 2.
+    Interpretability: the pipeline of representing a datum via an enriched persistent set and then with a dendrogram with values in an editable space, is in itself complex and needs to be carefully designed.
+Beyond the construction, it is also crucial to evaluate how well the edit distance $d_{E}$ captures variability between dendrograms.
+For instance, if the editable space is $\mathbb{R}_{\geq 0}$, it is clear that editing a positively weighted tree with our edits, always gives a positively weighted tree.
+However, when working with functions defined on display posets, Pegoraro (2024b) shows that only a subset of dendrograms corresponds to actual functions.
+Therefore, additional care is needed to ensure that $d_{E}$ meaningfully captures the variability between such functions.
+  3. 3.
+    Stability: closely related to interpretability is the issue of establishing stability guarantees specific to the chosen pipeline, akin to the results found in Pegoraro and Secchi (2024); Pegoraro (2024b).
+These results are essential to ensure that small perturbations in the data lead to correspondingly small changes in the associated dendrograms under $d_{E}$.
+5.3 Stability Properties
+In Section 5.2, we mention that establishing stability results is of pivotal importance to validate theoretically a pipeline involving dendrograms.
+Since some stability results have already been obtained in Pegoraro (2024a, b), we briefly describe them in order to give a qualitative idea of how the metrics $d_{E}$ behaves in different contexts.
+In Pegoraro (2024a); Pegoraro and Secchi (2024) $d_{E}$ is extended (via some non-trivial steps) to merge trees.
+Building on such definition, as already mentioned in the introduction, it is therein shown that:
+$d_{I}(T_{f},T_{g})\leq d_{E}(T_{f},T_{g})\leq 2\cdot\text{size}(T_{f})\cdot\text{size}(T_{g})\parallel f-g\parallel_{\infty}$ (2)
+with:
+  • $f,g$ being two tame (Chazal et al., 2016) real valued functions defined on the same path-connected topological space;
+  • $T_{f}$ and $T_{g}$ are the merge trees associated to the sublevel set filtrations of $f$ and $g$;
+  • $d_{I}$ is the interleaving distance between merge trees (Morozov et al., 2013);
+  • $d_{E}$ is an adaptation of metric we define in this work, so that it induces a metric between merge trees;
+  • $\text{size}(T_{f})$ being the cardinality of the edge set of $T_{f}$.
+Equation 2 represents a stability property in line with our expectations: since our metric is a summation of all the costs of the local modifications needed to align two trees we have that the to control the distance between trees with the sup norm between functions, we need a constant that scales linearly with the size of the trees.
+The interleaving distance, instead, is focused on determining the maximal local modifications to turn one tree into the other and it is well known to be universally stable, in the sense that is the biggest distance which is always bounded by the sup norm between the functions.
+Moreover, we have already pointed out that the $1$-Wasserstein distance between persistence diagrams satisfies an analogous relation w.r.t.
+the bottleneck distances:
+$d_{B}(D_{f},D_{g})\leq W_{1}(D_{f},D_{g})\leq(\#D_{f}+\#D_{g})\parallel f-g\parallel_{\infty}.$
+As a consequence, our metric is better at picking up differences (analogously to Wasserstein distances w.r.t bottleneck distance for persistence diagrams).
+On the other hand, this also means that is it more sensitive to noise.
+Building on these considerations, in Pegoraro (2024a) it is argued that such stability properties may be preferable to universal stability properties, in which the operator from functions to topological representation is $1$-Lipschitz.
+Moreover, such comparison is framed in analogy to the bias-variance tradeoff in statistical modeling.
+We further observe that, since merge trees can be viewed as a special case of Reeb graphs, the same considerations we made for $d_{I}$ also apply to all the metrics discussed in Bollen (2022)—including those introduced in Bauer et al.
+(2014); Di Fabio and Landi (2016); Bauer et al.
+(2020)—due to their metric equivalence with the interleaving distance.
+Lastly, Pegoraro (2024b) considers dendrograms enriched with Betti numbers of (path-connected components of) sublevel sets of functions, similarly to Equation 1. It is therein shown that:
+$d_{E}(\varphi_{f}^{\Theta_{p}},\varphi_{g}^{\Theta_{p}})\leq(\#\text{Crit}(f)+\#\text{Crit}(g))2C\parallel f-g\parallel_{\infty},$
+with:
+  • $f,g$ being two tame (Chazal et al., 2016) real valued functions defined on the same path-connected topological space;
+  • $\varphi_{f}^{\Theta_{p}}$ and $\varphi_{g}^{\Theta_{p}}$ are the dendrograms associated to (enriched) the persistent sets of the sublevel set filtrations of $f$ and $g$;
+  • $d_{E}$ is an adaptation of metric we define in this work, so that it induces a metric between functions defined on display posets of persistent sets;
+  • $C>0$ is an upper bound to the values of the Betti numbers;
+  • $\text{Crit}(f)$ is the set of values where the homology of the sublevel sets of $f$ changes, in any dimension.
+To conclude, we note that, if we consider the persistent homology of the sublevel set filtration of $f$ and $g$, in dimensions $i=0,\ldots,N$, and employ the persistence diagrams $D_{f}^{i}$ and $D_{g}^{i}$, we have:
+$\sum_{i=0}^{N}W_{1}(D^{i}_{f},D^{i}_{g})\leq(\#\text{Crit}(f)+\#\text{Crit}(g))\parallel f-g\parallel_{\infty}.$
+6 Computing the Edit Distance
+In these last section we develop an algorithm to compute the edit distance for dendrograms.
+6.1 Decomposition Properties
 Following ideas found in Hong et al.
@@
 Then, in Section 6.2, we formulate each of the split-up calculations as binary optimization problems, which are then aggregated recursively in a bottom-up fashion in Section 6.3·
-Name $T_{2}$ the only representative without order $2$ vertices inside the equivalence class of $T$.
-One can always suppose that a dendrogram is given without order $2$ vertices.
+Name $T_{2}$ the only representative without degree $2$ vertices inside the equivalence class of $T$.
+One can always suppose that a dendrogram is given without degree $2$ vertices.
 Thus, for notational convenience, from now on we suppose $T=T_{2}$ and $T^{\prime}=T^{\prime}_{2}$.
-To help us in the calculations define a particular subset of mappings $M_{2}(T,T^{\prime})\subset\operatorname*{Mapp}(T,T^{\prime})$.
-Definition 20
-A mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ has maximal ghostings if $(v,"G")\in M$ if and only if $v$ is of order $2$ after the deletions in $T$ and, similarly ("$G",w)\in M$ if and only if $w$ is of order $2$ after the deletions in $T^{\prime}$.
-A mapping $M\in\operatorname*{Mapp}(T,T^{\prime})$ has minimal deletions if $(v,"D")\in M$ only if neither $v$ nor $father(v)$ are of order $2$ after applying all the other deletions in $T$ and, similarly, ("$D",w)\in M$ only if $w$ is not of order $2$ after applying all the other deletions in $T^{\prime}$.
-We collect all mappings with maximal ghostings in the set $M_{\mathcal{G}}(T,T^{\prime})$ and paths with minimal deletions in the set $M_{\mathcal{D}}(T,T^{\prime})$.
-Lastly: $M_{2}(T,T^{\prime}):=M_{\mathcal{G}}(T,T^{\prime})\bigcap M_{\mathcal{D}}(T,T^{\prime})$.
-In other words with mappings in $M_{2}(T,T^{\prime})$ we are always eliminating all the order $2$ vertices which arise from deletions and we are not deleting edges which we can shrink.
-The following lemma then applies.
-Lemma 1
-$\min\{cost(M)\mid M\in\operatorname*{Mapp}(T,T^{\prime})\}=\min\{cost(M)\mid M\in M_{2}(T,T^{\prime})\}$
-In addition to that, we consider some particular subsets of $E_{T}\times E_{T^{\prime}}$ which play a fundamental role in what follows.
+We consider some particular subsets of $E_{T}\times E_{T^{\prime}}$ which play a fundamental role in what follows.
 Recall that, using $E_{T}\cong V_{T}-\{r_{T}\}$, we can induce $\pi_{T}:E_{T}\times E_{T^{\prime}}\rightarrow V_{T}$.
@@
   • (A1)
-    the points in $\pi_{T}(M^{*})$ form antichains in $V_{T}$ (and the same for $\pi_{T^{\prime}}(M^{*})$ in $V_{T^{\prime}}$), with respect to the partial order given by $father>son$.
+    the points in $\pi_{T}(M^{*})$ form antichains in $V_{T}$ (and the same for $\pi_{T^{\prime}}(M^{*})$ in $V_{T^{\prime}}$), with respect to the partial order given by $parent>child$.
 This means that any two distinct vertices of $T$ (respectively of $T^{\prime}$) which appear in $M^{*}$ are incomparable with respect to “$>$”;
@@
 Consider now $M^{*}\in\mathcal{C}^{*}(T,T^{\prime})$.
-Starting from such set of couples we build a set of edits which form a “partial” mapping between $T$ and $T^{\prime}$: each couple $(x,y)\in M^{*}$ means that we do not care of what lies below $x\in V_{T}$ and $y\in V_{T^{\prime}}$ and we need to define edits only for the other vertices.
+Starting from such set of pairs we build a set of edits which form a “partial” mapping between $T$ and $T^{\prime}$: each pair $(x,y)\in M^{*}$ means that we do not care of what lies below $x\in V_{T}$ and $y\in V_{T^{\prime}}$ and we need to define edits only for the other vertices.
 The vertices below $x$ and $y$ will be taken care separately.
@@
 $\Omega_{M^{*}}(x)=\min\{LCA(x,x^{\prime})\mid x^{\prime}\in\pi_{T}(M^{*})\text{ and }x\neq x^{\prime}\}$
-And the same holds for $y\in V_{T^{\prime}}\cap\pi_{T^{\prime}}(M^{*})$.
-For ease of notation we will often avoid explicit reference to $M^{*}$ and write directly $\Omega(x)$.
-Note that $\zeta_{x}^{\Omega(x)}$ is a segment (see 9).
+And the same holds for $y\in V_{T^{\prime}}\cap\pi_{T^{\prime}}(M^{*})$. For ease of notation we will often avoid explicit reference to $M^{*}$ and write directly $\Omega(x)$.
 With these bits of notation, given $M^{*}\in\mathcal{C}^{*}(T,T^{\prime})$, we build the “partial” mapping $\alpha(M^{*})$: is a mapping that ignores all the vertices which lie below $x\in V_{T}$ and $y\in V_{T}$ if $(x,y)\in M^{*}$.
@@
   2. 2.
-    if there is not $x\in V_{T}$ such that $v<\Omega(x)$ or $v>\Omega(x)$, then $(v,"D")\in\alpha(M^{*})$;
+    if there is not $x\in V_{T}$ such that $v<\Omega(x)$ or $v>\Omega(x)$, then $(v,\mathfrak{D})\in\alpha(M^{*})$;
   3. 3.
-    if there is $x\in V_{T}$ such that $v>\Omega(x)$ then $(v,"D")\in\alpha(M^{*})$;
+    if there is $x\in V_{T}$ such that $v>\Omega(x)$ then $(v,\mathfrak{D})\in\alpha(M^{*})$;
   4. 4.
@@
     1. (a)
-      if $v\in\zeta_{x}^{\Omega(x)}$ then $(v,"G")\in\alpha(M^{*})$
+      if $v\in\zeta_{x}^{\Omega(x)}$ then $(v,\mathfrak{G})\in\alpha(M^{*})$
     2. (b)
-      if $v<v_{i}$ for some $v_{i}\in\zeta_{x}^{\Omega(x)}=\{v_{0}<v_{1}<\ldots<v_{n}\}$ then $(v,"D")\in\alpha(M^{*})$;
+      if $v<v_{i}$ for some $v_{i}\in\zeta_{x}^{\Omega(x)}=\{v_{0}<v_{1}<\ldots<v_{n}\}$ then $(v,\mathfrak{D})\in\alpha(M^{*})$;
     3. (c)
@@
 This means that each $v\in V_{T}$ satisfies one and only one of the above conditions and so $\alpha(M^{*})$ is well defined.
-The idea behind $\alpha(M^{*})$ is that, for all couples $(x,y)\in M^{*}$, we want to make the ghostings to turn the segments $\zeta_{x}^{\Omega(x)}$ and $\zeta_{y}^{\Omega(y)}$ respectively into the edges $(x,\Omega(x))$ and $(y,\Omega(y))$, and then shrink one in the other.
+The idea behind $\alpha(M^{*})$ is that, for all pairs $(x,y)\in M^{*}$, we want to make the ghostings to turn the paths $\zeta_{x}^{\Omega(x)}$ and $\zeta_{y}^{\Omega(y)}$ respectively into the edges $(x,\Omega(x))$ and $(y,\Omega(y))$, and then shrink one in the other.
 As we already anticipated, $\alpha(M^{*})$ takes care of all the vertices in $V_{T}$ and $V_{T^{\prime}}$, a part from the sets $\cup_{(x,y)\in M^{*}}\{x^{\prime}\in E_{T}\mid x^{\prime}<x\}$ and $\cup_{(x,y)\in M^{*}}\{y^{\prime}\in E_{T^{\prime}}\mid y^{\prime}<y\}$.
@@
 $d_{E}(T,T^{\prime})=\min_{M^{*}\in\mathcal{C}^{*}(T,T^{\prime})}\sum_{(x,y)\in M^{*}}d_{E}(sub_{T}(x),sub_{T^{\prime}}(y))+cost(\alpha(M^{*})).$ (3)
-Figure 7: Given two weighted trees $T$ (left) and $T^{\prime}$ (right) - which are the same of Figure 3(a) - we consider $T_{x}=sub_{T}(x)$ and $G_{y}=sub_{T^{\prime}}(y)$ and we use 3 to compute $d_{E}(T_{x},T_{y})$, as in the algorithm in Section 6.2.
+Figure 5: Given two weighted trees $T$ (left) and $T^{\prime}$ (right) - which are the same of Figure 3(a) - we consider $T_{x}=sub_{T}(x)$ and $T^{\prime}_{y}=sub_{T^{\prime}}(y)$ and we use 3 to compute $d_{E}(T_{x},T^{\prime}_{y})$, as in the algorithm in Section 6.2.
 The set $M^{*}=\{(v,w),(v^{\prime},w^{\prime}),(v^{\prime\prime},w^{\prime\prime})\}$ satisfies (A1),(A2).
 The set $\alpha(M^{*})$ is made by the deletions/insertions indicated by the red edges, the ghostings/splittings indicated by the yellow and the shrinkings given by edges of the same color, different from red.
-To obtain $\widetilde{T}_{x}$ and $\widetilde{T}^{\prime}_{y}$ as in 7 all the black vertices covered by shaded regions must be deleted.
-6.2 Computing the Edit Distance for Trees: Dynamical Binary Linear Programming problems
+To obtain $\widetilde{T}_{x}$ and $\widetilde{T}^{\prime}_{y}$ as in 6 all the black vertices covered by shaded regions must be deleted.
+6.2 Dynamical Binary Linear Programming problems
 We want to use 3 to write a dynamical, binary linear optimization algorithm to calculate $d_{E}$: by translating 3 into a Binary Linear Programming (BLP) problem, we obtain a single step in a bottom-up procedure.
@@
 Along with keeping the notation defined in Section 6.1, define $T_{x}:=sub_{T}(x)$ and $T_{y}:=sub_{T^{\prime}}(y)$, $N_{x}:=\dim(T_{x})=\#E_{T}$ and $N_{y}:=\dim(T_{y})=\#E_{T^{\prime}}$.
-Moreover, given $v\in V_{T_{x}}$, the sequence $v_{0}=v<v_{1}<\ldots<r_{T}$ indicates the points in the segment $\zeta_{v}$.
+Moreover, given $v\in V_{T_{x}}$, the sequence $v_{0}=v<v_{1}<\ldots<r_{T}$ indicates the points in the path $\zeta_{v}$.
 Thus $v_{i}$ will be a vertex $v_{i}>v$.
@@
 We would like to find $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ minimizing Equation (3) for $T_{x}$ and $T_{y}$, but this is a difficult task.
-In fact, as evident in the construction of $\alpha(M^{*})$, a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ has the role of pairing segments: if $(v,w)\in M^{*}$, then the segments $\zeta_{v}^{\Omega(v)}$ and $\zeta_{w}^{\Omega(w)}$ are shrunk one on the other by $\alpha(M^{*})$.
-However, the points $\Omega(v)$ and $\Omega(w)$ depend on the whole set $M^{*}$, and not simply on the couple $(v,w)$.
-Modelling such global dependence gives rise to non-linear relationships between coupled points, and so leading to a non linear cost function, in terms of points interactions, to be minimized.
+In fact, as evident in the construction of $\alpha(M^{*})$, a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$ has the role of pairing paths: if $(v,w)\in M^{*}$, then the paths $\zeta_{v}^{\Omega(v)}$ and $\zeta_{w}^{\Omega(w)}$ are shrunk one on the other by $\alpha(M^{*})$.
+However, the points $\Omega(v)$ and $\Omega(w)$ depend on the whole set $M^{*}$, and not simply on the pair $(v,w)$.
+Modeling such global dependence gives rise to non-linear relations between paired points, and so leading to a non linear cost function, in terms of points interactions, to be minimized.
 For this reason we “weaken” the last term in Equation (3), allowing also mappings different from $\alpha(M^{*})$ to be built from $M^{*}$.
@@
 where $\beta(M^{*})$ is such that:
-  • $\beta(M^{*})\in M_{\mathcal{G}}(\widetilde{T}_{x},\widetilde{T}_{y})$ (using the notation established in 7, replacing $T$ and $T^{\prime}$ with $T_{x}$ and $T_{y}$ respectively);
-  • the set of vertices coupled by $\beta(M^{*})$ is exactly $M^{*}$: $M^{*}=\beta(M^{*})\bigcap V_{T_{x}}\times V_{T_{y}}$.
-Since, by construction $M^{*}=\alpha(M^{*})\bigcap V_{T}\times V_{T^{\prime}}$ and by 7, $\alpha(M^{*})\in M_{2}(\widetilde{T}_{x},\widetilde{T}_{y})$, minimizing Equation (2) or Equation (3) gives the same results.
+  • $\beta(M^{*})\in M_{\mathcal{G}}(\widetilde{T}_{x},\widetilde{T}_{y})$ (using the notation established in 6, replacing $T$ and $T^{\prime}$ with $T_{x}$ and $T_{y}$ respectively);
+  • the set of vertices paired by $\beta(M^{*})$ is exactly $M^{*}$: $M^{*}=\beta(M^{*})\bigcap V_{T_{x}}\times V_{T_{y}}$.
+Since, by construction $M^{*}=\alpha(M^{*})\bigcap V_{T}\times V_{T^{\prime}}$ and by 6, $\alpha(M^{*})\in M_{2}(\widetilde{T}_{x},\widetilde{T}_{y})$, minimizing Equation (3) or Equation (4) gives the same results.
 6.2.3 Setup and Variables
@@
 We use $\delta$ to indicate the matrix of variables $(\delta^{v,w}_{i,j})_{v,w,i,j}$.
-The mapping $\beta(M^{*})$ is built according to the variables $\delta^{v,w}_{i,j}$ with value $1$: we write a constrained optimization problem such that having $\delta^{v,w}_{i,j}=1$ means pairing the segments $\zeta_{v}^{v_{i+1}}$ (that is, the segment which starts with $(v,v_{1})$ and ends with $(v_{i},v_{i+1})$) and $\zeta_{w}^{w_{j+1}}$, and shrinking one in the other in the induced mapping.
-In order to pair and shrink the segments $\zeta_{v}^{v_{i+1}}=\{v=v_{0},v_{1},\ldots,v_{i+1}\}$ and $\zeta_{w}^{w_{j+1}}$ we need to collect some edits in the set $\beta(M^{*})$ adding the following edits:
+The mapping $\beta(M^{*})$ is built according to the variables $\delta^{v,w}_{i,j}$ with value $1$: we write a constrained optimization problem such that having $\delta^{v,w}_{i,j}=1$ means pairing the paths $\zeta_{v}^{v_{i+1}}$ (that is, the path which starts with $(v,v_{1})$ and ends with $(v_{i},v_{i+1})$) and $\zeta_{w}^{w_{j+1}}$, and shrinking one in the other in the induced mapping.
+In order to pair and shrink the paths $\zeta_{v}^{v_{i+1}}=\{v=v_{0},v_{1},\ldots,v_{i+1}\}$ and $\zeta_{w}^{w_{j+1}}$ we need to collect some edits in the set $\beta(M^{*})$ adding the following edits:
   • all the points $v_{k}\in\zeta_{v}^{v_{i+1}}$ with $10<k<i+$ 1 are ghosted, that is $(v_{k},G)\in\beta(M^{*})$;
@@
 6.2.4 Constraints
-Clearly, the set of $\delta^{v,w}_{i,j}$ with value $1$ does not always produce $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$: for instance segments could be paired multiple times.
+Clearly, the set of $\delta^{v,w}_{i,j}$ with value $1$ does not always produce $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$: for instance paths could be paired multiple times.
 To avoid such issues, we build a set of constraints for the variable $\delta$.
@@
 If $\delta\in\mathcal{K}$:
-  • the couples $(v,w)$ such that $\delta^{v,w}_{i,j}=1$ define a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$;
+  • the pairs $(v,w)$ such that $\delta^{v,w}_{i,j}=1$ define a set $M^{*}\in\mathcal{C}^{*}(T_{x},T_{y})$;
   • the edits induced by all $\delta^{v,w}_{i,j}=1$ give a mapping $\beta(M^{*})$ in $M_{``G^{\prime\prime}}(\widetilde{T}_{x},\widetilde{T}_{y})$.
-With $\widetilde{T}_{x},\widetilde{T}_{y}$ being obtained from $T_{x}$ and $T_{y}$ as in 7.
+With $\widetilde{T}_{x},\widetilde{T}_{y}$ being obtained from $T_{x}$ and $T_{y}$ as in 6.
 Remark 6
@@
 Having built a mapping $\beta(M^{*})$ using the binary variables, we want to define a cost functions which computes the cost of such mapping depending on $\delta$.
-As before, consider $v\in E_{T_{x}}$ and interpret $\delta^{v,w}_{i,j}=1$ as coupling the segments $\zeta_{v}^{v_{i+1}}$ and $\zeta_{w}^{w_{j+1}}$; then $v$ is coupled with some $w\in E_{T_{y}}$ if $C(v):=\sum_{i,w,j}\delta^{v,w}_{i,j}=1$ and is ghosted if $G(v):=\sum_{\{i,v^{\prime}\mid v\in\zeta_{v^{\prime}}^{v^{\prime}_{i}}\}}\sum_{w,j}\delta^{v^{\prime},w}_{i,j}=1$.
+As before, consider $v\in E_{T_{x}}$ and interpret $\delta^{v,w}_{i,j}=1$ as pairing the paths $\zeta_{v}^{v_{i+1}}$ and $\zeta_{w}^{w_{j+1}}$; then $v$ is paired with some $w\in E_{T_{y}}$ if $C(v):=\sum_{i,w,j}\delta^{v,w}_{i,j}=1$ and is ghosted if $G(v):=\sum_{\{i,v^{\prime}\mid v\in\zeta_{v^{\prime}}^{v^{\prime}_{i}}\}}\sum_{w,j}\delta^{v^{\prime},w}_{i,j}=1$.
 The vertex $v$ is instead deleted if $D(v):=1-C(v)-G(v)=1$.
@@
 Note that the above operations are taken inside the editable space $X$.
-The function which computes the cost given by coupled points is therefore:
+The function which computes the cost given by paired points is therefore:
 $F^{C}(\delta):=\sum_{v,w,i,j}\Delta^{v,w}_{i,j}\cdot\delta^{v,w}_{i,j}$
@@
 where the “norm” of a tree $T$ is $\mid\mid T\mid\mid=\sum_{e\in E_{T}}d_{X}(\varphi(e),0)$.
-Finally, one must take into account the values of $d_{E}(T_{v},T_{w})$, whenever $v$ and $w$ are coupled; this information is contained in $(W_{xy})_{v,w}$:
+Finally, one must take into account the values of $d_{E}(T_{v},T_{w})$, whenever $v$ and $w$ are paired; this information is contained in $(W_{xy})_{v,w}$:
 $F^{S}(\delta):=\sum_{v,w}(W_{xy})_{v,w}\cdot\left(\sum_{i,j}\delta^{v,w}_{i,j}\right)$
@@
 $d_{E}(T_{x},T_{y})=\min_{\delta\in\mathcal{K}}F^{C}(\delta)+F^{D}(\delta)-F^{-}(\delta)+F^{S}(\delta)$ (7)
-Proof  The contribution of coupled points is $F^{C}(\delta)$ and the contribution of deleted points is $F^{D}(\delta)-F^{-}(\delta)$.
+Proof  The contribution of paired points is $F^{C}(\delta)$ and the contribution of deleted points is $F^{D}(\delta)-F^{-}(\delta)$.
 The cost of $\beta(M^{*})$ is: $F^{\beta}(\delta):=F^{C}(\delta)+F^{D}(\delta)-F^{-}(\delta)$.
-Lastly, $F^{S}(\delta)$ takes into account the value of $d_{E}(T_{v},T_{w})$, if $v$ and $w$ are coupled.
-By 3, combined with 8, the solution of the following optimization problem:
+Lastly, $F^{S}(\delta)$ takes into account the value of $d_{E}(T_{v},T_{w})$, if $v$ and $w$ are paired.
+By 3, combined with 7, the solution of the following optimization problem:
 $\min_{\delta\in\mathcal{K}}F^{S}(\delta)+F^{\beta}(\delta)$ (8)
@@
 A solution to Problem Equation 7 exists because the minimization domain is finite and there are admissible values; it is not unique in general.
-6.3 Computing the Edit Distance for Trees: Bottom-Up Algorithm
+6.3 Bottom-Up Algorithm
 In this section the results obtained in Section 6.1 and the formulation established in Section 6.2 are used to obtain the algorithm implemented to compute the metric $d_{E}$ between dendrograms.
@@
 Given $x\in V_{T}$, define $\operatorname*{len}(x)$ to be the number of vertices in $\zeta_{x}$ and $\operatorname*{len}(T)=\max_{v\in V_{T}}\operatorname*{len}(v)$.
-Therefore, $\operatorname*{lvl}(x)=\operatorname*{len}(T)-\operatorname*{len}(x)$.
-Lastly, $\operatorname*{lvl}_{T}(n)=\{v\in V_{T}\mid\operatorname*{lvl}(v)=n\}$
+Then, we set $\operatorname*{lvl}(x)=\operatorname*{len}(T)-\operatorname*{len}(x)$.
 The key property is that: $\operatorname*{lvl}(x)>\operatorname*{lvl}(v)$ for any $v\in sub(x)$.
-Thus, if $W_{xy}$ is known for any $x\in\operatorname*{lvl}_{T}(n)$ and $y\in\operatorname*{lvl}_{T^{\prime}}(m)$, then for any $v$, $w$ in $V_{T}$, $V_{T^{\prime}}$ such that $\operatorname*{lvl}(v)<n$ and $\operatorname*{lvl}(w)<m$, $W_{vw}$ is known as well.
+Thus, if $W_{xy}$ is known for any $x\in\operatorname*{lvl}_{T}^{-1}(n)$ and $y\in\operatorname*{lvl}_{T^{\prime}}^{-1}(m)$, then for any $v$, $w$ in $V_{T}$, $V_{T^{\prime}}$ such that $\operatorname*{lvl}(v)<n$ and $\operatorname*{lvl}(w)<m$, $W_{vw}$ is known as well.
 With this notation we can write down Algorithm 1.
@@
 Proof  In a full binary tree structure, at each level $l$ we have $2^{l}$ vertices.
-Let $L=\operatorname*{len}(T)$ and $L^{\prime}=len(T^{\prime})$.
-We have that, for any vertex $v\in V_{T}$ at level $l$, the cardinality of the path from $v$ to any of the leaves in $sub_{T}(v)$ is ${L-l}$ and the number of leaves in $sub_{T}(v)$ is $2^{L-l}$.
+Let $L=\operatorname*{len}(T)$ and $L^{\prime}=\operatorname*{len}(T^{\prime})$.
+We have that, for any vertex $v\in\operatorname*{lvl}(l)^{-1}$, the cardinality of the path from $v$ to any of the leaves in $sub_{T}(v)$ is ${L-l}$ and the number of leaves in $sub_{T}(v)$ is $2^{L-l}$.
 So, given $v\in V_{T}$ at level $l$ and $w\in V_{T^{\prime}}$ at level $l^{\prime}$, to calculate $d_{E}(sub_{T}(v),sub_{T^{\prime}}(w))$ (having already $W_{vw}$) we need to solve a binary linear problem with $2^{L-l}\cdot(L-l)\cdot 2^{L^{\prime}-l^{\prime}}\cdot(L^{\prime}-l^{\prime})$ variables and $2^{L-l}+2^{L^{\prime}-l^{\prime}}$ linear constraints.
@@
 Remark 8
-We end the manuscript with a remark about the computational cost of $d_{E}$ as described in 10.
-We found this result surprising to some extent.
-If we consider the latest implementation of the classical edit distance between unlabelled and unordered trees obtained with BLP (Hong et al., 2017), then the complexity of the two metrics is very similar: the classical edit distance can be computed by solving $O(N\cdot M)$ BLP problems with $O(N\cdot M)$ variables and $O(\log_{2}(M)+\log_{2}(N))$ constraints - $O(N+M)$ if we count also the constraints restricting the integer variables to $\{0,1\}$, as the authors of Hong et al.
+We end the manuscript with a remark about the computational cost of $d_{E}$ as described in 9.
+We find this result surprising to some extent.
+If we consider a recent implementation of the classical edit distance between unlabeled and unordered trees obtained with BLP (Hong et al., 2017), then the complexity of the two metrics is very similar: the classical edit distance can be computed by solving $O(N\cdot M)$ BLP problems with $O(N\cdot M)$ variables and $O(\log_{2}(M)+\log_{2}(N))$ constraints - $O(N+M)$ if we count also the constraints restricting the integer variables to $\{0,1\}$, as the authors of Hong et al.
 (2017) do.
@@
 7 Conclusions
-We develop a novel framework to work with graphs that arise as topological summaries in the context of topological data analysis.
+We develop a novel framework to work with tree that arise as topological summaries in the context of topological data analysis, possibly enriched with some abstract weights as a result of different pipelines.
 These kind of summaries are increasingly frequent and require some particular care to be properly handled by a metric.
-We define an edit distance employing some novel edit operations, proving that such definition induces a metric up to a certain equivalence relationship between graphs.
-For situations in which the graphs possess the structure of a tree then, by exploiting the ordering properties of trees, we produce a binary linear programming algorithm to compute such metric.
+We define an edit distance employing some novel edit operations, proving that such definition induces a metric up to a certain equivalence relation between weighted trees.
+By exploiting the ordering properties of trees, we produce a binary linear programming algorithm to compute such metric.
 The modifications that we make w.r.t.
@@
 Still, the computational procedure we describe here is used in separate applied works, as mentioned in the introduction.
-These additional modifications are essential for the stability properties assessed in Pegoraro (2024a), justifying the higher computational cost w.r.t.
+These additional modifications are essential for the stability properties assessed in Pegoraro and Secchi (2024); Pegoraro (2024b), justifying the higher computational cost w.r.t.
 other edit distances for merge trees, which, instead, are unstable.
@@
   • we think that the properties of the editable spaces can be relaxed; however, the algorithm presented in this manuscript may need to be adapted to the properties of the chosen weight space;
-  • Reeb Graphs can be investigated theoretically and algorithmically, perhaps following the decomposition presented in Stefanou (2020);
-  • likewise, computational procedures and theoretic investigations should be considered for particular finite sub-posets for $\mathbb{R}^{n}$;
-  • as motivated by 11 we believe that polynomial time approximations can be obtained for computing $d_{E}$ between dendrograms;
+  • we would like to extend this edit distance outside tree-shaped graphs, encompassing Reeb graphs (Biasotti et al., 2008; De Silva et al., 2016) and their combinatorial approximation, namely mapper graphs (Singh et al., 2007) are studied (Carriere et al., 2018; Carriere and Oudot, 2018), perhaps following the decomposition presented in Stefanou (2020).
+We also mention that works on GEDs like (Ambauen et al., 2003; Lerouge et al., 2016) could be of great help in developing such generalization;
+  • as motivated by 8 we believe that polynomial time approximations can be obtained for computing $d_{E}$ between dendrograms;
   • further applications and case studies exploiting the generality of the weights we consider on the edges should be considered and explored.
@@
 This work was carried out as part of my PhD Thesis, under the supervision of Professor Piercesare Secchi.
+I also acknowledge the support of the Wallenberg AI, Autonomous Systems and Software Program (WASP), and of the SciLifeLab and Wallenberg National Program for Data-Driven Life Science (DDLS), which fund the project: Topological Data Analysis of Functional Genome to Find Covariation Signatures.
 Outline of the Appendix
@@
 Here we present in details the first steps of the Algorithm 1, used to calculate the distance between two merge trees.
-We consider the following couple of weighted trees.
+We consider the following pair of weighted trees.
 Let $(T,w_{T})$ be the tree given by: $V_{T}=\{a,b,c,d,r_{T}\}$, $E_{T}=\{(a,d),(b,d),(d,r_{T}),(c,r_{T})\}$ and $w_{T}(a)=w_{T}(b)=w_{T}(d)=1$, $w_{T}(c)=5$; the tree $(T^{\prime},w_{T^{\prime}})$ instead, is defined by: $V_{T^{\prime}}=\{a^{\prime},b^{\prime},c^{\prime},d^{\prime},r_{T^{\prime}}\}$, $E_{T^{\prime}}=\{(a^{\prime},d^{\prime}),(b^{\prime},d^{\prime}),(d^{\prime},r_{T^{\prime}}),(c^{\prime},r_{T^{\prime}})\}$ and $w_{T^{\prime}}(a)=1$, $w_{T^{\prime}}(b)=w_{T^{\prime}}(c)=2$ and $w_{T^{\prime}}(d)=3$.
@@
 Step: $n=m=0$
-This step is trivial since we only have couples between leaves, like $(a,a^{\prime})$, which have trivial subtrees and thus $d_{E}(sub_{T}(a),sub_{T^{\prime}}(a^{\prime}))=0$.
+This step is trivial since we only have pairs between leaves, like $(a,a^{\prime})$, which have trivial subtrees and thus $d_{E}(sub_{T}(a),sub_{T^{\prime}}(a^{\prime}))=0$.
 Step: $n=m=1$
 The points $x\in V_{T}$ with $\operatorname*{lvl}_{T}(x)\leq 1$ are $\{a,b,c,d\}$ and the points $y\in V_{T^{\prime}}$ with $\operatorname*{lvl}_{T^{\prime}}(y)\leq 1$ are $\{a^{\prime},b^{\prime},c^{\prime},d^{\prime}\}$.
-Thus the couples $(x,y)$ which are considered are: $(d,d^{\prime})$, $(d,a^{\prime})$, $(d,b^{\prime})$, $(d,c^{\prime})$ and $(a,d^{\prime})$, $(b,d^{\prime})$, $(c,d^{\prime})$.
-The couples between leaves, like $(a,a^{\prime})$ have already been considered.
-Couple: $(d,d^{\prime})$
+Thus the pairs $(x,y)$ which are considered are: $(d,d^{\prime})$, $(d,a^{\prime})$, $(d,b^{\prime})$, $(d,c^{\prime})$ and $(a,d^{\prime})$, $(b,d^{\prime})$, $(c,d^{\prime})$.
+The pairs between leaves, like $(a,a^{\prime})$ have already been considered.
+Pair: $(d,d^{\prime})$
 Let $T_{d}=sub_{T}(d)$ and $T_{d^{\prime}}=sub_{T^{\prime}}(d^{\prime})$.
@@
 A solution is given by $\delta_{0,0}^{a,a^{\prime}}=\delta_{0,0}^{b,b^{\prime}}=1$ and $\delta_{0,0}^{a,b^{\prime}}=\delta_{0,0}^{b,a^{\prime}}=0$, which entails $F^{C}(\delta)=1$, $F^{D}(\delta)=0$, $F^{-}(\delta)=0$ and $F^{S}(\delta)=0$ and $d_{E}(T_{d},T_{d^{\prime}})=1$.
-Couple: $(d,a^{\prime})$
+Pair: $(d,a^{\prime})$
 Obviously: $d_{E}(sub_{T}(d),sub_{T^{\prime}}(a^{\prime}))=\mid\mid sub_{T}(d)\mid\mid$.
-All the couples featuring a leaf and an internal vertex (that is, a vertex which is not a leaf), such as $(d,b^{\prime})$, $(a,d^{\prime})$ etc.
+All the pairs featuring a leaf and an internal vertex (that is, a vertex which is not a leaf), such as $(d,b^{\prime})$, $(a,d^{\prime})$ etc.
 behave similarly.
@@
 The points $x\in V_{T}$ with $\operatorname*{lvl}_{T}(x)\leq 2$ are $\{a,b,c,d,r_{T}\}$ and the points $y\in V_{T^{\prime}}$ with $\operatorname*{lvl}_{T^{\prime}}(y)\leq 2$ are $\{a^{\prime},b^{\prime},c^{\prime},d^{\prime},r_{T^{\prime}}\}$.
-Thus the couples $(x,y)$ which are considered are $(d,r_{T^{\prime}})$, $(r_{T},d^{\prime})$, $(r_{T},r_{T^{\prime}})$ and then the trivial ones: $(r_{T},a^{\prime})$, $(r_{T},b^{\prime})$, $(r_{T},c^{\prime})$ and $(a,r_{T^{\prime}})$, $(b,r_{T^{\prime}})$, $(c,r_{T^{\prime}})$.
-Some couples have already been considered and thus are not repeated.
-Couple: $(d,r_{T^{\prime}})$
+Thus the pairs $(x,y)$ which are considered are $(d,r_{T^{\prime}})$, $(r_{T},d^{\prime})$, $(r_{T},r_{T^{\prime}})$ and then the trivial ones: $(r_{T},a^{\prime})$, $(r_{T},b^{\prime})$, $(r_{T},c^{\prime})$ and $(a,r_{T^{\prime}})$, $(b,r_{T^{\prime}})$, $(c,r_{T^{\prime}})$.
+Some pairs have already been considered and thus are not repeated.
+Pair: $(d,r_{T^{\prime}})$
 Let $T_{d}=sub_{T}(d)$ and $T^{\prime}=sub_{T^{\prime}}(r_{T^{\prime}})$.
@@
 This value of $\delta$ entails $F^{C}(\delta)=2+1$, $F^{D}(\delta)=w_{T^{\prime}}(a^{\prime})+w_{T^{\prime}}(b^{\prime})=1+2$, $F^{-}(\delta)=\mid\mid sub_{T^{\prime}}(d^{\prime})\mid\mid=3$ and $F^{S}(\delta)=d_{E}(sub_{T}(a),sub_{T^{\prime}}(d^{\prime}))=\mid\mid sub_{T^{\prime}}(d^{\prime})\mid\mid=3$, and thus $d_{E}(T_{d},T^{\prime})=3+3-3+3=6$.
-Couple: $(r_{T},d^{\prime})$
-This and the other couples are left to the reader.
+Pair: $(r_{T},d^{\prime})$
+This and the other pairs are left to the reader.
 B Proofs
@@
 By Bubenik and Elchesen (2022) we have:
-$W_{1}(D,D^{\prime})=W_{1}(\sum_{x\in D}\{x\},\sum_{y\in D^{\prime}}\{y\})=\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in D}k(x)-\sum_{y\in D^{\prime}}k(y)).$
+$W_{1}([D],[D^{\prime}])=W_{1}(\sum_{x\in[D]}\{x\},\sum_{y\in[D^{\prime}]}\{y\})=\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in[D]}k(x)-\sum_{y\in[D^{\prime}]}k(y)).$
 At this point $(P3)$ and $(P4)$ follow easily by, respectively:
-  • replacing $D^{\prime}$ with the empty diagram and noting:
-    $\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in D}k(x))=\sum_{x\in D}\min_{y\in\Delta}\parallel x-y\parallel_{1};$
-  • observing that when we have $W_{1}(D+D^{\prime\prime},D^{\prime}+D^{\prime\prime})$ the contributions by the elements in $D^{\prime\prime}$ when computing:
-    $\sum_{x\in D+D^{\prime\prime}}k(x)-\sum_{y\in D^{\prime}+D^{\prime\prime}}k(y)$
+  • replacing $[D^{\prime}]$ with the empty diagram and noting:
+    $\sup_{k\in\text{Lip}(\mathbb{R}^{2}_{\Delta})}(\sum_{x\in[D]}k(x))=\sum_{x\in[D]}\min_{y\in\Delta}\parallel x-y\parallel_{1};$
+  • observing that when we have $W_{1}([D+D^{\prime\prime}],[D^{\prime}+D^{\prime\prime}])$ the contributions by the elements in $[D^{\prime\prime}]$ when computing:
+    $\sum_{x\in[D+D^{\prime\prime}]}k(x)-\sum_{y\in[D^{\prime}+D^{\prime\prime}]}k(y)$
     appear in both sums, erasing themselves out.
 $\blacksquare$
-Proof of 3.
-We prove $s+s^{\prime}\in\mathcal{S}(G)$ and that it is well defined on the equivalence classes of $\mathcal{S}(G)$.
-Because of the intersection condition then $s+s^{\prime}$ is injective.
-By construction it is also continuous.
-Moreover it is automatically an homeomorphism onto its image in a small enough neighbourhood of any point except for $1/2$.
-Let $(a,b)$ be an open neighbourhood of $1/2$ in $[0,1]$.
-We know that $s+s^{\prime}$ restricted to $(a,1/2]$ and to $[1/2,b)$ is in both cases an embedding.
-Which implies that $s+s^{\prime}(1/2)$ is going to be an internal point in $(s+s^{\prime})((a,b))$, which is then an open set.
-Thus $s+s^{\prime}$ is an homeomorphism also if restricted to $(a,b)$.
-Associativity follows easily.
-In case $G$ is a directed graph we are done as considering $s^{\prime\prime}\sim s$ and $s^{\prime\prime\prime}\sim s^{\prime}$ would still give:
-  • $s(1)=s^{\prime}(0)$;
-  • $Im(s)\cap Im(s^{\prime})=s(1)$.
-In case $G$ is an undirected graph, then $Im(s^{\prime\prime})\cap Im(s^{\prime\prime\prime})=s^{\prime\prime}(1)=s(1)$ holds in any case.
-However we may have $s^{\prime\prime}(1)\neq s^{\prime\prime\prime}(0)$.
-But, if that is the case, it is enough to change $s^{\prime\prime}(t)$ with $s^{\prime\prime}(1-t)$ and/or $s^{\prime\prime\prime}(t)$ with $s^{\prime\prime\prime}(1-t)$.
-$\blacksquare$
-Proof of 1.
-To lighten the notation we use the following symbols:
-  • the symbols $e_{d}$ and $e^{\prime-1}_{d}$ respectively mean that the edge $e$ is deleted and $e^{\prime}$ is inserted;
-  • if instead we write $s_{g}$ and $s^{\prime-1}_{g}$ we imply that the segment $s$ is involved in a ghosting operation and the segment $s^{\prime}$ is the result of the splitting of an edge;
-  • lastly, the shrinking induced by $(e,e^{\prime})$ is called $e_{w,w^{\prime}}$ with $w$ being the original weight function, and $w^{\prime}$ the weight function after the shrinking.
-Note that all edits are defined on a single edge a part from ghostings and splittings, which in general are associated to a segment.
-Suppose that $\gamma$ is a finite edit path.
-This means that $\gamma$ is the composition of a finite set of edits.
-We indicate such ordered composition with $\gamma=\prod_{i=0}^{N}e_{i}$ with $e_{i}$ edit operation.
-We would like to change the order of the edit operations without raising the cost and changing the extremes of the edit path.
-This is not always possible.
-However we can work it around in the useful cases using properties (P1)-(P4).
-In particular, we would like to know when we can commute a generic edit $e_{i}$ in the following situations:
-  • $e_{d}\circ e_{i}$ and $e_{i}\circ e^{-1}_{d}$
-  • $s_{g}\circ e_{i}$ and $e_{i}\circ e^{-1}_{g}$.
-Moreover we want to reduce the edit path to max one shrinking/deletion/insertion for any edge of $G$ and $G^{\prime}$.
-We divide the upcoming part of the proof in subsections, each devoted to different combinations of edits.
-Inverse Operations
-We point out the following fact which will be used in the proof.
-Suppose we have a couple of edits $e_{i}\circ e_{j}$ that can be replaced with the couple $e^{\prime}_{i}\circ e^{\prime}_{j}$.
-Now consider $e^{-1}_{j}\circ e^{-1}_{i}$:
-$e^{-1}_{j}\circ e^{-1}_{i}=(e_{i}\circ e_{j})^{-1}=(e^{\prime}_{i}\circ e^{\prime}_{j})^{-1}=e^{\prime-1}_{j}\circ e^{\prime-1}_{i}.$ (8)
-$e_{d}$ and $e_{d}^{-1}$
-Note that, when we delete or insert one edge $e$, we are modifying the graph only at the level of its open star $\mathcal{N}(e)$. Consider then the following scenarios:
-  • $e_{d}\circ s_{g}$ so that the ghostings induced by $s$ result in the edge $e$, which is then deleted.
-This composition of edits can be safely replaced with the deletions of all the edges in $s$, conserving the length of the path by (P3).
-If the ghostings defined by $s$ do not involve $e$ then we can safely commute the operations.
-  • $e_{d}\circ s_{g}^{-1}$, with $e$ being a new edge originated by the splitting induced by $s_{g}^{-1}$.
-This composition of edits can clearly be replaced with a shrinking of the original edge and then a suitably defined splitting (which avoids the edge $e$) without changing the path length - by (P3) and (P4).
-  • $e_{d}\circ e_{w,w^{\prime}}$ can be replaced by $e_{d}$ potentially diminishing the length of the path, but surely not raising it by (P1).
-  • $e_{w,w^{\prime}}\circ e_{d}^{-1}$ can be replaced with an insertion directly with weight $w(e)$, possibly shortening the path (P1).
-  • $s^{\prime}_{g}\circ e_{d}^{-1}$follows from Equation 8.
-  • $s^{\prime-1}_{g}\circ e_{d}^{-1}$ follows from Equation 8.
-$s_{g}$ and $s_{g}^{-1}$
-Like in the previous case, we only take into account transformations regarding edges in the open star of the ghostings/splittings.
-  • $s_{g}\circ s^{\prime}_{g}$, with $s$ and $s^{\prime}$ being segments whose interiors have non-empty intersection, can be merged in a unique ghosting by (P2).
-Note that, in case $G$ and $G^{\prime}$ are directed graphs, then the directions of the two segments are compatible.
-Otherwise the intersection would be empty.
-We avoid repeating this same comment in the upcoming points.
-  • $s_{g}\circ s^{\prime-1}_{g}$, with $s$ and $s^{\prime}$ being segments whose interiors have non-empty intersection, can commute provided we define carefully the splittings $s^{\prime-1}_{g}$ by (P2), - avoiding the splittings which lie in the intersection.
-  • $s_{g}\circ e_{w,w^{\prime}}$ with $e\subset Im(s)$.
-This means that we are shrinking an edge before merging it with others via some ghostings.
-However, we can achieve the same result, without increasing the path length, by making the ghostings at first, and then shrinking the resulting edge - by (P1)-(P4).
-  • $e_{w,w^{\prime}}\circ s_{g}^{-1}$ follows from Equation 8.
-$e_{w,w^{\prime}}$
-  • $v_{w^{\prime},w^{\prime\prime}}\circ e_{w,w^{\prime}}$ can be replaced by $e_{w,w^{\prime\prime}}$ which is either conserving or shortening the path (P1).
-  • note that $v_{w,w^{\prime}}^{-1}=v_{w^{\prime},w}$.
-Thanks to these properties we can take a given path $\gamma=\prod_{i=0,...,N}e_{i}$ and modify the edit operations in order to obtain the following situation:
-  • the first operations are all in the form $e_{d}$; this can be achieved because $e_{d}\circ-$ can be always rearranged so that the deletions are always done first, potentially shortening the path.
-Of course there can be only one deletion for each edge of $G$;
-  • then we have all the edits in the form $s_{g}$; since $s_{g}\circ-$ is exchangeable when we do not have $s_{g}$ pre-composed by some deletion, this is not a problem.
-Observe that all ghostings (and so all splittings) involving segments whose interiors have non empty intersection can be joined in a unique ghosting operation;
-  • in the same way we can put last all the paths in the form $e_{d}^{-1}$ and before them the ones in the form $s_{g}^{-1}$;
-  • in the middle we are left with the shrinking paths, which by construction are on edges which were segments in the original graphs.
-Since we can substitute $e_{w,w^{\prime}}\circ v_{w^{\prime},w^{\prime\prime}}$ with $e_{w,w^{\prime\prime}}$, we can obtain just one single shrinking for each edge and this means that the shrinkings are pairing uniquely segments in the original graphs.
-Thus the edit path
-$\overline{\gamma}=(\gamma_{d}^{G^{\prime}})^{-1}\circ(\gamma_{g}^{G^{\prime}})^{-1}\gamma_{s}^{G}\circ\gamma_{g}^{G}\circ\gamma_{d}^{G},$
-with:
-  • $\gamma_{d}^{G}=\prod e_{d}$
-  • $\gamma_{g}^{G}=\prod s_{g}$
-  • $\gamma_{s}^{G}=\prod e_{w,w^{\prime}}$
-  • $(\gamma_{g}^{G^{\prime}})^{-1}=\prod s_{g}^{-1}$
-  • $(\gamma_{d}^{G^{\prime}})^{-1}=\prod e_{d}^{-1}$
-is such that $\gamma(T)=\overline{\gamma}(G)=G^{\prime}$ and $cost(\overline{\gamma})\leq cost(\gamma)$.
-The key point is that $\overline{\gamma}$ can be easily realized as a mapping whose couples are given by the segments paired by the shrinkings.
-$\blacksquare$
-Proof of 6.
+Proof of 5.
 For the construction of the merge tree associated to a persistent set see Pegoraro (2024b).
@@
 Thus we have $V-\{r_{T_{S}}\}\hookrightarrow D_{S}$.
-Moreover, this embedding is an embedding of posets if we consider $T_{S}$ with the relationship generated by $child<father$.
+Moreover, this embedding is an embedding of posets if we consider $T_{S}$ with the relation generated by $child<parent$.
 Consider now $(t_{j},p)\in D_{S}-V-\{r_{T_{S}}\}$: this happens if there are $t_{i}<t_{j}$ and $t_{k}>t_{j}$ such that for every $\varepsilon_{i}>0$ such that $t_{j}-\varepsilon_{i}>t_{i}$ and for every $\varepsilon_{k}>0$ such that $t_{j}+\varepsilon_{k}<t_{k}$, we have #$\#S(t_{j}\leq t_{j}+\varepsilon_{k})^{-1}(S(t_{j}\leq t_{j}+\varepsilon_{k})(p))=1$ and #$\#S(t_{j}-\varepsilon_{i}\leq t_{j})^{-1}(p)=1$.
@@
 The same for any edge $(t_{j},p)\rightarrow(t,s)$ with $(t,s)\neq(t_{k},S(t_{i}\leq t_{k})(p))$.
-Thus $(t_{j},p)$ induces an order $2$ vertex in $\mathcal{G}(D_{S})$.
+Thus $(t_{j},p)$ induces a degree $2$ vertex in $\mathcal{G}(D_{S})$.
 To conclude the proof we just need to prove that $\mathcal{G}(D_{S})$ is a tree.
-In fact, if this is the case, then we have a poset map $V-\{r_{T}\}\hookrightarrow V_{\mathcal{G}(D_{S})}$ so that all the vertices not in the image of this map are of order $2$.
+In fact, if this is the case, then we have a poset map $V-\{r_{T}\}\hookrightarrow V_{\mathcal{G}(D_{S})}$ so that all the vertices not in the image of this map are of degree $2$.
 I.e.
-the map is an isomorphism of posets onto the image, which is the poset obtained by removing all order $2$ vertices from $\mathcal{G}(D_{S})$.
+the map is an isomorphism of posets onto the image, which is the poset obtained by removing all degree $2$ vertices from $\mathcal{G}(D_{S})$.
 We know that $\mathcal{G}(D_{S})$ is a DAG which has, by construction one maximal element.
@@
 Proof of 1.
-Any order $2$ vertex which is not ghosted is paired with another order $2$ vertex. Ghosting both of them does not increase the cost of the mapping.
+Any degree $2$ vertex which is not ghosted is paired with another degree $2$ vertex. Ghosting both of them does not increase the cost of the mapping.
 $\blacksquare$
-Proof of 7.
+Proof of 2.
+To lighten the notation we use the following symbols:
+  • the edit induced by $(v,\mathfrak{D})$ is called $v_{d}$ and $v^{-1}_{d}$ stands for $(\mathfrak{D},v)$.
+  • the edit induced by $(v,\mathfrak{G})$ is called $v_{g}$ and $v^{-1}_{g}$ stands for $(\mathfrak{G},v)$.
+  • the edit induced by $(v,v^{\prime})$ is called $v_{\varphi,\varphi^{\prime}}$ with $\varphi$ being the original weight function, and $\varphi^{\prime}$ the weight function after the shrinking.
+We know that the set of finite edit paths between two dendrograms is nonempty.
+Suppose that $\gamma$ is a finite edit path.
+This means that $\gamma$ is the composition of a finite set of edits.
+We indicate such ordered composition with $\gamma=\prod_{i=0}^{N}e_{i}$ with $e_{i}$ edit operation.
+We would like to change the order of the edit operations without raising the cost and changing the endpoints of the edit path.
+This is not always possible.
+However we can work it around in the useful cases using properties (P1)-(P4).
+In particular, we would like to know when we can commute a generic edit $e_{i}$ in the following situations:
+  • $v_{d}\circ e_{i}$ and $e_{i}\circ v^{-1}_{d}$
+  • $v_{g}\circ e_{i}$ and $e_{i}\circ v^{-1}_{g}$.
+Moreover we want to reduce the edit path to max one edit for any vertex of $T$ and $T^{\prime}$.
+We divide the upcoming part of the proof in subsections, each devoted to different combinations of edits.
+Inverse Operations
+We point out the following fact, which will be used often in the proof.
+Suppose we have a pair of edits $e_{i}\circ e_{j}$ that can be replaced with the pair $e^{\prime}_{i}\circ e^{\prime}_{j}$.
+Now consider $e^{-1}_{j}\circ e^{-1}_{i}$:
+$e^{-1}_{j}\circ e^{-1}_{i}=(e_{i}\circ e_{j})^{-1}=(e^{\prime}_{i}\circ e^{\prime}_{j})^{-1}=e^{\prime-1}_{j}\circ e^{\prime-1}_{i}.$ (9)
+$v_{d}$ and $v_{d}^{-1}$
+When we delete or insert one vertex, we are modifying the tree structure at the level of its parent and its children.
+Therefore, we are only taking into account operations involving the parent of the considered vertex, on the vertex himself or on the children of the deleted/inserted vertex.
+  • $v_{d}\circ v^{\prime}_{g}$: if $v$ is a child of $v^{\prime}$, can be safely replaced with $v_{d}\circ v^{\prime}_{d}$.
+Instead of ghosting the parent and then deleting the whole edge, we can delete both edges one by one; conserving the length of the path (P3).
+If, instead, $v$ is parent of $v^{\prime}$, we can safely commute the operations.
+  • $v_{d}\circ v_{g}^{-1}$: the two edits can be replaced with $v^{\prime}_{\varphi,\varphi^{\prime}}$ with $v^{\prime}$ parent of $v$ (after the insertion) and $\varphi^{\prime}$ properly defined not to raise the cost of the path.
+In fact we are inserting $v$ on an edge and then deleting it.
+This can obviously be achieved by shrinking the original edge (without changing the path length - (P4)).
+  • $v_{d}\circ v^{\prime-1}_{g}$: if $v^{\prime}$ is parent of $v$ after the splitting, we can again replace the two edits with shrinking: instead of inserting a point in an edge, and deleting then the edge below, we can directly shrink the original edge (P4).
+If, instead, $v^{\prime}$ is inserted below $v$, we have the same situation, but seen from the point of view of the vertex which is the child of $v^{\prime}$ after the splitting.
+  • $v_{d}\circ v_{\varphi,\varphi^{\prime}}$: the two edits can be replaced by $v_{d}$ potentially diminishing the length of the path, but surely not raising it (P1).
+  • $v^{\prime}_{g}\circ v_{d}^{-1}$: if $v^{\prime}$ is the parent of $v$, this edit can be replaced with just $v^{\prime}_{\varphi,\varphi^{\prime}}$ with appropriate weights: we are inserting an edge under a vertex which (in this case) becomes of degree two and is ghosted.
+We can directly modify the edge without changing the length of the path (P4).
+If, instead, $v^{\prime}$ becomes a child of $v$ after the insertion, we can simply shrink $v^{\prime}$ to obtain the same result without raising the cost (P4).
+  • $v^{\prime-1}_{g}\circ v_{d}^{-1}$: if $v^{\prime}$ is on the edge inserted with $v_{d}^{-1}$, we cannot commute these two edits, but we can replace them with two insertions: instead of inserting an edge and then splitting it, we can directly insert two smaller edges; without changing the cost of the path (P3).
+In all other situations, we can commute the two operations.
+  • $v_{\varphi,\varphi^{\prime}}\circ v_{d}^{-1}$: the two edits can be replaced with an insertion directly with weight $\varphi^{\prime}$, possibly shortening the path (P1).
+  • $v^{\prime-1}_{d}\circ v_{d}$: suppose that the parent of $v^{\prime}$, after its insertion, is the parent of $v$ before we apply both edits.
+If the children of $v^{\prime}$ are different from the children of $v$, this operations do not commute.
+If the children are the same, they can be changed with a shrinking of $v$, reducing the length of the path by at most $cost(v^{\prime-1}_{d})+cost(v_{d})$ (P1).
+$v_{g}$ and $v_{g}^{-1}$
+Like in the previous case, we only take into account transformations concerning the parent and the child of the added/ghosted degree two vertex.
+  • $v_{g}\circ v^{\prime}_{g}$: always commute (P2).
+  • $v_{g}\circ v^{\prime-1}_{g}$: always commute, provided that we define carefully the splitting $v^{\prime-1}_{g}$ (P2).
+  • $v_{g}\circ v_{\varphi,\varphi^{\prime}}$: with such operations, we are shrinking a vertex before ghosting it.
+However, we can achieve the same result, without increasing the path length, by ghosting the vertex at first, and then shrinking its child (P1)-(P4).
+  • $v^{\prime}_{\varphi,\varphi^{\prime}}\circ v_{g}^{-1}$: if either $v^{\prime}=v$, or $v$ is the parent of $v^{\prime}$, we can replace those edits with an appropriate shrinking of the vertex which is the child of $v$ after the splitting and an appropriate insertion of $v^{\prime}$, without changing the length of the path (P3)-(P4).
+  • $v_{g}\circ v^{\prime}_{d}$: if $v$ is the parent of $v^{\prime}$, such edits do not commute and can’t be replaced by a similar operation which inverts ghosting and deletion.
+$v_{w,w^{\prime}}$
+  • $v_{\varphi^{\prime},\varphi^{\prime\prime}}\circ v_{\varphi,\varphi^{\prime}}$: those edit can be replaced by $v_{\varphi,\varphi^{\prime\prime}}$ which is either conserving or shortening the path (P1).
+  • $v_{\varphi,\varphi^{\prime}}^{-1}$: we can proceed as in the previous point, since: $v_{\varphi,\varphi^{\prime}}^{-1}=v_{\varphi^{\prime},\varphi}$.
+Thanks to these properties, we can take a given path $\gamma=\prod_{i=0,...,N}e_{i}$ and modify the edit operations in order to obtain the following situation:
+  • the first operations are all in the form $v_{d}$; this can be achieved because $v_{d}\circ-$ can be always rearranged, potentially by changing the path (as shown before) and shortening it.
+Of course, there can be only one deletion for each vertex of $T$;
+  • then we have all the edits in the form $v_{g}$; since $v_{g}\circ-$ is exchangeable any time but when we have $v_{g}\circ v^{\prime}_{d}$, this is not a problem.
+Observe that all degree two vertices which were not deleted can be ghosted (at most one time);
+  • in the same way we can put last all the paths in the form $v_{d}^{-1}$ and before them $v_{g}^{-1}$.
+All the new vertices appearing with the insertion of edges and the splitting of edges with degree two vertices are all nodes which remain in $T^{\prime}$ and which are not further edited;
+  • in the middle we are left with the shrinking paths.
+Since we can substitute $v_{\varphi,\varphi^{\prime}}\circ v_{\varphi^{\prime},\varphi^{\prime\prime}}$ with $v_{\varphi,\varphi^{\prime\prime}}$, we can obtain at most one transformation on each vertex.
+Thus, we can find at least one edit path of the form:
+$\overline{\gamma}=(\gamma_{d}^{T^{\prime}})^{-1}\circ(\gamma_{g}^{T^{\prime}})^{-1}\gamma_{s}^{T}\circ\gamma_{g}^{T}\circ\gamma_{d}^{T}$
+with:
+  • $\gamma_{d}^{T}=\prod v_{d}$;
+  • $\gamma_{g}^{T}=\prod v_{g}$;
+  • $\gamma_{s}^{T}=\prod v_{\varphi,\varphi^{\prime}}$;
+  • $(\gamma_{g}^{T^{\prime}})^{-1}=\prod v_{g}^{-1}$;
+  • $(\gamma_{d}^{T^{\prime}})^{-1}=\prod v_{d}^{-1}$.
+As for the paths $\gamma_{M}$ described in Section 4.2, all the permutations of the edits inside each $\gamma_{d}^{T},\gamma_{g}^{T},\ldots,$ does not change its endpoints and its cost.
+Thus, $\overline{\gamma}$ is such that $\gamma(T)=\overline{\gamma}(T)=T^{\prime}$ (up to isomorphism) and $cost(\overline{\gamma})\leq cost(\gamma)$.
+The key point is that $\overline{\gamma}$ can be easily realized as a mapping made by the following elements:
+  • $(v,\mathfrak{D})$, $\forall v_{d}\in\gamma_{d}^{T}$;
+  • $(v,\mathfrak{G})$, $\forall v_{g}\in\gamma^{T}_{g}$;
+  • $(v,v^{\prime})$, $\forall v_{\varphi,\varphi^{\prime}}\in\gamma_{s}^{T}$, where $v^{\prime}$ is the edge in $T^{\prime}$ (more precisely, in $\gamma_{g}^{T^{\prime}}\circ\gamma_{d}^{T^{\prime}}(T^{\prime})$) associated to the shrinking $v_{\varphi,\varphi^{\prime}}$;
+  • $(\mathfrak{G},v)$, $\forall v^{-1}_{g}\in(\gamma_{g}^{T^{\prime}})^{-1}$;
+  • $(\mathfrak{D},v)$, $\forall v^{-1}_{d}\in(\gamma_{d}^{T^{\prime}})^{-1}$.
+$\blacksquare$
+Proof of 6.
 Condition (M2) coincide with condition (A2).
@@
 However, in this case, $v$ does not appear in $\widetilde{T}$, and thus (M1) is satisfied.
-Moreover, all and only order $2$ vertices, after the deletions, are ghosted, and (M4) follows .
+Moreover, after the deletions, all degree $2$ vertices are ghosted, and (M4) follows.
 $\blacksquare$
@@
 Let $M\in M_{2}(T,T^{\prime})$ such that $d_{E}(T,T^{\prime})=cost(M)$.
-We note that $father>son$ induces a partial order relationship also on the pairs given by coupled points in $M$: $(x,y)>(v,w)$ if $x>v$ and $y>w$.
+We note that $parent>child$ induces a partial order relation also on the pairs given by paired points in $M$: $(x,y)>(v,w)$ if $x>v$ and $y>w$.
 In fact, by property (M3), $x>v$ if and only if $y>w$.
-So we can select $(x_{i},y_{i})$, the maxima with respect to this partial order relationship.
+So we can select $(x_{i},y_{i})$, the maxima with respect to this partial order relation .
 Thus, we obtain $(x_{0},y_{0})$,…,$(x_{n},y_{n})$ which form an antichain (both in $V_{T}$ and $V_{T^{\prime}}$).
@@
 Let $\bar{x}=LCA(x_{i},x_{j})$.
-Since $\bar{x}>x_{i},x_{j}$, it is not coupled in $M$.
-Since $x_{i}$ and $x_{j}$ are coupled, $\bar{x}$ cannot be ghosted, so it is deleted in $M$.
+Since $\bar{x}>x_{i},x_{j}$, it is not paired in $M$.
+Since $x_{i}$ and $x_{j}$ are paired, $\bar{x}$ cannot be ghosted, so it is deleted in $M$.
 Any point $x$ above $\bar{x}$ is deleted for the same reasons.
@@
 $d(x_{i}\odot x,y_{i})\leq d(x_{i}\odot x,y_{i}\odot x)+d(y_{i}\odot x,y_{i})=d(x_{i},y_{i})+d(x,0)$
-by properties (P1)-(P4). Since $\alpha(M^{*})\in M_{2}(\widetilde{T},\widetilde{T}^{\prime})$ by 7, we have:
+by properties (P1)-(P4). Since $\alpha(M^{*})\in M_{2}(\widetilde{T},\widetilde{T}^{\prime})$ by 6, we have:
 $\sum_{(x,y)\in M^{*}}d_{E}(sub_{T}(x),sub_{T^{\prime}}(y))+cost(\alpha(M^{*}))\leq cost(M)$
@@
 $\blacksquare$
-Proof of 8.
+Proof of 7.
 Having fixed a leaf $l$, the constraint in Equation (5) allows for at most one path $\zeta_{v}^{v_{i+1}}\subset\zeta_{l}$ to be kept after the edits induced by all the variables equal to $1$.
@@
 Which means that for any $a\in V_{T_{x}}$, we are forcing that the vertex $a$ can be an internal vertex or lower extreme of at most one path $\zeta_{v}^{v_{i+1}}$ such that $\delta_{i,j}^{v,w}=1$.
-In other words if two “kept” segments $\zeta_{v}^{v_{i+1}}$ and $\zeta_{v^{\prime}}^{v^{\prime}_{i^{\prime}+1}}$ (i.e.
+In other words if two “kept” paths $\zeta_{v}^{v_{i+1}}$ and $\zeta_{v^{\prime}}^{v^{\prime}_{i^{\prime}+1}}$ (i.e.
 with $\delta_{i,j}^{v,w}=\delta_{i^{\prime},j}^{v^{\prime},w}=1$) intersect each other, it means that they just share the upper extreme $v_{i+1}=v^{\prime}_{i^{\prime}+1}$.
 These facts together imply that (if the constraints are satisfied) the edits induced on $T_{x}$ by $\delta^{v,w}_{i,j}=1$ and $\delta^{v^{\prime},w^{\prime}}_{i^{\prime},j^{\prime}}=1$ are always compatible: if $v^{\prime\prime}\in V_{sub_{T}(v_{i})}$ then it is not touched by (the edits induced by) $\delta^{v^{\prime},w^{\prime}}_{i^{\prime},j^{\prime}}=1$ (and the same exchanging the role of $v^{\prime}$ and $v$), if $v^{\prime\prime}$ is equal or above $v_{i+1}$ and/or $v^{\prime}_{i^{\prime}+1}$, then it is deleted in any case.
-Lastly, by noticing that if $\delta^{v,w}_{i,j}\in\Phi(v^{\prime})$ then $\delta^{v,w^{\prime}}_{i,j^{\prime}}\in\Phi(v^{\prime})$ for all other possible $w^{\prime}$ and $j^{\prime}$, we see that every segment $\zeta_{v}^{v_{i}}$ is paired with at most one segment $\zeta_{w}^{w_{j}}$, and viceversa.
-As a consequence, for any vertex $v^{\prime}$ in any of the tree structures, at most one point on the segment $\zeta_{v^{\prime}}$ is coupled in $M^{*}$, guaranteeing the antichain condition.
+Lastly, by noticing that if $\delta^{v,w}_{i,j}\in\Phi(v^{\prime})$ then $\delta^{v,w^{\prime}}_{i,j^{\prime}}\in\Phi(v^{\prime})$ for all other possible $w^{\prime}$ and $j^{\prime}$, we see that every path $\zeta_{v}^{v_{i}}$ is paired with at most one path $\zeta_{w}^{w_{j}}$, and viceversa.
+As a consequence, for any vertex $v^{\prime}$ in any of the tree structures, at most one point on the path $\zeta_{v^{\prime}}$ is paired in $M^{*}$, guaranteeing the antichain condition.
 Moreover, any point of $T_{x}$ which is in $\pi_{T_{x}}(M^{*})$ is assigned to one and only one point of $T^{\prime}_{y}$ and viceversa.